### v1

# Diversity as a Reward: Fine-Tuning LLMs on a Mixture of Domain-Undetermined Data

Abstract
Fine-tuning large language models (LLMs) using diverse datasets is crucial for enhancing their overall performance across various domains. In practical scenarios, existing methods based on modeling the mixture proportions of data composition often struggle with data whose domain labels are missing, imprecise or non-normalized, while methods based on data selection usually encounter difficulties in balancing multi-domain performance. To address these challenges, in this paper, we study the role of data diversity in enhancing the overall abilities of LLMs by empirically constructing contrastive data pools and theoretically deriving explanations for both inter- and intra-diversity. Building upon the insights gained, we propose a new method that gives the LLM a dual identity: an output model to cognitively probe and select data based on diversity reward, as well as an input model to be tuned with the selected data. Extensive experiments show that the proposed method notably boosts performance across domain-undetermined data and a series of foundational downstream tasks when applied to various advanced LLMs. We release our code and hope this study can shed light on the understanding of data diversity and advance feedback-driven data-model co-development for LLMs.
Machine Learning, ICML
1 Introduction
The advancement of large language models (LLMs), exemplified by open-source models such as the Llama3 series (Dubey et al., 2024), Qwen2 series (Yang et al., 2024), and the DeepSeek-V3 series (Liu et al., 2024a), along with closed-source models such as GPT-4 (Achiam et al., 2023), has revolutionized artificial intelligence by enhancing capabilities in foundational abilities such as common sense, reasoning, mathematics, and coding. Fine-tuning these models further optimizes their usability by enhancing their performance and aligning with specific human instructions and preferences  (Taori et al., 2023; Rafailov et al., 2023).
To cultivate comprehensive capabilities in LLMs, fruitful studies have explored preferable trade-offs between quality, quantity, and diversity of their training data  (Li et al., 2024b; Chen et al., 2024b; Qin et al., 2024; Zhao et al., 2024a). For example, methods focusing on data selection (Li et al., 2024a; Xia et al., 2024; Wang et al., 2023) and mixture (Ge et al., 2024) demonstrate promising capabilities to enhance model performance, particularly through semantic diversity  (Lu et al., 2024; Liu et al., 2024b).
However, real-world applications frequently encounter unlabeled data and difficulties in domain labeling (Ge et al., 2023), posing challenges for data mixture methodologies that take the domain tags as a prior, as well as the data selection approaches, which often prioritize quality over diversity especially with data sourced from quite different domains.
To leverage the best of both worlds, in this work, we propose a new fine-tuning method for LLMs named DaaR, which encourages the given model to learn diversity as a reward signal, and then to autonomously select domain-undetermined training data to achieve a theoretically informed trade-off to maximize diversity for overall model performance enhancement.
Our investigation begins by diving into the semantic diversity of LLM data, considering how to explicitly model it and how it can influence the overall performance of LLMs. We systematically construct contrastive data pools and conduct extensive empirical examinations of different distributional performances across various foundational LLM capabilities. Based on these observations, we provide theoretical discussions on a general assumption that a mixture of underlying component LLMs determines the data mixing effect, and the explanations regarding inter- and intra-diversity modeled from a data semantic perspective.
Figure 1: The t-SNE visualization of embeddings for data samples with different distributions on Qwen2-7B. (a) The data pool of all 40,000 samples, (b) Randomly selected subset, (c) Distribution of data farthest from other domain centroids based on Inter-Diversity, (d) Distribution of data closest to other domain centroids based on Inter-Diversity, (e) Distribution of data closest to its own domain centroid based on Inter-Diversity, (f) Distribution of data farthest from its own domain centroid based on Inter-Diversity.
Our theoretical insights illuminate that optimal solutions for overall model performance enhancement cannot be explicitly applied to domain-undetermined data. To address this, we integrate an external multi-layer perceptron (MLP) structure into the LLM to be tuned, acting as a trainable probe module conditioned on the inherent knowledge and weights of the given LLM. The probe module takes semantic entropy as a proxy measure of diversity, and output reward scores, and select suitable data subsets aligned with the LLM’s underlying mixture distributions for better diversity. This is achieved by training the probe module on synthetically generated domain-aware data by the model itself, and using the data after selection to further self-fine-tuning. Extensive experiments on various state-of-the-art (SOTA) LLMs like the Qwen and LLaMA series verify the effectiveness of DaaR in enhancing overall model capabilities over many SOTA baseline methods.
Our contributions are threefold:
  • We empirically and theoretically demonstrate how data plays a crucial role to enhance LLMs in semantic diversity, providing generalized modeling based on mixture of component models.
  • We propose a lightweight, self-rewarding method for selecting data in domain-undetermined scenarios, which utilizes synthesized model-aware data to maximize domain diversity and fine-tuning performance.
  • We verify that the proposed methods can effectively balance seven diverse benchmarks for improved overall performance when applied to representative LLMs, whereas other SOTA methods struggle in such challenging scenarios. Our code is open-sourced  [^1] https://github.com/modelscope/data-juicer/tree/DaaR to foster more in-depth understanding and future progress.
2 Related Work
Data Selection
Fine-tuning is a pivotal training paradigm for enhancing LLMs’ task-specific and domain-specific capabilities. Research has shown that a small set of instruction pairs can enable LLMs to follow major instructions effectively (Zhou et al., 2024; Chen et al., 2024a). This fine-tuning can be achieved through rule-based methods, which focus on attributes such as Error L2-Norm (Paul et al., 2021) and token length (Raffel et al., 2020). More recently, model-based heuristics have been explored, including methods based on instruction-following difficulty (Li et al., 2024a), GPT scoring (Chen et al., 2024d; Wettig et al., 2024), data model selection (Engstrom, 2024), and influence scores derived from loss (Xia et al., 2024).
Diversity & Data Mixing
A fundamental principle for LLMs is being able to handle diverse human requests, underscoring data diversity as essential for effective fine-tuning (Yu et al., 2022; Ding et al., 2023). Data diversity encompasses aspects such as data deduplication (Abbas et al., 2023), the coverage scope of tags (Lu et al., 2024), model-based diversity evaluation (Liu et al., 2024b; Zhang et al., 2024), and scaling properties (Song et al., 2024). Typically, data mixing approaches (Albalak et al., 2023; Ye et al., 2024; Ge et al., 2024) focus on adjusting the proportional weights of different domains to enhance model capabilities.
Our Position
From a modeling perspective, our work shares a focus with many data mixing methodologies regarding diversity measurement. However, we address a more challenging setting relevant to real-world applications, where available datasets often lack domain labels. This situation presents challenges for existing mixing methods in balancing label granularity and normalization, and in determining which domains should be considered as candidates for improved solution spaces and tuning performance.
Methodologically, our approach aligns with the data selection category but differs in both the learning process (utilizing a rewarding model) and the selection criteria (employing semantic entropy). Additionally, we revisit the applicability of existing selection methods in multi-domain scenarios. This area is under-explored and requires new insights to avoid biased modeling of diverse domain distributions—a topic we further investigate in this paper.
3 Motivational Observations
In this section, we will explore how data with different domain-specific diversity affects model capabilities through experimental data pools, featuring contrastive data distributions in terms of inter-domain and intra-domain diversity.
3.1 Seed Data Pools and Basic Setting
Data Pools and Sources
To explore how selecting data samples from extensive and diverse data repositories affects the foundational capabilities of LLMs, we construct various data pools and consistently fine-tune LLMs on them. We then analyze the performance changes and attribute these changes to the different controlled data pools.
The seed data pool is sourced from following datasets: Dolly-15k (Conover et al., 2023) for common sense, Cot-en (Chung et al., 2024) for reasoning, Math-Instruct (Yue et al., 2024) for mathematics, and Code-Alpaca (Chaudhary, 2023) for coding. Each dataset was randomly tailored to 10,000 entries, resulting in a combined data pool of 40,000 entries. Following instruction tuning practices (Zhou et al., 2024; Liu et al., 2024b), we then uniformly sample 8,000 data entries as a referenced data pool for random baseline. In subsequent sections, we will introduce how to construct other comparative pools with the same size to random pool.
Benchmarks
Aligned with representative capabilities of leading open-source LLMs, we select the following widely used evaluation sets: NQ (Kwiatkowski et al., 2019) and TriviaQA (Joshi et al., 2017) for common sense, Hellaswag (Zellers et al., 2019) for reasoning, GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) for mathematics, MBPP (Austin et al., 2021) and HumanEval (Chen et al., 2021) for coding. To evaluate the comprehensive performance of LLMs across domains, we employ the average metric (Avg) as the primary evaluation criterion.
Models & Implementation
To ensure the effectiveness and applicability of our empirical findings, we employ the Qwen2 series (Qwen2-7B & Qwen2.5-7B) (Yang et al., 2024) and the Llama3.1-8B (Dubey et al., 2024) as representative SOTA base models to be fine-tuned. All experiments are conducted under identical training and evaluation protocols with two independent repetitions. Full platform, training and evaluation details are provided in Appendix B.
3.2 Data Pools with Contrastive Distributions
To systematically analyze the impact of domain-specific diversity patterns on model capabilities, we propose a contrastive construction with three phases: (A) Foundational Definitions, (B) Diversity Metric Formulation, and (C) Distribution Synthesis.
(A) Foundational Definitions
Let the composite dataset $\mathcal{D}=\bigcup_{k=1}^{K}\mathcal{D}_{k}$ comprise $K=4$ distinct domains, where each domain subset $\mathcal{D}_{k}$ contains $N_{k}=|\mathcal{D}_{k}|$ samples. We represent each data instance through its semantic embedding $\mathbf{x}_{i}^{(k)}\in\mathbb{R}^{d}$ extracted from the Embedding layer of the pretrained LLM, capturing high-dimensional semantic features. The domain centroid $\mathcal{C}_{k}$ serves as the semantic prototype:
$\small\mathcal{C}_{k}=\frac{1}{N_{k}}\sum_{i=1}^{N_{k}}\mathbf{x}_{i}^{(k)}.$ (1)
This centroid-based representation enables geometric interpretation of domain characteristics in the embedding space. We dissect data diversity into two complementary aspects:
(B.1) Inter-Diversity
It quantifies the diversity between distinct domains through centroid geometry. For sample $\mathbf{x}_{i}^{(k)}$, its cross-domain similarity is measured by:
$\small\phi_{\text{inter}}(\mathbf{x}_{i}^{(k)})=\sum_{\begin{subarray}{c}j=1\\ j\neq k\end{subarray}}^{K}\frac{\mathbf{x}_{i}^{(k)}\cdot\mathcal{C}_{j}}{\|\mathbf{x}_{i}^{(k)}\|\|\mathcal{C}_{j}\|}.$ (2)
The global inter-diversity metric $\Phi_{\text{inter}}$ computes the expected pairwise centroid distance:
$\small\Phi_{\text{inter}}=\mathbb{E}_{k\neq l}\left[\|\mathcal{C}_{k}-\mathcal{C}_{l}\|_{2}\right]=\frac{1}{\binom{K}{2}}\sum_{k=1}^{K-1}\sum_{l=k+1}^{K}\|\mathcal{C}_{k}-\mathcal{C}_{l}\|_{2}.$ (3)
This formulation reflects a key insight: maximizing $\Phi_{\text{inter}}$ encourages domain separation, while minimization leads to overlapping representations. Fig. 1 demonstrates this continuum through t-SNE projections – high $\Phi_{\text{inter}}$ manifests as distinct cluster separation with clear margins (Fig. 1.(c)), whereas low values produce entangled distributions (Fig. 1.(d)). Full analysis is detailed in Appendix D.1.
(B.2) Intra-Diversity
Focusing solely on the separation between different domains may hinder the model’s ability to learn the knowledge specific to a given domain. Hence we measure variation within each domain. We calculate sample similarity to its domain center:
$\small\phi_{\text{intra}}(\mathbf{x}_{i}^{(k)})=\frac{\mathbf{x}_{i}^{(k)}\cdot\mathcal{C}_{k}}{\|\mathbf{x}_{i}^{(k)}\|\|\mathcal{C}_{k}\|}.$ (4)
And the domain-level variance metric is defined as:
$\small\Phi_{\text{intra}}^{(k)}=\frac{1}{N_{k}}\sum_{i=1}^{N_{k}}\|\mathbf{x}_{i}^{(k)}-\mathcal{C}_{k}\|_{2}^{2}.$ (5)
Controlled manipulation of $\Phi_{\text{intra}}$ reveals critical trade-offs: lower variance (tight clusters near $\mathcal{C}_{k}$) enhances domain-specific feature learning but risks over-specialization. Higher variance improves robustness at the cost of potential cross-domain interference. The visualization in Fig. 1 (e-f) illustrates this scenario that concentrated distributions exhibit sharp marginal peaks, while dispersed variants show overlapping density regions.
(C) Distribution Synthesis
For each domain $\mathcal{D}_{k}$, we compute sample-wise diversity scores $\{\phi_{\text{inter}}(\mathbf{x}_{i}^{(k)})\}_{i=1}^{N_{k}}$ and $\{\phi_{\text{intra}}(\mathbf{x}_{i}^{(k)})\}_{i=1}^{N_{k}}$. The construction proceeds via partition each $\mathcal{D}_{k}$ into 20% intervals based on the percentiles of $\phi_{\text{inter}}$ for inter-diversity control, and partition the $\phi_{\text{intra}}$ scores into 20% quantile intervals for intra-diversity control.
The 20% interval results in five choices of data selection per domain, parameterizing the trade-off between diversity preservation and domain specificity. As demonstrated in Appendix D.2 (Fig. 10 and Fig. 11), this quantization process induces measurable distribution shifts.
3.3 Experimental Observations
Table 1 presents comprehensive evaluations across seven benchmarks, where the notation Inter-Diversity (X-Y) indicates samples ranked in the top (100-Y)to (100-X)of cross-domain similarity scores. Due to space constraints, we present only the results for the top 20%, middle 20%, and bottom 20%. More results can be found in Appendix D.7.
Diversity-Aware Performance: Our diversity-controlled selections reveal two critical observations:
  • Varied Improvement Patterns: Both models demonstrate marked improvements over Raw distributions across all diversity conditions, but the effects of their improvements vary. For Llama3.1-8B, Inter-D (80-100) achieves 38.98 average accuracy (+3.12 over Raw), outperforming the Random baseline by 1.71, while Inter-D (0-20) is below Random of 0.09.
  • Model-dependent Performance Peak: Each model exhibits distinct optimal operating points along the diversity spectrum. Llama3.1-8B reaches peak performance at Inter-D (80-100) and Intra-D (80-100), suggesting complementary benefits from both diversity types. Qwen2-7B peaks in inter-type selection at low inter-diversity, while it peaks in intra-type selection at high intra-diversity.
These results show the promising potential of diversity-aware data selection, motivating us to further understand the performance variance more formally and propose principled solutions to adaptively achieve the performance peaks.
Practical Challenges
Despite existing positive improvements on overall performance, the optimal distribution parameters exhibit model-dependent variability. This parameter sensitivity suggests the existence of multiple local optima in the diversity-performance landscape. Two constraints merit consideration for real-world applications. (1) Label Dependency: The studied heuristic strategies Inter-Diversity and Intra-Diversity currently require domain-labeled data for centroid calculation. (2) Distribution Transiency: The optimal diversity parameters (e.g., 80-100 vs. 40-60) show sensitivity across tasks and models, necessitating automated and potentially costly configuration search.
Table 1: Performance of Llama3.1-8B and Qwen2-7B on various downstream task benchmarks under different constructed Inter-Diversity (Inter-D) and Intra-Diversity (Intra-D) distributions.
Models Distribution Common Sense Reasoning Mathematics Coding Avg
NQ TriviaQA Hellaswag GSM8K MATH MBPP HumanEval
Llama3.1-8B Raw 14.13 65.90 74.62 54.80 7.90 5.00 28.66 35.86
Random 21.99 64.83 74.72 55.70 14.50 5.10 24.09 37.27
Inter-D (0-20) 19.28 65.79 74.44 54.90 6.50 4.30 35.06 37.18
Inter-D (40-60) 23.70 65.14 74.86 56.40 17.15 5.00 24.40 38.09
Inter-D (80-100) 23.76 64.43 75.20 56.40 15.05 4.50 33.54 38.98
Intra-D (0-20) 22.08 65.08 75.00 54.70 16.20 4.40 33.54 38.71
Intra-D (40-60) 22.12 64.74 74.87 54.00 16.00 6.00 27.44 37.88
Intra-D (80-100) 19.78 64.77 74.51 56.50 13.00 5.20 37.50 38.75
Qwen2-7B Raw 8.03 59.58 73.00 78.00 5.70 5.00 60.98 41.47
Random 13.28 58.27 73.00 75.35 35.36 52.20 63.72 53.02
Inter-D (0-20) 15.18 59.28 73.34 74.50 34.94 53.10 68.60 54.13
Inter-D (40-60) 14.62 58.58 73.35 72.50 34.50 52.50 61.28 52.47
Inter-D (80-100) 9.30 57.72 73.14 74.60 28.00 51.30 63.42 51.07
Intra-D (0-20) 12.64 58.54 73.35 75.10 8.75 51.10 61.59 48.72
Intra-D (40-60) 15.24 58.57 73.12 74.70 32.50 51.80 64.02 52.85
Intra-D (80-100) 11.91 57.88 73.29 75.00 36.05 52.50 66.16 53.25
4 Theoretical Analysis and Insights
In this section, we first formalize the optimization dynamics of multi-domain data aggregation for enhancing LLMs’ comprehensive capabilities. Building on the empirical observations from Sec. 3, we then establish theoretical connections between distributional diversity and emergent model behaviors. Finally, we demonstrate the fundamental limitation of explicit diversity optimization in undetermined-domain scenarios.
4.1 Ideal Optimization Formulation
During the training process of LLMs, various sources of labeled data are often collected to enhance models’ capabilities across multiple dimensions. Let $\mathcal{K}=\{1,\ldots,K\}$ represent the set of labeled domains, where $K$ is determined by real-world task requirements (e.g., $K=4$ in our experiments). For each domain $k\in\mathcal{K}$, data samples $\mathbf{x}_{i}^{(k)}\in\mathcal{D}_{k}$ are generated according to a distribution $\mathcal{D}_{k}$. The data distributions are assumed mutually distinct across different domains, leading to an ideal I.I.D. per-domain hypothesis: each domain $k$ corresponds to an independent model $h_{k}\in\mathcal{H}$. The standard optimization goal thus can be:
$\small\forall k\in\mathcal{K},\quad\min_{h_{k}\in\mathcal{H}}\mathcal{L}_{\mathcal{D}_{k}}(h_{k}),$ (6)
where $\mathcal{L}_{\mathcal{D}_{k}}(h_{k})=\mathbb{E}_{(x,y)\sim\mathcal{D}_{k}}\left[l(h_{k}(x),y)\right]$ is the empirical risk given a specific loss function $l$.
Counterintuitive Findings
If the distributions $\{\mathcal{D}_{k}\}_{k=1}^{K}$ are strictly I.I.D., the risk $\mathcal{L}_{\mathcal{D}_{k}}(h_{k})$ should depend only on $\mathcal{D}_{k}$. However, numerous previous studies (Zhao et al., 2024a; Tirumala et al., 2023) have demonstrated the presence of synergistic and antagonistic effects between different datasets. This is further demonstrated in Sec. 3 (Fig. 1) that certain domain mixtures lead to catastrophic performance degradation (e.g. Intra-D (0-20) in MATH with Qwen2). This contradicts the above I.I.D. hypothesis, suggesting that the simplified and conventional formulation in Eq (6) fails to capture cross-domain interference.
4.2 Introducing Mixture of Underlying Distributions
Inspired by the counterintuitive findings, we posit that each domain distribution $\mathcal{D}_{k}$ arises from the mixture of $M$ latent distributions $\{\tilde{\mathcal{D}}_{m}\}_{m=1}^{M}$ representing foundational LLM capabilities ($M\ll K$ in practice). This leads to our core statistical assumption:
Assumption 4.1 (Latent Capability Structure).
For each observed domain $k\in\mathcal{K}$, its data distribution decomposes into $M$ latent capability distributions:
$\small\mathcal{D}_{k}=\sum_{m=1}^{M}\pi_{km}\tilde{\mathcal{D}}_{m},\quad\sum_{m=1}^{M}\pi_{km}=1,$ (7)
where $\tilde{\mathcal{D}}_{m}$ indicates the $m$-th foundational capability. The weights $\pi_{km}$ reflect domain-specific capaility composition.
To analyze how data optimization interacts with latent capability distributions, we further posit that each foundational capability admits an optimal configuration:
Assumption 4.2 (Capability Optimality).
Each foundational capability admits a unique optimal predictor:
$\small\exists h_{\theta_{m}^{*}}\in\mathcal{H}\text{ s.t. }\theta_{m}^{*}=\mathop{\mathrm{argmin}}_{\theta}\mathbb{E}_{(x,y)\sim\tilde{\mathcal{D}}_{m}}[l(h_{\theta}(x),y)].$ (8)
The loss $l$ is strongly convex, holding for cross-entropy or MSE, where $h_{\theta_{m}}$, termed the component model, represents a model’s manifestation of specific foundational capabilities.
This leads to our key result formalizing how real-world training data activates latent capabilities via component models:
Proposition 4.3.
Under Assumption 4.1-4.2, the minimizer of the cross-domain average risk
$\small\min_{\{\theta_{m}\}}\mathbb{E}_{k\sim\mathcal{K}}\left[\mathcal{L}_{\mathcal{D}_{k}}\left(\textstyle\sum_{m=1}^{M}\pi_{km}h_{\theta_{m}}\right)\right]$ (9)
admits a solution $\{\theta_{m}^{*}\}_{m=1}^{M}$ where each $h_{\theta_{m}^{*}}$ optimizes the corresponding foundational capability $\tilde{\mathcal{D}}_{m}$.
Remark Proposition 4.3 reveals an underlying fact: optimal or suboptimal solutions can be achieved through data selection to enhance the overall model performance. This is further validated by the role of the coefficients $\pi_{km}$, which act as capability selectors to configure domain-specific behavior, demonstrating the impact of data composition.
4.3 Diversity Influence on the Optimization
Building upon Proposition 4.3, we analyze the intrinsic relationship between diversity and capability composition. The mixture coefficients $\pi_{km}$ inherently govern both inter- and intra-domain diversity formulated in Sec. 3.2, differing in their geometric properties in the latent capability space.
Centroids with Coefficients Let $\mathcal{C}_{m}\in\mathbb{R}^{d}$ denote the centroid vector of latent capability $\tilde{\mathcal{D}}_{m}$ in the embedding space. The domain centroid $\mathcal{C}_{k}$ can be expressed as:
$\small\mathcal{C}_{k}=\sum_{m=1}^{M}\pi_{km}\mathcal{C}_{m}.$ (10)
This linear combination implies that the inter- and intra-diversity are determined by $\pi_{km}$ configurations:
Proposition 4.4 (Diversity Decomposition).
The inter-diversity metric decomposes into:
$\small\Phi_{\text{inter}}=\sum_{m=1}^{M}\sum_{n=1}^{M}\lambda_{m}\lambda_{n}\|\mathcal{C}_{m}-\mathcal{C}_{n}\|_{2},$ (11)
where $\lambda_{m}=\mathbb{E}_{k}[\pi_{km}]$. And the intra-diversity satisfies:
$\small\Phi_{\text{intra}}^{(k)}\leq\sum_{m=1}^{M}\pi_{km}\|\mathcal{C}_{m}-\mathcal{C}_{k}\|_{2}^{2}+\mathbb{E}_{m\sim\pi_{k}}[\text{Var}(\tilde{\mathcal{D}}_{m})].$ (12)
The inter-diversity result follows from substituting $\mathcal{C}_{k}=\sum\pi_{km}\mathcal{C}_{m}$ into $\mathbb{E}_{k\neq l}\|\mathcal{C}_{k}-\mathcal{C}_{l}\|$ and applying Jensen’s inequality. The intra-diversity bound combines the variance decomposition within each latent capability. Full proof is provided in Appendix C.2.
4.4 Theoretical Insights
Collectively, results in this section provide a unified framework to explain the varied patterns and existence of peak performance observed in Sec. 3.3, providing theoretical insights and guidance for designing our following algorithm dealing with undetermined data.
Diversity’s Influence on LLMs’ Capability: Based on Proposition 4.4, we conjecture that the Inter-/Intra-Diversity, which can be explicitly modeled by $\pi$, also governs $\pi$-variation control for component model mixing and thus the final overall performance. Proposition 4.3 further suggests that improper diversity levels may induce model suboptimality by constraining overall capability, consistent with the distribution-dependent empirical results in Sec. 3.
Robust Diversity Selection Strategies: Based on Proposition 4.4, it can be observed that averaging the $\pi_{km}$ helps to mitigate the missing of underlying optimal predictor $\theta_{m}^{*}$. From this view, the random baseline can gain relatively stable improvements in expectation. A more robust yet simple strategy can be derived: ensemble across candidate models trained on multiple data pools with distribution-varied inter- and intra-diversity, e.g., via techniques like voting or model weights averaging. This can be corroborated by results in Table 1, such as the case for Llama3.1-8B.
Challenges in Undetermined-Domain Scenarios: However, all the selection strategies discussed so far face intractable solutions in undetermined domains with explicit domain labels, as Proposition 4.4 technically revealed: (1) ambiguous domain boundaries, e.g., blended instruction types; (2) partial overlaps in latent capabilities; and (3) absence of centroid priors $\{\mathcal{C}_{m}\}$. These intrinsic constraints motivate our model-aware diversity reward in the following Sec. 5, which harnesses domain-diversity awareness into reward modeling without requiring explicit parameterization for such structural assumptions.
5 DaaR: Diversity as a Reward
To address the challenges identified and leverage the insights gained in previous Sec. 3 and Sec. 4, we establish a data selection method DaaR guided by diversity-aware reward signals. It comprises three key components illustrated in subsequent sections: (1) model-aware centroid synthesis, (2) two-stage training with reward probe, and (3) diversity-driven data selection.
5.1 Model-Aware Training Data
Model-Aware Centroid Construction
The proposed method initiates with centroid self-synthesis through a two-phase generation process to address two fundamental challenges: (1) eliminating dependency on human annotations through automated domain prototyping, and (2) capturing the base model’s intrinsic feature space geometry for model-aware domain separation.
  • Phase 1 - Seed Generation: For each domain $k$, generate 5 seed samples $\mathcal{S}_{k}^{(0)}$ via zero-shot prompting with domain-specific description templates, establishing initial semantic anchors. The prompt details and ablation of choices on sample number are provided in Appendix B.7.
  • Phase 2 - Diversity Augmentation: Iteratively expand $S_{k}^{(t)}$ through context-aware generation, conditioned on a sliding window buffer with 3 random anchors sampled from the $(t-1)$ iteration $S_{k}^{(t-1)}$. The generated sample $\mathbf{x^{\prime}}$ will be admitted during the iteration via geometric rejection sampling:
    $\max_{\mathbf{x}\in S_{k}^{(t-1)}}\cos(\mathcal{M}_{\text{ebd}}(\mathbf{x^{\prime}}),\mathcal{M}_{\text{ebd}}(\mathbf{x}))<\tau=0.85,$ (13)
    where $\mathcal{M}_{\text{ebd}}(\cdot)$ indicates the embedding layer outputs of the given LLM. This process terminates when $|S_{k}|=30$. More analysis and ablation on these choices are provided in the Appendix. In Appendix D.3, we found that the adopted number of samples sufficiently leads to a stable convergence and larger data quantity does not impact the final centroids. In Appendix D.4, we found that this synthetic data has the ability to produce domain-representative data with clear distinction. Interestingly, the generated content consistently exhibits the greatest divergence between common sense, reasoning, and coding domains across model architectures and parameters.
The domain centroid is then computed from the final augmented set $\mathcal{S}_{k}$ using the model’s knowledge prior:
$\mathcal{C}_{k}=\frac{1}{|\mathcal{S}_{k}|}\sum_{\mathbf{x}_{i}\in\mathcal{S}_{k}}\mathcal{M}_{\text{ebd}}(\mathbf{x}_{i}).$ (14)
This captures the LLM’s intrinsic feature space geometry while eliminating dependency on human annotations.
Domain-Aware Clustering
We then automatically construct pseudo-labels for the given data samples based on the previously synthesized centroids $\{C_{k}\}_{k=1}^{K}$. Take them as fixed prototypes, we perform constrained k-means clustering in the embedding space:
$\arg\min_{\{S_{k}\}}\sum_{k=1}^{K}\sum_{\tilde{x}\in S_{k}}\|\mathcal{M}_{\text{ebd}}(\tilde{x})-\mathcal{C}_{k}\|^{2}.$ (15)
This produces the seed dataset $\mathcal{D}_{\text{probe}}$ containing pseudo-labels $\{\tilde{y}_{i}\}_{i=1}$ where $\tilde{y}_{i}\in\{1,\ldots,K\}$, with model-induced and embedding-derived domain label assignments.
5.2 Training for Self-Rewarding Abilities
Stage 1: Domain Discrimination
The proposed DaaR then establishes model-aware domain discrimination abilities through a multi-layer perceptron (MLP) probe module, $\psi_{\text{dom}}$, attached to the layer-3 of the LLMs $\mathcal{M}_{3}(\tilde{x})$. The probe will be trained meanwhile all the parameters of the LLM are frozen. This achieves a preferable balance between effectiveness and cost, with detailed analysis regarding the choice of Layer-3 presented in Appendix D.5. Specifically, with pseudo-label $\tilde{y}$, we can compute domain probabilities as:
$p_{k}(\tilde{x})=\text{softmax}\left(\psi_{\text{dom}}(\mathcal{M}_{3}(\tilde{x}))\right),\quad\psi_{\text{dom}}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{K},$ (16)
where $\psi_{\text{dom}}$ is optimized via cross-entropy loss function:
$\mathcal{L}_{\text{dom}}=-\frac{1}{|\mathcal{D}_{\text{probe}}|}\sum_{(\tilde{x},\tilde{y})\in\mathcal{D}_{\text{probe}}}\sum_{k=1}^{K}\mathbb{I}_{[k=\tilde{y}]}\log p_{k}(\tilde{x}),$ (17)
where $\mathbb{I}_{[k=\tilde{y}]}$ denotes the indicator function. We employ single-sample batches with the AdamW optimizer to prevent gradient averaging across domains. Training consistently converges and achieves 92.7% validation accuracy on domain classification, as shown in Fig. 2 (a).
Figure 2: Training loss and validation process of the two stages of DaaR on Qwen2-7B, detailed results are in Appendix D.6.
Stage 2: Diversity Rewarding
Building on the stabilized domain probe module, we quantify sample-level diversity through predictive entropy:
$H(\tilde{x})=-\sum_{k=1}^{K}p_{k}(\tilde{x})\log p_{k}(\tilde{x}).$ (18)
To enable efficient reward computation during data selection, we then train another 5-layer MLP $\psi_{\text{div}}$ to directly estimate $H(\tilde{x})$ from $\mathcal{M}_{3}(\tilde{x})$:
$\hat{H}(\tilde{x})=\psi_{\text{div}}(\mathcal{M}_{3}(\tilde{x})),\quad\psi_{\text{div}}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{+}.$ (19)
This diversity probe module $\psi_{\text{div}}$ shares $\psi_{\text{dom}}$’s architecture up to its final layer (replaced with single-output regression head), trained using entropy-scaled MSE:
$\mathcal{L}_{\text{div}}=\frac{1}{|\mathcal{D}_{\text{probe}}|}\sum_{\tilde{x}\in\mathcal{D}_{\text{probe}}}\left(\hat{H}(\tilde{x})-H(\tilde{x})\right)^{2}.$ (20)
This module is also well-converged as shown in Fig. 2 (b).
Table 2: Evaluation results of Llama3.1-8B, Qwen2-7B, and Qwen2.5-7B across various downstream task benchmarks. DaaR demonstrates superiority in average performance (AVG) compared to other baselines.
Models Distribution Common Sense Reasoning Mathematics Coding Avg
NQ TriviaQA Hellaswag GSM8K MATH MBPP HumanEval
Llama3.1-8B Raw 14.13 65.90 74.62 54.80 7.90 5.00 28.66 35.86
Random 21.99 64.83 74.72 55.70 14.50 5.10 24.09 37.27
Instruction Len 15.34 63.60 73.73 54.00 15.40 3.60 30.80 36.64
Alpagasus (Chen et al., 2024d) 21.57 64.37 74.87 55.20 17.65 4.60 16.16 36.34
Instag-Best (Lu et al., 2024) 18.12 64.96 74.01 55.70 15.50 4.80 37.81 38.70
SuperFilter (Li et al., 2024a) 22.95 64.99 76.39 57.60 6.05 2.60 40.55 38.73
Deita-Best (Liu et al., 2024b) 15.58 64.97 74.21 55.00 13.05 4.60 34.46 37.41
DaaR (Ours) 20.08 64.55 74.88 54.8 15.30 4.70 37.50 38.83
Qwen2-7B Raw 8.03 59.58 73.00 78.00 5.70 5.00 60.98 41.47
Random 13.28 58.27 73.00 75.35 35.36 52.20 63.72 53.02
Instruction Len 8.62 58.44 72.86 73.30 27.05 53.10 63.72 51.01
Alpagasus (Chen et al., 2024d) 13.67 57.94 73.04 73.90 32.30 51.40 63.41 52.24
Instag-Best (Lu et al., 2024) 9.51 58.50 73.06 74.70 35.35 51.90 64.70 52.53
SuperFilter (Li et al., 2024a) 19.16 58.98 72.99 73.70 30.10 52.40 58.85 52.31
Deita-Best (Liu et al., 2024b) 16.41 57.80 72.70 76.10 29.05 52.40 64.63 52.73
DaaR (Ours) 16.88 57.58 73.03 75.40 38.1 52.00 64.94 53.99
Qwen2.5-7B Raw 8.84 58.14 72.75 78.20 9.10 7.40 78.05 44.64
Random 11.46 57.85 73.08 78.90 13.15 62.50 71.65 52.65
Instruction Len 11.34 58.01 72.79 78.00 15.80 62.30 68.12 52.34
Alpagasus (Chen et al., 2024d) 10.40 57.87 72.92 77.20 18.75 61.80 65.55 52.07
Instag-Best (Lu et al., 2024) 11.08 58.40 72.79 76.40 16.40 62.90 70.43 52.63
SuperFilter (Li et al., 2024a) 13.54 58.51 72.89 79.30 11.35 39.50 65.25 48.62
Deita-Best (Liu et al., 2024b) 10.50 58.17 73.14 74.60 16.60 62.00 72.26 52.47
DaaR (Ours) 15.83 58.65 72.48 80.20 16.70 64.20 68.29 53.76
Data Selection: After training the module $\psi_{\text{div}}$, we can use its output to select data samples. Building on the theoretical insights in Sec. 4.4, data points that are closer to other centroids and more dispersed within their own centroid are more beneficial for enhancing the comprehensive capabilities of the model. Therefore, we use the predicted entropy-diversity score as a reward, selecting the top 20% with the highest scores as the final data subset for fine-tuning.
5.3 Empirical Validation and Main Results
To validate the efficacy of DaaR, we conduct experiments comparing more SOTA methods on the data pools in Sec. 3.1 with critical modifications: all domain-specific labels are deliberately stripped. This constraint mimics more challenging real-world scenarios and precludes direct comparison with data mixture methods requiring domain label prior.
Baselines
We use the following data selection methods for comprehensive evaluation: (1) Random Selection: traditional random sampling; (2) Instruction Len: measuring instruction complexity by token count  (Cao et al., 2023); (3) Alpagasus (Chen et al., 2024d): using ChatGPT for direct quality scoring of instruction pairs; (4-5) Instag (Lu et al., 2024): semantic analysis approach with Instag-C (complexity scoring via tag quantity) and Instag-D (diversity measurement through tag set expansion); (6) SuperFilter (Li et al., 2024a): response-loss-based complexity estimation using compact models; (7-9) Deita (Liu et al., 2024b): model-driven evaluation with Deita-C (complexity scoring), Deita-Q (quality scoring), and Deita-D (diversity-aware selection). Detailed configurations for these baselines are in Appendix B.4.
5.3.1 Overall Performance
The experimental results are presented in Table 2, where Instag-Best and Deita-Best represent the optimal variants from their respective method families. Our experiments clearly demonstrate the effectiveness of DaaR across three major language models and seven challenging benchmarks. A detailed analysis is provided below.
High-Difficulty Scenario: The task of balanced capability enhancement proves particularly challenging for existing methods. While some baselines achieve strong performance on specific tasks (e.g., SuperFilter’s 40.55 on HumanEval for Llama3.1), they suffer from catastrophic performance drops in other domains (e.g., SuperFilter’s 6.05 on MATH). Only three baseline methods perform better than Random selection, with notably severe degradation applied in Qwen-series models, all of which fall below the Random performance. We hypothesize this stems from over-specialization – excessive focus on narrow capability peaks at the expense of broad competence (visualized in Appendix B.5). Our method’s robustness stems from preventing extreme distribution shifts through diversity constraints.
Universal Superiority: DaaR establishes new SOTA averages across all models, surpassing the best baselines by +0.14 (Llama3.1), +0.97 (Qwen2), and +1.11 (Qwen2.5). The proposed method uniquely achieves dual optimization in critical capabilities: Mathematical Reasoning: Scores 38.1 MATH (Qwen2) and 16.70 MATH (Qwen2.5), with 7.4% and 27.0% higher than respective random baselines. Coding Proficiency: Maintains 64.94 HumanEval (Qwen2) and 64.20 MBPP (Qwen2.5) accuracy with <1% degradation from peak performance. This demonstrates DaaR’s ability to enhance challenging STEM capabilities while preserving core competencies, a critical advancement over specialized but unstable baselines.
Cost-Efficiency and Flexibility: Compared to baseline methods requiring GPT-based evaluators (Alpagasus, Instag) or full LLaMA-7B inference (Deita), DaaR achieves superior efficiency through data-model co-optimizations. Our method makes the LLM ability of self-rewarding from dedicated data synthesis, and operates on frozen embeddings from layer 3 (vs. full 32-layer inference in comparable methods), reducing computational overhead while maintaining capability integrity. The lightweight 5-layer MLP probe module requires only 9GB GPU memory during training (vs. 18$\sim$24GB for LLM-based evaluators) and adds merely 76M parameters. This plug-and-play feature enables seamless integration of DaaR with existing LLMs without additional dependency management.
6 Conclusion and Future Work
In this paper, we propose a new approach to fine-tuning LLMs with data that lacks clear domain labels, using diversity as a guiding principle. Our method allows LLMs to automatically select and benefit from diverse datasets. By measuring semantic diversity with entropy, we employ a self-reward mechanism built upon the given LLM, identifying data that best fits the model’s natural tendencies in terms of its underlying knowledge distribution. Our experiments with various SOTA LLMs show notable superiority of the method over SOTA methods, highlighting the potential of data diversity to enhance model overall performance.
This research demonstrates feasibility of reward modeling on LLM data diversity and deepens the understanding of its utilities, especially when domain labels are missing or uncertain. Future work could include developing methods to efficiently adjust diversity measures and learning algorithm, thus supporting LLMs in self-evolving towards artificial general intelligence in dynamic environments.
Appendix
Our appendix is organized as follows:
  • Sec. A: The symbols used in this paper.
  • Sec. B: The details about experiments and DaaR implementation, including the construction of data pools (Sec. B.1), the benchmarks (Sec. B.2), the training & evaluation implementation (Sec. B.3), the implementation of baselines (Sec. B.4), the visualization of baseline’s selected data samples (Sec. B.5), the implementation of the synthetic data generation (Sec. B.6), the implementation of seed generation & clustering (Sec. B.7), and the prompts for the generation stage (Sec. B.8).
  • Sec C: The full proof for theoretical results are presented in Section 4.
  • Sec D: The more experimental results, including the visualization results of embeddings on Llama3.1-8B & Qwen2.5-7B (Sec D.1), the visualization results of inter- & intra-diversity distributions (Sec D.2), the similarity analysis in generated centroid quantities (Sec. D.3), the similarity analysis in generated centroid domains (Sec. D.4), the DaaR layer selection protocol (Sec. D.5), the comprehensive training dynamics (Sec. D.6), the validation of inter-diversity and intra-diversity (Sec. D.7) and the complete results of DaaR with baselines on all LLMs (Sec. D.8).
Appendix A Notation
For ease of reading and reference, we present all mathematical symbols used in this paper in Table 3.
Table 3: Symbol Notation
Symbol Description Symbol Description
$\mathcal{D}$ Composite dataset $\mathcal{D}_{k}$ Distinct domain dataset
$x_{i}^{(k)}$ Raw $i$-th data sample in domain-$k$ $\mathbf{x}_{i}^{(k)}$ Embedded $i$-th data sample in domain-$k$
$N_{k}$ Number of $\mathcal{D}_{k}$ data samples $\mathcal{C}_{k}$ Domain centroid of $k$
$\phi_{\text{inter}}$ Sample-level cross-domain similarity $\Phi_{\text{inter}}$ Global inter-diversity
$\phi_{\text{intra}}$ Sample similarity to its domain centroid $\Phi_{\text{intra}}^{(k)}$ Domain-level variance of domain-$k$
$\mathcal{K}$ Set of labeled domains $h_{k}$ Independent model of domain $k$
$l$ Loss function $\mathcal{L}_{\mathcal{D}_{k}}$ Empirical risk
$\tilde{\mathcal{D}}_{m}$ Latent distribution of foundational capabilities $\pi_{km}$ Mixture weights
$h_{\theta_{m}}$ Component model $\theta_{m}^{*}$ Optimal predictor
$\mathcal{S}_{k}^{(0)}$ Generated centroids’ seed sample $\mathcal{S}_{k}^{(t)}$ Generated centroid sample at $t$-th round
$\mathcal{M}_{\text{ebd}}$ Embedding layer of LLM $\mathcal{M}_{3}$ Layer-3 of LLM
$\mathcal{D}_{\text{probe}}$ DaaR training set $\tau$ A similarity threshold
$\tilde{x}_{i}$ Data sample in $\mathcal{D}_{\text{probe}}$ $\tilde{y}_{i}$ Pseudo-label of $\tilde{x}_{i}$
$\psi_{\text{dom}}$ A MLP-based domain classifier $\psi_{\text{div}}$ A MLP-based entropy predictor
$p_{k}(\tilde{x})$ Domain predicted probability by $\psi_{\text{dom}}$ $\hat{H}(\mathbf{x})$ Predictive entropy by $\psi_{\text{dom}}$
$\mathcal{L}_{\text{dom}}$ Cross-entropy loss of $\psi_{\text{dom}}$ $\mathcal{L}_{\text{div}}$ MSE loss of $\psi_{\text{dom}}$
Appendix B Experiments and Method Details
B.1 Construction of Data Pool
To investigate data selection from large data pools and its impact on the mixture of downstream tasks, we construct a data pool with distinct properties to mimic practical settings. We select the following datasets to evaluate specific abilities:
  • Common Sense: Dolly-15K (Conover et al., 2023) with 15,011 samples, an open source dataset of instruction-following records generated by thousands of Databricks employees in several of the behavioral categories.
  • Reasoning: Cot-en (Chung et al., 2024) with 74,771 samples, is created by formatting and combining nine CoT datasets released by FLAN.
  • Mathematics: Math-Instruct (Yue et al., 2024) with 262,039 samples, is compiled from 13 math rationale datasets, six of which are newly curated by this work. It uniquely ensures extensive coverage of diverse mathematical fields.
  • Coding: Code-Alpaca (Chaudhary, 2023) with 20,016 samples, is constructed from real-world code examples, providing a rich set of tasks designed to guide models in generating accurate and functional code.
Each dataset was initially filtered and randomly reduced to 10,000 entries, resulting in a combined data pool of 40,000 entries. Specifically, for the Math-Instruct dataset, due to its inclusion of CoT and certain coding capabilities, we extract a highly mathematics-related subset and use regular expressions to filter out the coding-related content (including ’program’, ’python’, ’def’, ’import’, ’print’, ’return’), ensuring it remains within the domain of mathematics.
B.2 Benchmarks
To evaluate the models’ true capabilities and performance across different domains, we follow the approach of major open-source LLMs (e.g., the Llama3 series (Dubey et al., 2024) and Qwen2 series (Yang et al., 2024)) and select the following widely used evaluation sets. All evaluations were conducted on the OpenCompass platform [^2] https://opencompass.org.cn/.
  • Common Sense: NQ (Kwiatkowski et al., 2019) and TriviaQA (Joshi et al., 2017), which cover factual knowledge-based questions of varying difficulty.
  • Reasoning: HellaSwag (Zellers et al., 2019), which effectively evaluates the model’s comprehensive reasoning ability.
  • Mathematics: GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) benchmarks, which encompass problems ranging from elementary to competition-level difficulty.
  • Coding: MBPP (Austin et al., 2021) and HumanEval (Chen et al., 2021), which include evaluations of basic coding abilities in Python. We use the average of various metrics to demonstrate the models’ overall performance across different domains.
Considering the numerous evaluation tasks, utilizing the complete evaluation set would result in significant time expenditure. To accelerate the evaluation process while maintaining fairness and accuracy, we randomly tailor the original evaluation sets into evaluation subsets, as detailed in Table 4. All experiments were conducted using this consistent setup to ensure the fairness of the experiments.
Table 4: Number of samples in various evaluation benchmarks’ datasets.
Number of Samples Common Sense Reasoning Mathematics Coding
NQ Triviaqa Hellaswag GSM8K MATH MBPP HumanEval
Original 3,610 8,837 10,042 1,319 5,000 974 164
Utilized 3,610 5,000 10,042 500 1,000 500 164
B.3 Training and Evaluation Details
Platform
We implement our approaches using PyTorch (Paszke et al., 2019) v2.4.1, coupled with PEFT v0.12.0 and the Transformers library (Wolf et al., 2020) v4.45.2. Experiments are conducted on a computing platform equipped with four NVIDIA A100 GPUs (40GB), with pre-trained LLMs loaded as 16-bit floating-point numbers. The specific data-model development processes are completed in Data-Juicer Sandbox (Chen et al., 2024c), via integration with the ms-swift (Zhao et al., 2024b) training repository, and the OpenCompass (Contributors, 2023) evaluation repository.
Training Details
In our experimental setup, we employ Low-Rank Adaptation (LoRA) (Hu et al., 2021) adapters for the fine-tuning process, utilizing a LoRA-rank of 8 and a LoRA-alpha of 16. The learning rate was consistently maintained at $5\times 10^{-5}$ across all experiments to ensure uniformity in training dynamics. We utilize a batch size of 4 and set the maximum sequence length to 2048 tokens to accommodate the model’s capacity. To optimize the training process, a warmup ratio of 0.05 was applied and a validation ratio of 0.03 was used. The training was conducted over a single epoch, balancing computational efficiency with the need for effective model adaptation. Following some effective instruction-tuning work (Zhou et al., 2024; Lu et al., 2024), we set the size of our subset to 8,000 entries, which constitutes 20% of the data pool.
Evaluation Details
Following the guidelines provided by OpenCompass (Contributors, 2023), we adhered to the default settings for our evaluation process. We select the hf-type as base and utilized a batch size of 16 to ensure efficient processing. For most tasks, we employ the gen mode, while for the Hellaswag task, we opt for the ppl mode to better assess perplexity.
B.4 Implementation of Baselines
We select the following representative methods and works related to data selection as our baselines to evaluate their performance in the context of a mixture of downstream tasks.
Random Selection(Rand)
A random selection of 8,000 data samples from the data pool was made, which to some extent reflects the distribution characteristics of the original data pool.
Instruction Length (IL)
The length of the instruction can be considered a measure of input complexity. It is widely believed (Cao et al., 2023; Zhao et al., 2023) that more complex data is beneficial for enhancing model capabilities. Therefore, we select a subset of 8,000 entries with the maximum number of words (based on spaces) in the concatenation of Instruction and Input as part of the filtering process.
AlpaGasus (Chen et al., 2024d)
Using prompts to directly score and annotate the quality of the data leverages the powerful cognitive abilities of LLMs for evaluation and selection. Based on the original work, we use the GPT-3.5-Turbo API to score the data with the following prompt:
[…]
We would like to request your feedback on the performance of AI assistant in response to the instruction and the given input displayed following.
Instruction: [Instruction]
Input: [Input]
Response: [Response]
User Prompt:
Please rate according to the accuracy of the response to the instruction and the input. Each assistant receives a score on a scale of 0 to 5, where a higher score indicates a higher level of accuracy. The generated scores can be precise to decimal points, not just in increments of 0.5. Please first output a single line containing the value indicating the scores. In the subsequent line, please provide a comprehensive explanation of your evaluation, avoiding any potential bias.
The distribution of direct ratings for the 40,000 data pool is shown in Fig. 3. Although the original paper’s prompts were strictly followed and efforts were made to minimize potential bias, most scores still clustered around 5. Since we require a uniform selection of 8,000 data samples, we randomly select the subset with a rating of 5 to serve as the baseline data.
Figure 3: The distribution of the score by Alpagasus.
Instag (Lu et al., 2024)
first utilizes ChatGPT to tag the samples based on semantics and intentions, then trains a LLaMA-based tagger on the ChatGPT tags to tag data. They use the number of tags as a proxy for complexity. We directly use ChatGPT-3.5-Turbo as a tagger to achieve better performance. Following the original paper, the prompt is as follows.
[…]
You are a tagging system that provides useful tags for instruction intentions to distinguish instructions for a helpful AI assistant. Below is an instruction:
[begin]
{Instruction + Input}
[end]
User Prompt:
Please provide coarse-grained tags, such as ”Spelling and Grammar Check” and ”Cosplay”, to identify main intentions of above instruction. Your answer should be a list including titles of tags and a brief explanation of each tag. Your response have to strictly follow this JSON format: [”tag”: str, ”explanation”: str]. Please response in English.
A total of 19,585 tags were assigned across 40,000 data samples, with the distribution shown below. Subsequently, based on the procedures outlined in the original text, tags were deduplicated as follows: 1) filter out long-tail tags that appear fewer than $\alpha$ times in the entire annotated dataset, and 2) transform all tags to lowercase to mitigate the influence of capitalization.
Figure 4: The distribution of the tags by Instag.
  • Instag Complexity (Instag-C): To reduce redundancy, we apply a threshold of $\alpha=5$, resulting in a set of 1,948 valid tags. Following the definition of Complexity, we select the top 8,000 entries with the highest tag counts. Specifically, there are 8,211 entries with more than three tags, so we include all records with more than four tags and randomly supplement from those with exactly four tags until reaching a total of 8,000 entries.
  • Instag Diversity (Instag-D): For Diversity, we use $\alpha=1$ to reduce redundancy. The algorithm employed involves sorting all data in descending order based on the number of tags, and prioritizing records with more tags. To ensure dataset diversity, a tag is only added to the final dataset if it increases the overall size of the tag set. This approach captures a broader range of unique tags, thereby enhancing the diversity and representativeness of the dataset.
SuperFilter (Li et al., 2024a)
introduces a method that astonishingly utilizes a small GPT2 model to successfully filter out the high-quality subset from the existing GPT4-generated instruction tuning dataset. The core concept is the instruction-following difficulty (IFD) score. By meticulously following this methodology, we utilize GPT-2 to select 8,000 data samples.
Deita (Liu et al., 2024b)
is an open-sourced project designed to facilitate Automatic Data Selection for instruction tuning in LLMs. It delves into the relationships between data complexity, quality, and diversity, and develops a series of methodologies. Specifically, we utilize it as the following three baselines:
  • Deita Complexity (Deita-C): Enhances instruction complexity by evolving examples with techniques like adding constraints. ChatGPT ranks these evolved samples for complexity, and the scores train a LLaMA-1-7B model to predict complexity in new instructions, refining complexity assessment. We utilize its LLaMA-1-7B-based complexity-scorer to evaluate the data pool and select the top 20% of them to reach 8,000 entries.
  • Deita Quality (Deita-Q): Enhances response quality by prompting ChatGPT to iteratively improve helpfulness, relevance, depth, creativity, and detail. After five iterations, ChatGPT ranks and scores the responses, providing nuanced quality distinctions. These scores train an LLaMA-1 7B model to predict quality scores for new instruction-response pairs. We utilize its LLaMA-1-7B-based quality scorer to evaluate the data pool and select the top 20% of them to reach 8,000 entries.
  • Deita Deita (Deita-D): Data-Efficient Instruction Tuning for Alignment, it selects data by combining complexity and quality into an evol score, prioritizing higher scores. It uses the REPR FILTER to iteratively add samples, ensuring diversity and avoiding redundancy, resulting in a balanced dataset of complexity, quality, and diversity. Following its methodology, we gain 8,000 entries as a baseline.
B.5 Visualization of Baseline’s Selected Data
We project all baseline-selected data samples through Llama3.1-8B’s, Qwen2-7B’s and Qwen2.5-7B’s embedding layer and visualize their distributions via t-SNE dimensionality reduction in Fig. 5, Fig. 6 and Fig. 7 respectively. This visualization provides intuitive insights into how different data selection strategies handle domain-mixed data. Three key observations emerge:
  • 1
    Certain methods (notably Instruction Length, Deita-C, and Deita-Q baselines) exhibit significant distributional bias against coding-domain samples, correlating with their poor HumanEval performance in Table 10.
  • 2
    The visualization reveals inherent challenges in maintaining original data distributions when handling domain-undetermined samples.
  • 3
    Despite observable distribution shifts, not all methods show proportional performance degradation, aligning with our theoretical analysis in Section 4 about the link between latent distribution and LLM’s fundamental capabilities.
Figure 5: t-SNE visualization of data samples selected by different baselines using Llama3.1-8B embeddings. While original labels were removed during training, we preserve them in the legend for interpretability.
Figure 6: t-SNE visualization of data samples selected by different baselines using Qwen2-7B embeddings. While original labels were removed during training, we preserve them in the legend for interpretability.
Figure 7: t-SNE visualization of data samples selected by different baselines using Qwen2.5-7B embeddings. While original labels were removed during training, we preserve them in the legend for interpretability.
B.6 Implementation of the Synthetic Data Generation
For the generation of synthetic data, we utilize the model.generate function from the PyTorch library. Key parameters included setting max-new-tokens to 2048 to control the length of the output, enabling do-sample to True to allow sampling, and configuring top-p to 0.95 and temperature to 0.9 to ensure diversity in the generated content. During the content extraction phase, we employ regular expressions to efficiently extract and structure the desired information.
B.7 Implementation of Seed Generation & Clustering
Due to the difficulty of pre-trained LLMs in providing accurate and coherent responses, we utilize their corresponding Instruct versions for centroid data generation. Specifically, we apply Llama3.1-8B-Instruct for Llama3.1-8B, Qwen2-7B-Instruct for Qwen2-7B, and Qwen2.5-7B-Instruct for Qwen2.5-7B. The domain description and the zero-shot prompt used in the Seed Generation phase are as follows. When clustering reward training data based on the seed, we select 5,000 entries that match the distribution of the fine-tuning data but do not overlap with it. This practice has proven that even a small amount of data can effectively train a reward probe module of DaaR.
[…]
Reasoning: Reasoning involves the ability to think logically about a situation or problem, to draw conclusions from available information, and to apply knowledge in new situations.
Mathematics: Mathematical skills include the ability to perform calculations, understand mathematical concepts, solve hard and professional math problems, and apply mathematical reasoning.
Coding: Design and generate specific code programs, or apply algorithms and data structures, with code generation in the Output.
[…]
Generate 5 different instruction pairs related to this field with various lengths. Maintain the index format: Instruction [1 … 5].
The response should include three parts:
1. Instruction: A clear command or question that can be understood by the assistant.
2. Input: Any information provided to help it understand the instruction. If there is no need to generate, just keep it empty.
3. Output: The expected answer or action.
Keep the generated content focused on {selected_domain}. And do not involve {unselected_domains} related knowledge.
B.8 Prompts for the Generation Stage
[…]
Generate only an instruction pair related to this field. The response should include three parts:
Instruction: A clear command or question that can be understood by the assistant.
Input: Any information provided to help it understand the instruction. If there is no need to generate, just keep empty.
Output: The expected answer or action.
Keep the generated content focused on {selected_domain}. Do not involve {unselected_domain} related knowledge.
Note that you should generate content strongly unrelated and different to these examples to ensure diversity in the generated output:
Counterexample: {}
The format of the generated content should be: Instruction: [], Input: [], Output: [].
Appendix C Proof of Section 4
C.1 Proof of Proposition 4.3
Lemma C.1 (Cross-Domain Risk Decomposition).
Under Assumption 4.1, the average cross-domain risk becomes:
$\mathbb{E}_{k\sim\mathcal{K}}\mathcal{L}_{\mathcal{D}_{k}}\left(\textstyle\sum_{m=1}^{M}\pi_{km}h_{\theta_{m}}\right)=\sum_{m=1}^{M}\lambda_{m}\mathbb{E}_{\tilde{\mathcal{D}}_{m}}\left[l\left(\textstyle\sum_{m^{\prime}=1}^{M}\pi_{km^{\prime}}h_{\theta_{m^{\prime}}}(x),y\right)\right],$ (21)
where $\lambda_{m}=\mathbb{E}_{k}[\pi_{km}]$ represents global activation of capability $m$.
Proof.
Under the finite variance condition $\mathbb{E}_{\tilde{\mathcal{D}}_{m}}[\|l\|_{2}^{2}]<\infty$, Fubini’s theorem holds, and results in:
$\mathbb{E}_{k}\mathcal{L}_{\mathcal{D}_{k}}$ $=\mathbb{E}_{k}\mathbb{E}_{(x,y)\sim\mathcal{D}_{k}}\left[l\left(\textstyle\sum_{m}\pi_{km}h_{\theta_{m}}(x),y\right)\right]$
$=\mathbb{E}_{k}\mathbb{E}_{m\sim\pi_{k}}\mathbb{E}_{\tilde{\mathcal{D}}_{m}}\left[l\left(\textstyle\sum_{m^{\prime}}\pi_{km^{\prime}}h_{\theta_{m^{\prime}}}(x),y\right)\right]$
$=\sum_{m=1}^{M}\lambda_{m}\mathbb{E}_{\tilde{\mathcal{D}}_{m}}\left[l\left(\textstyle\sum_{m^{\prime}}\pi_{km^{\prime}}h_{\theta_{m^{\prime}}}(x),y\right)\right]$ (22)
Theorem C.2 (Global Optimality Condition).
Under Assumptions 4.1-4.2, the optimal predictors $\{h_{\theta_{m}^{*}}\}$ minimize the cross-domain risk when:
$\forall m,\ \theta_{m}^{*}=\mathop{\mathrm{argmin}}_{\theta}\mathbb{E}_{\tilde{\mathcal{D}}_{m}}[l(h_{\theta}(x),y)].$ (23)
Proof.
Taking derivatives of Lemma C.1’s right-hand side:
$\frac{\partial}{\partial\theta_{m}}\sum_{m^{\prime}=1}^{M}\lambda_{m^{\prime}}\mathbb{E}_{\tilde{\mathcal{D}}_{m^{\prime}}}[\cdot]=\lambda_{m}\mathbb{E}_{\tilde{\mathcal{D}}_{m}}\left[\frac{\partial l}{\partial h}\frac{\partial h_{\theta_{m}}}{\partial\theta_{m}}\right]+\sum_{m^{\prime}\neq m}\lambda_{m^{\prime}}\mathbb{E}_{\tilde{\mathcal{D}}_{m^{\prime}}}\left[\frac{\partial l}{\partial h}\frac{\partial h_{\theta_{m}}}{\partial\theta_{m}}\right].$
By strong convexity and optimality of $\theta_{m}^{*}$, all terms vanish when $\theta_{m}=\theta_{m}^{*}$. ∎
C.2 Proof of Proposition 4.4
Proof of Inter-Diversity
To prove the inter-diversity decomposition as given in Proposition 4.4, we begin by considering the inter-diversity metric $\Phi_{\text{inter}}$, defined as the expected pairwise distance between domain centroids from Eq (3):
$\Phi_{\text{inter}}=\mathbb{E}_{k\neq l}\left[\|\mathcal{C}_{k}-\mathcal{C}_{l}\|_{2}\right].$ (24)
Substituting the expression for each domain centroid (Eq (10)) $\mathcal{C}_{k}=\sum_{m=1}^{M}\pi_{km}\mathcal{C}_{m}$, we have:
$\|\mathcal{C}_{k}-\mathcal{C}_{l}\|_{2}=\left\|\sum_{m=1}^{M}\pi_{km}\mathcal{C}_{m}-\sum_{n=1}^{M}\pi_{ln}\mathcal{C}_{n}\right\|_{2}.$ (25)
Applying the triangle inequality and expanding, the distance becomes:
$=\left\|\sum_{m=1}^{M}\sum_{n=1}^{M}(\pi_{km}-\pi_{ln})\mathcal{C}_{m}+\pi_{ln}\mathcal{C}_{n}\right\|_{2}.$ (26)
Taking the expectation over all domain pairs $k\neq l$ and using linearity, we have:
$\Phi_{\text{inter}}=\mathbb{E}_{k\neq l}\left[\sum_{m=1}^{M}\sum_{n=1}^{M}(\pi_{km}\pi_{ln})\|\mathcal{C}_{m}-\mathcal{C}_{n}\|_{2}\right].$ (27)
Recognizing that $\lambda_{m}=\mathbb{E}_{k}[\pi_{km}]$ is the expected activation of capability $m$, we can simplify further:
$=\sum_{m=1}^{M}\sum_{n=1}^{M}\lambda_{m}\lambda_{n}\|\mathcal{C}_{m}-\mathcal{C}_{n}\|_{2}.$ (28)
This completes the proof, showing that the inter-diversity metric is determined by the expected pairwise distances between latent capability centroids, weighted by their activation probabilities.
Proof of Intra-Diversity
The intra-diversity metric for domain $k$, denoted as $\Phi_{\text{intra}}^{(k)}$, measures the variance of data samples within the domain relative to its centroid $\mathcal{C}_{k}$. We aim to show that:
$\Phi_{\text{intra}}^{(k)}\leq\sum_{m=1}^{M}\pi_{km}\|\mathcal{C}_{m}-\mathcal{C}_{k}\|_{2}^{2}+\mathbb{E}_{m\sim\pi_{k}}[\text{Var}(\tilde{\mathcal{D}}_{m})].$ (29)
Proof.
To prove this bound, we start by expressing the intra-diversity metric $\Phi_{\text{intra}}^{(k)}$ as the expected squared distance from each sample $\mathbf{x}_{i}^{(k)}$ to the domain centroid $\mathcal{C}_{k}$ from Eq (5):
$\Phi_{\text{intra}}^{(k)}=\frac{1}{N_{k}}\sum_{i=1}^{N_{k}}\|\mathbf{x}_{i}^{(k)}-\mathcal{C}_{k}\|_{2}^{2}.$ (30)
Using the definition of the domain centroid (Eq (10)) $\mathcal{C}_{k}=\sum_{m=1}^{M}\pi_{km}\mathcal{C}_{m}$ in , we can decompose the squared distance for each sample:
$\|\mathbf{x}_{i}^{(k)}-\mathcal{C}_{k}\|_{2}^{2}=\|\mathbf{x}_{i}^{(k)}-\sum_{m=1}^{M}\pi_{km}\mathcal{C}_{m}\|_{2}^{2}.$ (31)
Applying the variance decomposition, we have:
$\|\mathbf{x}_{i}^{(k)}-\mathcal{C}_{k}\|_{2}^{2}=\|\mathbf{x}_{i}^{(k)}-\sum_{m=1}^{M}\pi_{km}\mathcal{C}_{m}\|_{2}^{2}=\sum_{m=1}^{M}\pi_{km}\|\mathbf{x}_{i}^{(k)}-\mathcal{C}_{m}\|_{2}^{2}+\sum_{m=1}^{M}\pi_{km}\|\mathcal{C}_{m}-\mathcal{C}_{k}\|_{2}^{2},$ (32)
where the first term represents the variance within each latent capability, and the second term accounts for the distance between latent capability centroids and the domain centroid.
By taking the expectation over the samples in domain $k$, we get:
$\Phi_{\text{intra}}^{(k)}\leq\sum_{m=1}^{M}\pi_{km}\|\mathcal{C}_{m}-\mathcal{C}_{k}\|_{2}^{2}+\mathbb{E}_{m\sim\pi_{k}}[\text{Var}(\tilde{\mathcal{D}}_{m})],$ (33)
where $\text{Var}(\tilde{\mathcal{D}}_{m})$ represents the variance within the $m$-th latent capability. This completes the proof of the intra-diversity bound. ∎
Appendix D More Experimental Results
D.1 Visualization of Embeddings on Llama3.1-8B & Qwen2.5-7B
We process concatenated data samples with “instruction”+“input”+“output” pairs through each LLM’s embedding layer, computing mean token embeddings for visualization. Beyond Qwen2-7B results in Fig. 1, we present the t-SNE visualization of data samples from Llama3.1-8 and Qwen2.5-7B with different distributions in Fig. 8 and Fig. 9. It is evident that, despite differences in architecture or model, the method of filtering data through Inter-Diversity and Intra-Diversity is effective. Notably, although the embedding distributions of different models are not identical, they exhibit similar behavior, indicating that the representations learned by the embedding layers are comparable.
Figure 8: The t-SNE visualization of embeddings for data samples with different distributions on Llama3.1-8B. (a) The data pool of all 40,000 samples, (b) Randomly selected subset, (c) Distribution of data farthest from other domain centroids based on Inter-Diversity, (d) Distribution of data closest to other domain centroids based on Inter-Diversity, (e) Distribution of data closest to its own domain centroid based on Inter-Diversity, (f) Distribution of data farthest from its own domain centroid based on Inter-Diversity.
Figure 9: The t-SNE visualization of embeddings for data samples with different distributions on Qwen2.5-7B. (a) The data pool of all 40,000 samples, (b) Randomly selected subset, (c) Distribution of data farthest from other domain centroids based on Inter-Diversity, (d) Distribution of data closest to other domain centroids based on Inter-Diversity, (e) Distribution of data closest to its own domain centroid based on Inter-Diversity, (f) Distribution of data farthest from its own domain centroid based on Inter-Diversity.
D.2 The Visualization of Inter- & Intra-Diversity Distribution Data
Using the Qwen2-7B model as an example, we construct data based on the Inter-Diversity and Intra-Diversity distribution methods, selecting a batch of data every 20%. The visualization process is shown in Fig. 10 and Fig. 11. As seen in the figures, the data gradually transitions from domain-aware diverse to domain-aware closed, indicating that our data construction method effectively controls the distribution of different data. However, it is important to note that t-SNE dimensionality reduction is used only for visualization purposes; during the reduction process, positional information is lost, and it cannot fully capture the true high-dimensional relationships between the data.
Figure 10: Data visualization (t-SNE) based on different Inter-Diversity distributions on Qwen2-7B.
Figure 11: Data visualization (t-SNE) based on different Intra-Diversity distributions on Qwen2-7B.
D.3 Similarity in Generated Centroid Quantities
To assess the impact of the amount of generated data on the accuracy of Domain Semantic Centroids, we select quantities of 10 and 30 samples for generation. Subsequently, we calculate the semantic cosine similarity between these two sets, with results presented in Table 5 (using Qwen2-7B and Llama3.1-8B as examples). The observed differences between varying data quantities were not significant, indicating that the generalization of the generated data is sufficient and further increasing the data quantity does not significantly enhance the accuracy of the centroids.
Table 5: Semantic cosine similarity between generated samples of size 10 and 30 across different domains.
Similarity of 10 & 30 Common Sense Reasoning Mathematics Coding
Qwen2-7B 0.9686 0.9866 0.9919 0.9881
Llama3.1-8B 0.9851 0.9795 0.9685 0.9895
D.4 Similarity in Generated Centroid Domains
To investigate whether there is a distinct separation between different domains generated, we compute the cosine similarity of semantic centroids across different domains. The experimental results are shown in Fig 12. It is evident that, compared to variations within the same domain at different data quantities, there are significant differences between different domains. This validates that this method of generating synthetic data can produce domain-representative data with clear distinction.
Notably, despite variations in model architecture and parameter count, the generated content consistently exhibits the greatest divergence between common sense, reasoning, and coding domains. Specifically, the discrepancy between common sense and coding, as well as reasoning and coding, is markedly pronounced. Conversely, the semantic difference between common sense and reasoning is relatively smaller. This pattern suggests that the models, although differing in complexity and size, exhibit varying sensitivity to different data types, highlighting their nuanced capability to distinguish between certain domain-specific characteristics while maintaining subtle distinctions in others.
Figure 12: Semantic cosine similarity across different domains for generated samples of size 10.
D.5 DaaR Layer Selection Protocol
To determine an appropriate layer for embedding DaaR into a large language model (LLM) that balances performance and computational cost, we first evaluate the capacity of different layers for domain awareness. Using Qwen2-7B as an example, we conduct t-SNE visualizations of Layer-5, Layer-10, Layer-15, Layer-20, Layer-25, and the last hidden layer, based on the average token vectors, as shown in Fig. 13.
Figure 13: t-SNE visualization of mean token embeddings across Layer-5, Layer-10, Layer-15, Layer-20, Layer-25, and the last hidden layer of Qwen2-7B.
Our observations indicate that even at the early Layer-5, the model begins to exhibit classification capabilities across different datasets. This prompted us to focus on the initial five layers to enhance learning outcomes. Considering that data clustering is inherently influenced by the embedding layer, a natural approach was to directly connect to the embedding layer for training. However, the training results showed persistently high training loss and a validation accuracy of only about 0.6, indicating suboptimal learning performance. Furthermore, related studies indicate that LLMs can learn semantic understanding and perception in the early layers (Wei et al., 2022), consequently, we select Layer-3 for DaaR.
Figure 14: Training loss and validation process of Embedding Layer on Qwen2-7B.
D.6 Comprehensive Training Dynamics
We present the training loss and validation results for two distinct models, Llama3.1-8B and Qwen2.5-7B, in Fig 15. It can be observed that across different models, the training process of DaaR method consistently ensures gradual convergence, achieving high domain predictability and calibrated entropy fitting.
Figure 15: Training loss and validation process of the two training stages of DaaR on Llama3.1-8B and Qwen2.5-7B, showing the model gradually converging.
D.7 Complete Validation Results of Inter-Diversity and Intra-Diversity
We present the complete experimental results of Llama3.1-8B, Qwen2-7B, and Qwen2.5-7B in the validation experiments in Tables 6, Tables 7, and Tables 8, respectively. It can be observed that for any complete dataset, the conclusions from Section 3.3 remain valid, specifically that the peak distribution of results is uneven, with significant differences among them.
Table 6: Validation Results of Inter-Diversity and Intra-Diversity on Llama3.1-8B across various benchmarks.
Llama3.1-8B Common Sense Reasoning Mathematics Coding Avg
NQ TriviaQA Hellaswag GSM8K MATH MBPP HumanEval
Raw 14.13 65.90 74.62 54.80 7.90 5.00 28.66 35.86
Rand 21.99 64.83 74.72 55.70 14.50 5.10 24.09 37.27
Inter-Diversity (0-20) 19.28 65.79 74.44 54.90 6.50 4.30 35.06 37.18
Inter-Diversity (20-40) 21.91 65.48 74.54 52.50 16.65 5.70 26.53 37.62
Inter-Diversity (40-60) 23.70 65.14 74.86 56.40 17.15 5.00 24.40 38.09
Inter-Diversity (60-80) 22.42 64.52 74.76 55.10 14.60 7.40 31.10 38.56
Inter-Diversity (80-100) 23.76 64.43 75.20 56.40 15.05 4.50 33.54 38.98
Intra-Diversity (0-20) 22.08 65.08 75.00 54.70 16.20 4.40 33.54 38.71
Intra-Diversity (20-40) 22.41 64.44 74.66 52.60 15.30 4.20 27.44 37.29
Intra-Diversity (40-60) 22.12 64.74 74.87 54.00 16.00 6.00 27.44 37.88
Intra-Diversity (60-80) 20.83 64.30 74.36 52.20 14.90 4.50 35.98 38.15
Intra-Diversity (80-100) 19.78 64.77 74.51 56.50 13.00 5.20 37.50 38.75
Table 7: Validation Results of Inter-Diversity and Intra-Diversity on Qwen2-7B across various benchmarks.
Qwen2-7B Common Sense Reasoning Mathematics Coding Avg
NQ TriviaQA Hellaswag GSM8K MATH MBPP HumanEval
Raw 8.03 59.58 73.00 78.00 5.70 5.00 60.98 41.47
Rand 13.28 58.27 73.00 75.35 35.36 52.20 63.72 53.02
Inter-Diversity (0-20) 15.18 59.28 73.34 74.50 34.94 53.10 68.60 54.13
Inter-Diversity (20-40) 13.77 58.42 73.18 73.60 32.55 53.00 64.33 52.69
Inter-Diversity (40-60) 14.62 58.58 73.35 72.50 34.50 52.50 61.28 52.47
Inter-Diversity (60-80) 14.31 58.60 73.33 74.80 33.90 51.40 60.68 52.43
Inter-Diversity (80-100) 9.30 57.72 73.14 74.60 28.00 51.30 63.42 51.07
Intra-Diversity (0-20) 12.64 58.54 73.35 75.10 8.75 51.10 61.59 48.72
Intra-Diversity (20-40) 14.17 58.78 73.10 74.10 29.20 52.10 63.41 52.12
Intra-Diversity (40-60) 15.24 58.57 73.12 74.70 32.50 51.80 64.02 52.85
Intra-Diversity (60-80) 14.02 57.40 73.06 75.20 32.20 53.50 66.77 53.16
Intra-Diversity (80-100) 11.91 57.88 73.29 75.00 36.05 52.50 66.16 53.25
Table 8: Validation Results of Inter-Diversity and Intra-Diversity on Qwen2.5-7B across various benchmarks.
Qwen2.5-7B Common Sense Reasoning Mathematics Coding Avg
NQ TriviaQA Hellaswag GSM8K MATH MBPP HumanEval
Raw 8.84 58.14 72.75 78.20 9.10 7.40 78.05 44.64
Rand 11.46 57.85 73.08 78.90 13.15 62.50 71.65 52.65
Inter-Diversity (0-20) 13.23 58.15 73.27 78.70 11.45 62.30 69.21 52.33
Inter-Diversity (20-40) 10.81 58.11 73.02 77.90 16.95 62.30 68.29 52.48
Inter-Diversity (40-60) 10.75 57.89 72.90 73.30 26.70 62.80 69.51 53.41
Inter-Diversity (60-80) 10.43 58.19 73.10 78.40 17.05 62.80 71.95 53.13
Inter-Diversity (80-100) 10.00 58.10 73.11 77.30 16.45 62.30 67.07 52.05
Intra-Diversity (0-20) 10.68 58.52 73.18 80.10 25.80 62.50 68.90 54.24
Intra-Diversity (20-40) 11.21 58.14 73.02 79.50 17.75 62.90 67.38 52.84
Intra-Diversity (40-60) 11.57 58.11 72.94 76.00 15.65 62.50 65.25 51.72
Intra-Diversity (60-80) 10.89 57.91 72.92 75.80 11.35 62.00 66.16 51.00
Intra-Diversity (80-100) 12.79 58.09 73.21 75.40 16.05 62.50 49.39 49.63
D.8 Complete Results of DaaR with Baselines
Additionally, we include the complete results of DaaR and the comparative baselines in the three tables: Table 9, Table 10 and Table 11. The results illustrate the challenges of the scenario, particularly for the Qwen2 series, where baseline methods struggle to outperform random selection. Furthermore, they demonstrate the robustness and effectiveness of our approach, consistently achieving the highest average scores across different models.
Table 9: Performance of DaaR with baselines on Llama3.1-8B across various benchmarks.
Llama3.1-8B Common Sense Reasoning Mathematics Coding Avg
NQ TriviaQA Hellaswag GSM8K MATH MBPP HumanEval
Raw 14.13 65.90 74.62 54.80 7.90 5.00 28.66 35.86
Rand 21.99 64.83 74.72 55.70 14.50 5.10 24.09 37.27
Instruction Len 15.34 63.60 73.73 54.00 15.40 3.60 30.80 36.64
Alpagasus (Chen et al., 2024d) 21.57 64.37 74.87 55.20 17.65 4.60 16.16 36.34
Instag-C (Lu et al., 2024) 18.12 64.96 74.01 55.70 15.50 4.80 37.81 38.70
Instag-D (Lu et al., 2024) 21.94 64.69 74.87 54.80 12.80 4.10 9.76 34.71
SuperFilter (Li et al., 2024a) 22.95 64.99 76.39 57.60 6.05 2.60 40.55 38.73
Deita-C (Liu et al., 2024b) 15.58 64.97 74.21 55.00 13.05 4.60 34.46 37.41
Deita-Q (Liu et al., 2024b) 19.57 64.22 75.15 54.00 7.20 4.20 28.35 36.10
Deita-D (Liu et al., 2024b) 20.97 63.32 75.10 54.90 7.00 4.00 31.71 36.71
DaaR (Ours) 20.08 64.55 74.88 54.8 15.30 4.70 37.50 38.83
Table 10: Performance of DaaR with baselines on Qwen2-7B across various benchmarks.
Qwen2-7B Common Sense Reasoning Mathematics Coding Avg
NQ TriviaQA Hellaswag GSM8K MATH MBPP HumanEval
Raw 8.03 59.58 73.00 78.00 5.70 5.00 60.98 41.47
Rand 13.28 58.27 73.00 75.35 35.36 52.20 63.72 53.02
Instruction Len 8.62 58.44 72.86 73.30 27.05 53.10 63.72 51.01
Alpagasus (Chen et al., 2024d) 13.67 57.94 73.04 73.90 32.30 51.40 63.41 52.24
Instag-C (Lu et al., 2024) 9.51 58.50 73.06 74.70 35.35 51.90 64.70 52.53
Instag-D (Lu et al., 2024) 12.87 57.48 72.80 74.40 33.75 51.80 64.02 52.45
SuperFilter (Li et al., 2024a) 19.16 58.98 72.99 73.70 30.10 52.40 58.85 52.31
Deita-C (Liu et al., 2024b) 8.94 58.07 73.06 73.90 35.55 52.90 62.20 52.09
Deita-Q (Liu et al., 2024b) 14.06 59.07 73.16 75.80 35.50 23.00 58.24 48.40
Deita-D (Liu et al., 2024b) 16.41 57.80 72.70 76.10 29.05 52.40 64.63 52.73
DaaR (Ours) 16.88 57.58 73.03 75.40 38.1 52.00 64.94 53.99
Table 11: Performance of DaaR with baselines on Qwen2.5-7B across various benchmarks.
Qwen2.5-7B Common Sense Reasoning Mathematics Coding Avg
NQ TriviaQA Hellaswag GSM8K MATH MBPP HumanEval
Raw 8.84 58.14 72.75 78.20 9.10 7.40 78.05 44.64
Rand 11.46 57.85 73.08 78.90 13.15 62.50 71.65 52.65
Instruction Len 11.34 58.01 72.79 78.00 15.80 62.30 68.12 52.34
Alpagasus (Chen et al., 2024d) 10.40 57.87 72.92 77.20 18.75 61.80 65.55 52.07
Instag-C (Lu et al., 2024) 10.81 58.45 73.27 76.00 13.30 61.80 68.29 51.70
Instag-D (Lu et al., 2024) 11.08 58.40 72.79 76.40 16.40 62.90 70.43 52.63
SuperFilter (Li et al., 2024a) 13.54 58.51 72.89 79.30 11.35 39.50 65.25 48.62
Deita-C (Liu et al., 2024b) 10.50 58.17 73.14 74.60 16.60 62.00 72.26 52.47
Deita-Q (Liu et al., 2024b) 11.24 57.83 72.97 78.50 12.95 38.10 67.68 48.47
Deita-D (Liu et al., 2024b) 10.48 57.81 73.05 77.20 15.25 52.90 69.21 50.84
DaaR (Ours) 15.83 58.65 72.48 80.20 16.70 64.20 68.29 53.76

### v2

@@
 In practical scenarios, existing methods based on modeling the mixture proportions of data composition often struggle with data whose domain labels are missing, imprecise or non-normalized, while methods based on data selection usually encounter difficulties in balancing multi-domain performance.
-To address these challenges, in this paper, we study the role of data diversity in enhancing the overall abilities of LLMs by empirically constructing contrastive data pools and theoretically deriving explanations for both inter- and intra-diversity.
+To address these challenges, in this work, we investigate the role of data diversity in enhancing the overall abilities of LLMs by empirically constructing contrastive data pools and theoretically deriving explanations.
 Building upon the insights gained, we propose a new method that gives the LLM a dual identity: an output model to cognitively probe and select data based on diversity reward, as well as an input model to be tuned with the selected data.
 Extensive experiments show that the proposed method notably boosts performance across domain-undetermined data and a series of foundational downstream tasks when applied to various advanced LLMs.
-We release our code and hope this study can shed light on the understanding of data diversity and advance feedback-driven data-model co-development for LLMs.
-Machine Learning, ICML
+We release our code and hope this study can shed light on the understanding of data diversity and advance feedback-driven data-model co-design for LLMs.
+11footnotetext: Equal contribution.
 1 Introduction
-The advancement of large language models (LLMs), exemplified by open-source models such as the Llama3 series (Dubey et al., 2024), Qwen2 series (Yang et al., 2024), and the DeepSeek-V3 series (Liu et al., 2024a), along with closed-source models such as GPT-4 (Achiam et al., 2023), has revolutionized artificial intelligence by enhancing capabilities in foundational abilities such as common sense, reasoning, mathematics, and coding.
-Fine-tuning these models further optimizes their usability by enhancing their performance and aligning with specific human instructions and preferences  (Taori et al., 2023; Rafailov et al., 2023).
-To cultivate comprehensive capabilities in LLMs, fruitful studies have explored preferable trade-offs between quality, quantity, and diversity of their training data  (Li et al., 2024b; Chen et al., 2024b; Qin et al., 2024; Zhao et al., 2024a).
-For example, methods focusing on data selection (Li et al., 2024a; Xia et al., 2024; Wang et al., 2023) and mixture (Ge et al., 2024) demonstrate promising capabilities to enhance model performance, particularly through semantic diversity  (Lu et al., 2024; Liu et al., 2024b).
-However, real-world applications frequently encounter unlabeled data and difficulties in domain labeling (Ge et al., 2023), posing challenges for data mixture methodologies that take the domain tags as a prior, as well as the data selection approaches, which often prioritize quality over diversity especially with data sourced from quite different domains.
-To leverage the best of both worlds, in this work, we propose a new fine-tuning method for LLMs named DaaR, which encourages the given model to learn diversity as a reward signal, and then to autonomously select domain-undetermined training data to achieve a theoretically informed trade-off to maximize diversity for overall model performance enhancement.
-Our investigation begins by diving into the semantic diversity of LLM data, considering how to explicitly model it and how it can influence the overall performance of LLMs.
-We systematically construct contrastive data pools and conduct extensive empirical examinations of different distributional performances across various foundational LLM capabilities.
-Based on these observations, we provide theoretical discussions on a general assumption that a mixture of underlying component LLMs determines the data mixing effect, and the explanations regarding inter- and intra-diversity modeled from a data semantic perspective.
-Figure 1: The t-SNE visualization of embeddings for data samples with different distributions on Qwen2-7B.
-(a) The data pool of all 40,000 samples, (b) Randomly selected subset, (c) Distribution of data farthest from other domain centroids based on Inter-Diversity, (d) Distribution of data closest to other domain centroids based on Inter-Diversity, (e) Distribution of data closest to its own domain centroid based on Inter-Diversity, (f) Distribution of data farthest from its own domain centroid based on Inter-Diversity.
-Our theoretical insights illuminate that optimal solutions for overall model performance enhancement cannot be explicitly applied to domain-undetermined data.
-To address this, we integrate an external multi-layer perceptron (MLP) structure into the LLM to be tuned, acting as a trainable probe module conditioned on the inherent knowledge and weights of the given LLM.
-The probe module takes semantic entropy as a proxy measure of diversity, and output reward scores, and select suitable data subsets aligned with the LLM’s underlying mixture distributions for better diversity.
-This is achieved by training the probe module on synthetically generated domain-aware data by the model itself, and using the data after selection to further self-fine-tuning.
-Extensive experiments on various state-of-the-art (SOTA) LLMs like the Qwen and LLaMA series verify the effectiveness of DaaR in enhancing overall model capabilities over many SOTA baseline methods.
-Our contributions are threefold:
-  • We empirically and theoretically demonstrate how data plays a crucial role to enhance LLMs in semantic diversity, providing generalized modeling based on mixture of component models.
-  • We propose a lightweight, self-rewarding method for selecting data in domain-undetermined scenarios, which utilizes synthesized model-aware data to maximize domain diversity and fine-tuning performance.
-  • We verify that the proposed methods can effectively balance seven diverse benchmarks for improved overall performance when applied to representative LLMs, whereas other SOTA methods struggle in such challenging scenarios.
-Our code is open-sourced  [^1] https://github.com/modelscope/data-juicer/tree/DaaR to foster more in-depth understanding and future progress.
-2 Related Work
+The rapid advancement of large language models (LLMs) — exemplified by open-source series such as LLaMA llama3series , Qwen qwen2series , and DeepSeek deepseekv3 , as well as closed-source models like GPT gpt4 — has transformed artificial intelligence by significantly enhancing capabilities in core areas such as common sense reasoning, mathematics, and code generation.
+Fine-tuning these models further improves their usability by aligning their behaviors with specific instructions and human preferences alpaca2023 ; dpo .
+To cultivate comprehensive capabilities in LLMs, prior studies have explored preferable trade-offs between quality, quantity, and diversity of their training data  li2024quantity ; djv2 ; qin2024surveycodev ; zhao2024beyond .
+For example, methods focusing on data selection li2024super ; xialess ; wang2023far and mixture ge2024bimix demonstrate promising capabilities to enhance model performance, particularly through semantic diversity  lu2023instag ; liu2023deita .
+However, real-world applications frequently encounter unlabeled data and difficulties in domain labeling ge2023openagi , posing challenges for data mixture methodologies that take the domain tags as a prior, as well as the data selection approaches, which often prioritize quality over diversity, especially with data sourced from quite different domains.
+To leverage the best of both worlds, in this work, we dive into the role of semantic diversity of LLM data.
+Our investigation begins with a systematic analysis of 40k instruction pairs across four capability domains.
+Through controlled experiments with contrastive data pools, we uncover two insights: (1) Optimal diversity thresholds for model performance vary significantly across architectures and task domains; (2) Current centroid-based diversity metrics lu2023instag fail to capture the dynamic interaction between inter-domain separation and intra-domain variance when labels are unavailable.
+These findings motivate our theoretical analysis, which establishes that traditional data mixing strategies assuming known domain weights ye2024data cannot achieve Pareto-optimal performance in label-free settings.
+We prove that when domain labels are unavailable, the optimal data selection strategy must simultaneously maximize the mutual information between selected samples and preserve the base model’s latent feature geometry.
+We operationalize these insights through DaaR, a self-supervised framework that learns diversity as a reward signal through three technical innovations: (1) Automatic synthesis of model-aware domain centroids via iterative embedding-space generation, (2) A lightweight MLP probe trained to predict semantic entropy using only the LLM’s frozen embeddings, and (3) Closed-loop fine-tuning where the model’s own diversity estimates guide subsequent data selection.
+Extensive evaluations across 7 benchmarks and 3 model families reveal that DaaR achieves new state-of-the-art (SOTA) average performance, consistently outperforming 9 baseline methods on high-difficulty tasks while maintaining computational efficiency.
+Crucially, our method requires no domain labels or external models - the LLM self-supervises its diversity estimation through geometric constraints in its native embedding space.
+Our core contributions for LLM data optimization are three-fold:
+  • Formal characterization of diversity’s dual role in LLM fine-tuning through controlled experiments with contrastive data pools.
+  • A new method with label-free diversity reward mechanism using embedding-space entropy prediction, theoretically grounded in mutual information maximization.
+  • Extensive empirical validations showing consistent gains across model families and tasks, with notable improvements in mathematical reasoning (+27%) and coding (+7.4%), whereas other baselines struggle in such challenging scenarios.
+Our code is released at https://github.com/modelscope/data-juicer/tree/DaaR to foster more data-centric research for LLMs.
+2 Preliminaries
+2.1 Related Works
 Data Selection
-Fine-tuning is a pivotal training paradigm for enhancing LLMs’ task-specific and domain-specific capabilities.
-Research has shown that a small set of instruction pairs can enable LLMs to follow major instructions effectively (Zhou et al., 2024; Chen et al., 2024a).
-This fine-tuning can be achieved through rule-based methods, which focus on attributes such as Error L2-Norm (Paul et al., 2021) and token length (Raffel et al., 2020).
-More recently, model-based heuristics have been explored, including methods based on instruction-following difficulty (Li et al., 2024a), GPT scoring (Chen et al., 2024d; Wettig et al., 2024), data model selection (Engstrom, 2024), and influence scores derived from loss (Xia et al., 2024).
+Fine-tuning is a pivotal training paradigm for enhancing LLMs’ domain-specific capabilities.
+Research has shown that a small set of instruction pairs can enable LLMs to follow major instructions effectively zhou2024lima ; djv1 .
+This fine-tuning can be achieved through rule-based methods, which focus on attributes such as Error L2-Norm paul2021el2n and token length raffel2020exploring .
+More recently, model-based heuristics have been explored, including methods based on instruction-following difficulty li2024super , GPT scoring chen2024alpagasus ; 2024qurating , data model selection 2024dsdm , and influence scores derived from loss xialess .
 Diversity & Data Mixing
-A fundamental principle for LLMs is being able to handle diverse human requests, underscoring data diversity as essential for effective fine-tuning (Yu et al., 2022; Ding et al., 2023).
-Data diversity encompasses aspects such as data deduplication (Abbas et al., 2023), the coverage scope of tags (Lu et al., 2024), model-based diversity evaluation (Liu et al., 2024b; Zhang et al., 2024), and scaling properties (Song et al., 2024).
-Typically, data mixing approaches (Albalak et al., 2023; Ye et al., 2024; Ge et al., 2024) focus on adjusting the proportional weights of different domains to enhance model capabilities.
+A fundamental principle for LLMs is being able to handle diverse human requests, underscoring data diversity as essential for effective fine-tuning yu2022can ; ding2023enhancing .
+Data diversity encompasses aspects such as data deduplication abbas2023semdedup , the coverage scope of tags lu2023instag , model-based diversity evaluation liu2023deita ; zhang2024harnessing , and scaling properties song2024scaling .
+Typically, data mixing approaches albalak2023efficient ; ye2024data ; ge2024bimix focus on adjusting the proportional weights of different domains to enhance model capabilities.
 Our Position
-From a modeling perspective, our work shares a focus with many data mixing methodologies regarding diversity measurement.
-However, we address a more challenging setting relevant to real-world applications, where available datasets often lack domain labels.
-This situation presents challenges for existing mixing methods in balancing label granularity and normalization, and in determining which domains should be considered as candidates for improved solution spaces and tuning performance.
-Methodologically, our approach aligns with the data selection category but differs in both the learning process (utilizing a rewarding model) and the selection criteria (employing semantic entropy).
-Additionally, we revisit the applicability of existing selection methods in multi-domain scenarios.
-This area is under-explored and requires new insights to avoid biased modeling of diverse domain distributions—a topic we further investigate in this paper.
-3 Motivational Observations
-In this section, we will explore how data with different domain-specific diversity affects model capabilities through experimental data pools, featuring contrastive data distributions in terms of inter-domain and intra-domain diversity.
+This work relates to data mixing and selection fields through its focus on diversity quantification, but advances both of them in three key aspects: First, we address practical challenges in real-world applications where unlabeled datasets hinder existing methods’ ability.
+Second, our framework distinguishes by integrating reward-driven learning with semantic entropy-based selection criteria.
+Third, we critically examine the under-explored applicability of current selection paradigms in multi-domain contexts, revealing critical limitations in modeling diverse distribution patterns.
+2.2 Problem Setup
+In this paper, we consider learning from domain-undetermined data mixtures to enhance LLM downstream performance in target domains.
+Let $\{\mathcal{D}_{k}\}_{k=1}^{K}$ denote $K$ target domain distributions, where each $\mathcal{D}_{k}:\mathcal{X}\rightarrow[0,1]$ defines a probability distribution over samples $x\in\mathcal{X}$.
+For model parameters $\theta$, the domain-specific performance is measured by $\mathbb{E}_{x\sim\mathcal{D}_{k}}[\ell(x;\theta)]$.
+Our objective is to minimize the weighted multi-domain loss: $\sum_{k=1}^{K}\lambda_{k}\mathbb{E}_{x\sim\mathcal{D}_{k}}[\ell(x;\theta)],$ where $\lambda_{1:K}$ are non-negative weights with $\sum_{k=1}^{K}\lambda_{k}=1$, reflecting application-specific priorities.
+For each domain $\mathcal{D}_{k}$, we assume only a natural language description $d_{k}$ is available while NO target domain samples are accessible, contrasting with conventional data selection methods requiring labeled examples xialess ; xie2023data ; renduchintala2024smart ; ye2024data in LLM training.
+We are provided with an domain-unlabeled dataset $\mathcal{D}=\{x_{i}\}_{i=1}^{N}$ containing domain-agnostic samples.
+A straightforward approach is direct fine-tuning using all $\mathcal{D}$, but it often yields suboptimal results (empirically evidenced in Table 2).
+Instead, we propose selecting a subset via binary vector $\alpha\in\{0,1\}^{N}$, where $\alpha_{i}=1$ indicates $x_{i}$ is selected.
+The LLM is trained by solving: $\min_{\theta}\sum_{i=1}^{N}\alpha_{i}\ell(x_{i};\theta)$, leading to the bi-level optimization:
+$\small\min_{\alpha}\sum_{k=1}^{K}\lambda_{k}\mathbb{E}_{x\sim\mathcal{D}_{k}}[\ell(x;\theta^{*}(\alpha))],\quad\text{s.t.}\ \ \theta^{*}(\alpha)=\arg\min_{\theta}\sum_{i=1}^{N}\alpha_{i}\ell(x_{i};\theta).$ (1)
+3 Diversity: A Critical Factor for LLM Fine-Tuning
+In this section, we investigate how domain-specific diversity in labeled training data modulates model capabilities through empirical analysis of well-designed data pools, complemented by a theoretical perspective that explains the observation and provides mechanistic insights.
+Figure 1: The illustration highlights our observations and the proposed DaaR method.
+Observation shows the t-SNE visualization of embeddings for data samples with different distributions, leading to varying LLM’s evaluation performance.
+In the domain-undetermined data selection scenario, DaaR introduces a dual-identity framework for LLMs: ① An output model to probe and select data based on diversity reward, ② An input model to be tuned with the selected data.
+The diversity probe is trained on model-aware synthetic data, enabling domain discrimination and diversity reward prediction.
 3.1 Seed Data Pools and Basic Setting
-Data Pools and Sources
-To explore how selecting data samples from extensive and diverse data repositories affects the foundational capabilities of LLMs, we construct various data pools and consistently fine-tune LLMs on them.
-We then analyze the performance changes and attribute these changes to the different controlled data pools.
-The seed data pool is sourced from following datasets: Dolly-15k (Conover et al., 2023) for common sense, Cot-en (Chung et al., 2024) for reasoning, Math-Instruct (Yue et al., 2024) for mathematics, and Code-Alpaca (Chaudhary, 2023) for coding.
+Data Pools and Sources To explore how selecting data samples from extensive and diverse data repositories affects the foundational capabilities of LLMs, we construct various data pools and consistently fine-tune LLMs on them.
+The seed data pool is sourced from the following datasets: Dolly-15k dolly-15k for common sense, Cot-en cot-en for reasoning, Math-Instruct math-instruct for mathematics, and Code-Alpaca code-alpaca for coding.
 Each dataset was randomly tailored to 10,000 entries, resulting in a combined data pool of 40,000 entries.
-Following instruction tuning practices (Zhou et al., 2024; Liu et al., 2024b), we then uniformly sample 8,000 data entries as a referenced data pool for random baseline.
-In subsequent sections, we will introduce how to construct other comparative pools with the same size to random pool.
-Benchmarks
-Aligned with representative capabilities of leading open-source LLMs, we select the following widely used evaluation sets: NQ (Kwiatkowski et al., 2019) and TriviaQA (Joshi et al., 2017) for common sense, Hellaswag (Zellers et al., 2019) for reasoning, GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) for mathematics, MBPP (Austin et al., 2021) and HumanEval (Chen et al., 2021) for coding.
+Following instruction tuning practices zhou2024lima ; liu2023deita , we then uniformly sample 8,000 data entries as a reference data pool for the random baseline.
+Benchmarks Aligned with representative capabilities of leading open-source LLMs, we select the following widely used evaluation sets: NQ nq and TriviaQA triviaqa for common sense, Hellaswag hellaswag for reasoning, GSM8K gsm8k and MATH math for mathematics, MBPP mbpp and HumanEval humaneval for coding.
 To evaluate the comprehensive performance of LLMs across domains, we employ the average metric (Avg) as the primary evaluation criterion.
-Models & Implementation
-To ensure the effectiveness and applicability of our empirical findings, we employ the Qwen2 series (Qwen2-7B & Qwen2.5-7B) (Yang et al., 2024) and the Llama3.1-8B (Dubey et al., 2024) as representative SOTA base models to be fine-tuned.
-All experiments are conducted under identical training and evaluation protocols with two independent repetitions.
-Full platform, training and evaluation details are provided in Appendix B.
+Models & Implementation We employ the Qwen2 series (Qwen2-7B & Qwen2.5-7B) qwen2series and the Llama3.1-8B llama3series as representative SOTA base models to be fine-tuned.
+All experiments are conducted under identical training and evaluation protocols with two independent repetitions. Full platform, training and evaluation details are provided in Appendix B.
 3.2 Data Pools with Contrastive Distributions
@@
 (A) Foundational Definitions
-Let the composite dataset $\mathcal{D}=\bigcup_{k=1}^{K}\mathcal{D}_{k}$ comprise $K=4$ distinct domains, where each domain subset $\mathcal{D}_{k}$ contains $N_{k}=|\mathcal{D}_{k}|$ samples.
-We represent each data instance through its semantic embedding $\mathbf{x}_{i}^{(k)}\in\mathbb{R}^{d}$ extracted from the Embedding layer of the pretrained LLM, capturing high-dimensional semantic features.
-The domain centroid $\mathcal{C}_{k}$ serves as the semantic prototype:
-$\small\mathcal{C}_{k}=\frac{1}{N_{k}}\sum_{i=1}^{N_{k}}\mathbf{x}_{i}^{(k)}.$ (1)
-This centroid-based representation enables geometric interpretation of domain characteristics in the embedding space. We dissect data diversity into two complementary aspects:
+Each domain dataset $\mathcal{D}_{k}$ contains $N_{k}=|\mathcal{D}_{k}|$ samples, and we represent each data sample through its semantic embedding $\mathbf{x}_{i}^{(k)}\in\mathbb{R}^{d}$ extracted from the embedding-layer of the pretrained LLM.
+The domain centroid $\mathcal{C}_{k}$ serves as the semantic prototype $\mathcal{C}_{k}=\frac{1}{N_{k}}\sum_{i=1}^{N_{k}}\mathbf{x}_{i}^{(k)}$.
+This centroid-based representation enables geometric interpretation of domain characteristics in the embedding space.
+We dissect data diversity into two complementary aspects:
 (B.1) Inter-Diversity
-It quantifies the diversity between distinct domains through centroid geometry. For sample $\mathbf{x}_{i}^{(k)}$, its cross-domain similarity is measured by:
-$\small\phi_{\text{inter}}(\mathbf{x}_{i}^{(k)})=\sum_{\begin{subarray}{c}j=1\\ j\neq k\end{subarray}}^{K}\frac{\mathbf{x}_{i}^{(k)}\cdot\mathcal{C}_{j}}{\|\mathbf{x}_{i}^{(k)}\|\|\mathcal{C}_{j}\|}.$ (2)
+It quantifies the diversity between distinct domains through centroid geometry.
+For sample $\mathbf{x}_{i}^{(k)}$, its cross-domain similarity is measured by $\phi_{\text{inter}}(\mathbf{x}_{i}^{(k)})$.
 The global inter-diversity metric $\Phi_{\text{inter}}$ computes the expected pairwise centroid distance:
-$\small\Phi_{\text{inter}}=\mathbb{E}_{k\neq l}\left[\|\mathcal{C}_{k}-\mathcal{C}_{l}\|_{2}\right]=\frac{1}{\binom{K}{2}}\sum_{k=1}^{K-1}\sum_{l=k+1}^{K}\|\mathcal{C}_{k}-\mathcal{C}_{l}\|_{2}.$ (3)
+$\phi_{\text{inter}}(\mathbf{x}_{i}^{(k)})=\sum_{\begin{subarray}{c}j=1\\ j\neq k\end{subarray}}^{K}\frac{\mathbf{x}_{i}^{(k)}\cdot\mathcal{C}_{j}}{\|\mathbf{x}_{i}^{(k)}\|\|\mathcal{C}_{j}\|},\quad\Phi_{\text{inter}}=\mathbb{E}_{k\neq l}\left[\|\mathcal{C}_{k}-\mathcal{C}_{l}\|_{2}\right]=\frac{1}{\binom{K}{2}}\sum_{k=1}^{K-1}\sum_{l=k+1}^{K}\|\mathcal{C}_{k}-\mathcal{C}_{l}\|_{2}.$ (2)
 This formulation reflects a key insight: maximizing $\Phi_{\text{inter}}$ encourages domain separation, while minimization leads to overlapping representations.
-Fig.
-1 demonstrates this continuum through t-SNE projections – high $\Phi_{\text{inter}}$ manifests as distinct cluster separation with clear margins (Fig.
-1.(c)), whereas low values produce entangled distributions (Fig.
-1.(d)).
-Full analysis is detailed in Appendix D.1.
+Observation of the upper part of Fig.
+1 (full-size version is in Appendix, Fig.
+16) demonstrates this continuum through t-SNE projections – high $\Phi_{\text{inter}}$ manifests as distinct cluster separation with clear margins (Fig.
+16.(c)), whereas low values produce entangled distributions (Fig.
+16.(d)).
+Full analysis is detailed in Appendix H.1.
 (B.2) Intra-Diversity
 Focusing solely on the separation between different domains may hinder the model’s ability to learn the knowledge specific to a given domain.
-Hence we measure variation within each domain.
-We calculate sample similarity to its domain center:
-$\small\phi_{\text{intra}}(\mathbf{x}_{i}^{(k)})=\frac{\mathbf{x}_{i}^{(k)}\cdot\mathcal{C}_{k}}{\|\mathbf{x}_{i}^{(k)}\|\|\mathcal{C}_{k}\|}.$ (4)
-And the domain-level variance metric is defined as:
-$\small\Phi_{\text{intra}}^{(k)}=\frac{1}{N_{k}}\sum_{i=1}^{N_{k}}\|\mathbf{x}_{i}^{(k)}-\mathcal{C}_{k}\|_{2}^{2}.$ (5)
-Controlled manipulation of $\Phi_{\text{intra}}$ reveals critical trade-offs: lower variance (tight clusters near $\mathcal{C}_{k}$) enhances domain-specific feature learning but risks over-specialization.
-Higher variance improves robustness at the cost of potential cross-domain interference.
+Hence we calculate sample similarity to its domain center in $\phi_{\text{intra}}(\mathbf{x}_{i}^{(k)})$, and the domain-level variance metric is defined as $\Phi_{\text{intra}}^{(k)}$.
+$\phi_{\text{intra}}(\mathbf{x}_{i}^{(k)})=\frac{\mathbf{x}_{i}^{(k)}\cdot\mathcal{C}_{k}}{\|\mathbf{x}_{i}^{(k)}\|\|\mathcal{C}_{k}\|},\quad\Phi_{\text{intra}}^{(k)}=\frac{1}{N_{k}}\sum_{i=1}^{N_{k}}\|\mathbf{x}_{i}^{(k)}-\mathcal{C}_{k}\|_{2}^{2}.$ (3)
+Controlled manipulation of $\Phi_{\text{intra}}$ reveals critical trade-offs: lower variance enhances domain-specific learning but risks over-specialization, while higher variance improves robustness with the risk of cross-domain interference.
 The visualization in Fig.
@@
 The 20% interval results in five choices of data selection per domain, parameterizing the trade-off between diversity preservation and domain specificity.
-As demonstrated in Appendix D.2 (Fig.
+As demonstrated in Appendix H.2 (Fig.
 18 and Fig.
-11), this quantization process induces measurable distribution shifts.
+19), this quantization process induces measurable distribution shifts.
 3.3 Experimental Observations
@@
 Due to space constraints, we present only the results for the top 20%, middle 20%, and bottom 20%.
-More results can be found in Appendix D.7.
-Diversity-Aware Performance: Our diversity-controlled selections reveal two critical observations:
-  • Varied Improvement Patterns: Both models demonstrate marked improvements over Raw distributions across all diversity conditions, but the effects of their improvements vary.
-For Llama3.1-8B, Inter-D (80-100) achieves 38.98 average accuracy (+3.12 over Raw), outperforming the Random baseline by 1.71, while Inter-D (0-20) is below Random of 0.09.
+Detailed performance results on various downstream tasks are shown in Appendix H.4.
+Our diversity-controlled selections reveal two observations:
+Table 1: Performance of Llama3.1-8B and Qwen2-7B on Avg with different constructed Inter-Diversity and Intra-Diversity distributions.
+Models Distribution Avg
+Llama3.1 Raw 35.86
+Full (40K) 37.22
+Random (8K) 37.27
+Inter-D (0-20) 37.18
+Inter-D (40-60) 38.09
+Inter-D (80-100) 38.98
+Intra-D (0-20) 38.71
+Intra-D (40-60) 37.88
+Intra-D (80-100) 38.75
+Qwen2 Raw 41.47
+Full (40K) 53.01
+Random (8K) 53.02
+Inter-D (0-20) 54.13
+Inter-D (40-60) 52.47
+Inter-D (80-100) 51.07
+Intra-D (0-20) 48.72
+Intra-D (40-60) 52.85
+Intra-D (80-100) 53.25
+  • Varied Improvement Patterns: Both models demonstrate marked improvements over Raw distributions across diversity conditions, but the effects of their improvements vary.
+For Llama3.1-8B, Inter-D (80-100) achieves 38.98 average accuracy (+3.12 over Raw), outperforming the Random baseline by 1.71, while Inter-D (0-20) is below Random.
   • Model-Dependent Performance Peak: Each model exhibits distinct optimal operating points along the diversity spectrum.
@@
 These results show the promising potential of diversity-aware data selection, motivating us to further understand the performance variance more formally and propose principled solutions to adaptively achieve the performance peaks.
-Practical Challenges
-Despite existing positive improvements on overall performance, the optimal distribution parameters exhibit model-dependent variability.
-This parameter sensitivity suggests the existence of multiple local optima in the diversity-performance landscape.
-Two constraints merit consideration for real-world applications.
+Remarks Despite existing positive improvements on overall performance, two constraints merit consideration for real-world applications.
+(1) Distribution Transiency: The optimal diversity parameters (e.g., 80-100 vs.
+40-60) show sensitivity across tasks and models, necessitating automated and potentially costly methods.
 (2) Label Dependency: The studied heuristic strategies Inter-Diversity and Intra-Diversity currently require domain-labeled data for centroid calculation.
-(2) Distribution Transiency: The optimal diversity parameters (e.g., 80-100 vs.
-40-60) show sensitivity across tasks and models, necessitating automated and potentially costly configuration search.
-Table 1: Performance of Llama3.1-8B and Qwen2-7B on various downstream task benchmarks under different constructed Inter-Diversity (Inter-D) and Intra-Diversity (Intra-D) distributions.
+3.4 Theoretical Perspective
+To explain why diversity serves as a crucial factor in LLM’s cross-domain performance and addresses distribution transiency, we present a theoretical analysis based on important sampling.
+We first formulate the data selection dynamics and then incorporate diversity into it, thereby revealing the mechanistic basis and offering insights for our method.
+Importance Sampling
+We assume that there exists an ideal target distribution $q$ that minimizes the multi-domain loss in Eq.
+(1).
+Hence, the process of data selection can be framed as aligning the source distribution $p$ (empirical distribution of $\mathcal{D}$) toward the optimal distribution $q$.
+To bridge this distribution shift, we adopt an importance sampling perspective xie2023data ; xie2023doremi ; xialess where the optimal subset corresponds to reweighting samples by the density ratio $\frac{q(x)}{p(x)}$:
+[…]
+The domains of LLM’s data are commonly defined by textual descriptions, which inherently constrain the semantic scope. Thus we make the following assumption regarding distributions $p$ and $q$:
+Assumption 3.1 (Shared Support).
+For any given domain index $c\in\{1,\dots,K\}$, the support of $p$ and $q$ is shared, formalized as: $p(x|c=k)=q(x|c=k),\quad k=1,\dots,K.$
+Assumption 3.1 implies that the discrepancy between the source distribution $p$ and the target distribution $q$ arises solely from differences in domain proportions, leading to the following proposition.
+Proposition 3.2.
+The diversity is entirely attributable to the relative weights assigned to each domain:
+$\frac{q(x)}{p(x)}=\sum_{k=1}^{K}\lambda_{k}p(c{=}k|x),\lambda_{k}:=\frac{q(c=k)}{p(c=k)}$ (5)
+The derivation is detailed in Appendix E. This decomposition reveals that the Inter-Diversity and Intra-Diversity are governed by $\lambda_{k}$ and $p(c{=}k|x)$, leading to the following analysis.
+Why Overlook Diversity is Suboptimal
+We define $k^{*}(x):=\arg\max_{k}p(c=k|x)$ as the prediction of the classifier $p(c=k|x)$.
+Consider a special case where the classifier yields deterministic predictions, i.e.
+$p(c=k^{*}(x)|x)=1,\forall x\in\mathcal{D}$.
+In this case, the importance weight reduces to $\frac{q(x)}{p(x)}=\sum_{k=1}^{K}\lambda_{k}p(c=k|x)=\lambda_{k^{*}(x)}$.
+Therefore, importance sampling can be applied within:
+$\small\sum_{x\in\mathcal{D}}\frac{q(x)}{p(x)}\ell(x;\theta)=\sum_{x\in\mathcal{D}}\lambda_{k^{*}(x)}\ell(x;\theta)=\sum_{k=1}^{K}\ \sum_{x\in\mathcal{D}:k^{*}(x)=k}\lambda_{k*(x)}\ell(x;\theta).$ (6)
+This formulation is commonly used when domain labels are available, under the implicit assumption that domains are mutually exclusive.
+However, such an assumption rarely holds in the context of LLM training, as cross-domain data may either synergize or conflict zhao2024beyond ; xie2023doremi (evidence in Table 1).
+Hence, neglecting diversity in data selection potentially leads to suboptimal overall performance.
+Guidance to Our Method
+Under the domain-undetermined scenario, our proposed method DaaR operationalizes the gained insights by (1) constructing pseudo-labels (Section 4.1) align with LLM-aware generated seed, (2) explicitly modeling the classifier $p(c{=}k|x)$ through a domain discrimination probe (Section 4.2), and (3) using the probe’s predictive entropy (Section 4.2) as a proxy for diversity, estimating the ability of classifier $p(c{=}k|x)$ to select data toward the target distribution $q$.
+4 DaaR: Diversity as a Reward
+To address the challenges identified in Sec.
+3.3 and leverage the insights gained in Sec.
+3.4, we establish a data selection method DaaR guided by diversity-aware reward signals.
+It comprises three key components illustrated in Fig.
+1: (1) model-aware centroid synthesis, (2) two-stage training with reward probe, and (3) diversity-driven data selection.
+4.1 Model-Aware Training Data
+Model-aware Centroid Construction
+The proposed method initiates with centroid self-synthesis through a two-phase generation process to address two fundamental challenges: (1) eliminating dependency on human annotations through automated domain prototyping, and (2) capturing the base model’s intrinsic feature space geometry for model-aware domain separation.
+  • Phase 1 - Seed Generation: For each domain $k$, generate seed samples $\mathcal{S}_{k}^{(0)}$ via zero-shot prompting with domain-specific description templates generated by LLM itself, alongside with minor injection from downstream task samples, establishing initial semantic anchors.
+  • Phase 2 - Diversity Augmentation: Iteratively expand $S_{k}^{(t)}$ through context-aware generation, conditioned on a sliding window buffer with 3 random anchors sampled from the $(t-1)$ iteration $S_{k}^{(t-1)}$.
+The generated sample $\mathbf{x^{\prime}}$ is retained through rejection sampling to enhance diversity.:
+    $\max_{\mathbf{x}\in S_{k}^{(t-1)}}\cos(\mathcal{M}_{\text{ebd}}(\mathbf{x^{\prime}}),\mathcal{M}_{\text{ebd}}(\mathbf{x}))<\tau,$ (7)
+    where $\mathcal{M}_{\text{ebd}}(\cdot)$ indicates the output of the embedding layer of the given LLM, with $\tau$ as the similarity threshold.
+This process terminates when it reaches the predetermined iteration.
+The implementation of all parameters, along with their corresponding ablation studies, is provided in Appendix C.1.
+The domain centroid is then computed from the final augmented set $\mathcal{S}_{k}$ using the model’s embedding, formalized as $\mathcal{C}_{k}=1/|\mathcal{S}_{k}|\cdot\sum_{\mathbf{x}_{i}\in\mathcal{S}_{k}}\mathcal{M}_{\text{ebd}}(\mathbf{x}_{i})$.
+This captures the LLM’s intrinsic feature space geometry while eliminating dependency on human annotations.
+Domain-Aware Clustering
+We then automatically construct pseudo-labels for the given data samples based on the previously synthesized centroids $\{C_{k}\}_{k=1}^{K}$.
+We perform constrained k-means clustering in the embedding space with $\arg\min_{\{S_{k}\}}\sum_{k=1}^{K}\sum_{\tilde{x}\in S_{k}}\|\mathcal{M}_{\text{ebd}}(\tilde{x})-\mathcal{C}_{k}\|^{2}$.
+This produces the seed dataset $\mathcal{D}_{\text{probe}}$ containing pseudo-labels $\{\tilde{y}_{i}\}_{i=1}$ where $\tilde{y}_{i}\in\{1,\ldots,K\}$, with model-induced and embedding-derived domain label assignments.
+4.2 Training for Self-Rewarding Abilities
+Entropy as a Diversity Proxy
+While diversity metrics in Eqs.
+(2)-(3) directly quantify cross-sample distribution, we aim to design a reward probe that outputs with sample-level diversity level.
+Since models require sample-level processing that is unable to leverage pairwise relationships, direct training for diversity leads to poor performance in Appendix F.4.
+We instead use softmax confidence scores and predictive entropy as an implicit diversity proxy, reflecting model-aware data discriminability.
+This enables effective diversity approximation through a two-stage framework.
+Stage 1: Domain Predictor
+The proposed DaaR establishes model-aware domain discrimination abilities through a multi-layer perceptron (MLP) probe module, $\psi_{\text{dom}}$, attached to the hidden layer $\mathcal{M}(\tilde{x})$ of the LLMs.
+The probe will be trained meanwhile all the parameters of the LLM are frozen, achieving a preferable balance between effectiveness and cost, with detailed analysis regarding the choice of layers presented in Appendix G.2.
+Specifically, with pseudo-label $\tilde{y}$, we can compute domain predicted probabilities as:
+$p_{k}(\tilde{x})=\text{softmax}\left(\psi_{\text{dom}}(\mathcal{M}(\tilde{x}))\right),\quad\psi_{\text{dom}}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{K},$ (8)
+$\psi_{\text{dom}}$ is optimized via cross-entropy loss $\mathcal{L}_{\text{dom}}=-\frac{1}{|\mathcal{D}_{\text{probe}}|}\sum_{(\tilde{x},\tilde{y})\in\mathcal{D}_{\text{probe}}}\sum_{k=1}^{K}\mathbb{I}_{[k=\tilde{y}]}\log p_{k}(\tilde{x})$, where $\mathbb{I}_{[k=\tilde{y}]}$ denotes the indicator function.
+We employ single-sample batches with the AdamW optimizer to prevent gradient averaging across domains.
+Training consistently converges and achieves 92.7% validation accuracy on domain classification, as shown in Fig.
+2 (a).
+Stage 2: Diversity Rewarding
+Building on the stabilized domain discrimination probe, we quantify sample-level diversity through predictive entropy $H(\tilde{x})$.
+And to enable efficient reward computation during data selection, we then train another 5-layer MLP $\psi_{\text{div}}$ to directly estimate $H(\tilde{x})$ from $\mathcal{M}(\tilde{x})$:
+$H(\tilde{x})=-\sum_{k=1}^{K}p_{k}(\tilde{x})\log p_{k}(\tilde{x}).\quad\hat{H}(\tilde{x})=\psi_{\text{div}}(\mathcal{M}(\tilde{x})),\psi_{\text{div}}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{+}.$ (9)
+This diversity probe module $\psi_{\text{div}}$ shares $\psi_{\text{dom}}$’s architecture up to its final layer that replaced with the regression head, trained using entropy-scaled MSE loss $\mathcal{L}_{\text{div}}=\frac{1}{|\mathcal{D}_{\text{probe}}|}\sum_{\tilde{x}\in\mathcal{D}_{\text{probe}}}(\hat{H}(\tilde{x})-H(\tilde{x}))^{2}$.
+The module is also well-converged as shown in Fig.
+2 (b).
+Figure 2: Training loss and validation process of the two stages of DaaR on Qwen2-7B, more detailed results in Appendix H.3.
+Data Selection: After training the module $\psi_{\text{div}}$, we can use its output to select data samples.
+Building on the theoretical insights in Sec.
+3.4, data points that are closer to other centroids and more dispersed within their own centroid are more beneficial for enhancing the comprehensive capabilities of the model.
+Therefore, we use the predicted entropy score as a reward, selecting the top 20% with the highest scores as the final data subset for fine-tuning.
+The analysis of stability is shown in Appendix G.8.
+5 Experiments
+To validate the efficacy of DaaR, we conduct comprehensive experiments on data pools and benchmarks in Sec.
+3.1, comparing SOTA baselines as follows with critical modifications: all domain-specific labels are deliberately stripped.
+This constraint mimics more challenging real-world scenarios and precludes direct comparison with data mixture methods requiring domain label prior.
+Baselines
+We use the following data selection methods for comprehensive evaluation: (1) Random Selection: traditional random sampling; (2) Instruction Len: measuring instruction complexity by token count  cao2023instruction ; (3) Alpagasus chen2024alpagasus : using ChatGPT for direct quality scoring of instruction pairs; (4-5) Instag lu2023instag : semantic analysis approach with Instag-C (complexity scoring via tag quantity) and Instag-D (diversity measurement through tag set expansion); (6) SuperFilter li2024super : response-loss-based complexity estimation using compact models; (7-9) Deita liu2023deita : model-driven evaluation with Deita-C (complexity scoring), Deita-Q (quality scoring), and Deita-D (diversity-aware selection).
+Detailed implementations for these baselines are in Appendix D.1.
+Table 2: Evaluation results of Llama3.1-8B, Qwen2-7B, and Qwen2.5-7B across various downstream task benchmarks. DaaR demonstrates superiority in AVG compared to other baselines.
 Models Distribution Common Sense Reasoning Mathematics Coding Avg
@@
 Llama3.1-8B Raw 14.13 65.90 74.62 54.80 7.90 5.00 28.66 35.86
-Random 21.99 64.83 74.72 55.70 14.50 5.10 24.09 37.27
-Inter-D (0-20) 19.28 65.79 74.44 54.90 6.50 4.30 35.06 37.18
-Inter-D (40-60) 23.70 65.14 74.86 56.40 17.15 5.00 24.40 38.09
-Inter-D (80-100) 23.76 64.43 75.20 56.40 15.05 4.50 33.54 38.98
-Intra-D (0-20) 22.08 65.08 75.00 54.70 16.20 4.40 33.54 38.71
-Intra-D (40-60) 22.12 64.74 74.87 54.00 16.00 6.00 27.44 37.88
-Intra-D (80-100) 19.78 64.77 74.51 56.50 13.00 5.20 37.50 38.75
+Full (40K) 21.92 65.11 73.62 51.70 7.60 4.10 36.50 37.22
+Random (8K) 21.99 64.83 74.72 55.70 14.50 5.10 24.09 37.27
+Instruction Len 15.34 63.60 73.73 54.00 15.40 3.60 30.80 36.64
+Alpagasus chen2024alpagasus 21.57 64.37 74.87 55.20 17.65 4.60 16.16 36.34
+Instag-Best lu2023instag 18.12 64.96 74.01 55.70 15.50 4.80 37.81 38.70
+SuperFilter li2024super 22.95 64.99 76.39 57.60 6.05 2.60 40.55 38.73
+Deita-Best liu2023deita 15.58 64.97 74.21 55.00 13.05 4.60 34.46 37.41
+DaaR (Ours) 20.08 64.55 74.88 54.80 15.30 4.70 37.50 38.83
 Qwen2-7B Raw 8.03 59.58 73.00 78.00 5.70 5.00 60.98 41.47
-Random 13.28 58.27 73.00 75.35 35.36 52.20 63.72 53.02
-Inter-D (0-20) 15.18 59.28 73.34 74.50 34.94 53.10 68.60 54.13
-Inter-D (40-60) 14.62 58.58 73.35 72.50 34.50 52.50 61.28 52.47
-Inter-D (80-100) 9.30 57.72 73.14 74.60 28.00 51.30 63.42 51.07
-Intra-D (0-20) 12.64 58.54 73.35 75.10 8.75 51.10 61.59 48.72
-Intra-D (40-60) 15.24 58.57 73.12 74.70 32.50 51.80 64.02 52.85
-Intra-D (80-100) 11.91 57.88 73.29 75.00 36.05 52.50 66.16 53.25
-4 Theoretical Analysis and Insights
-In this section, we first formalize the optimization dynamics of multi-domain data aggregation for enhancing LLMs’ comprehensive capabilities.
-Building on the empirical observations from Sec.
-3, we then establish theoretical connections between distributional diversity and emergent model behaviors.
-Finally, we demonstrate the fundamental limitation of explicit diversity optimization in undetermined-domain scenarios.
-4.1 Ideal Optimization Formulation
-During the training process of LLMs, various sources of labeled data are often collected to enhance models’ capabilities across multiple dimensions.
-Let $\mathcal{K}=\{1,\ldots,K\}$ represent the set of labeled domains, where $K$ is determined by real-world task requirements (e.g., $K=4$ in our experiments).
-For each domain $k\in\mathcal{K}$, data samples $\mathbf{x}_{i}^{(k)}\in\mathcal{D}_{k}$ are generated according to a distribution $\mathcal{D}_{k}$.
-The data distributions are assumed mutually distinct across different domains, leading to an ideal I.I.D.
-per-domain hypothesis: each domain $k$ corresponds to an independent model $h_{k}\in\mathcal{H}$.
-The standard optimization goal thus can be:
-$\small\forall k\in\mathcal{K},\quad\min_{h_{k}\in\mathcal{H}}\mathcal{L}_{\mathcal{D}_{k}}(h_{k}),$ (6)
-where $\mathcal{L}_{\mathcal{D}_{k}}(h_{k})=\mathbb{E}_{(x,y)\sim\mathcal{D}_{k}}\left[l(h_{k}(x),y)\right]$ is the empirical risk given a specific loss function $l$.
-Counterintuitive Findings
-If the distributions $\{\mathcal{D}_{k}\}_{k=1}^{K}$ are strictly I.I.D., the risk $\mathcal{L}_{\mathcal{D}_{k}}(h_{k})$ should depend only on $\mathcal{D}_{k}$.
-However, numerous previous studies (Zhao et al., 2024a; Tirumala et al., 2023) have demonstrated the presence of synergistic and antagonistic effects between different datasets.
-This is further demonstrated in Sec.
-3 (Fig.
-1) that certain domain mixtures lead to catastrophic performance degradation (e.g.
-Intra-D (0-20) in MATH with Qwen2).
-This contradicts the above I.I.D.
-hypothesis, suggesting that the simplified and conventional formulation in Eq (6) fails to capture cross-domain interference.
-4.2 Introducing Mixture of Underlying Distributions
-Inspired by the counterintuitive findings, we posit that each domain distribution $\mathcal{D}_{k}$ arises from the mixture of $M$ latent distributions $\{\tilde{\mathcal{D}}_{m}\}_{m=1}^{M}$ representing foundational LLM capabilities ($M\ll K$ in practice).
-This leads to our core statistical assumption:
-Assumption 4.1 (Latent Capability Structure).
-For each observed domain $k\in\mathcal{K}$, its data distribution decomposes into $M$ latent capability distributions:
-$\small\mathcal{D}_{k}=\sum_{m=1}^{M}\pi_{km}\tilde{\mathcal{D}}_{m},\quad\sum_{m=1}^{M}\pi_{km}=1,$ (7)
-where $\tilde{\mathcal{D}}_{m}$ indicates the $m$-th foundational capability. The weights $\pi_{km}$ reflect domain-specific capaility composition.
-To analyze how data optimization interacts with latent capability distributions, we further posit that each foundational capability admits an optimal configuration:
-Assumption 4.2 (Capability Optimality).
-Each foundational capability admits a unique optimal predictor:
-$\small\exists h_{\theta_{m}^{*}}\in\mathcal{H}\text{ s.t. }\theta_{m}^{*}=\mathop{\mathrm{argmin}}_{\theta}\mathbb{E}_{(x,y)\sim\tilde{\mathcal{D}}_{m}}[l(h_{\theta}(x),y)].$ (8)
-The loss $l$ is strongly convex, holding for cross-entropy or MSE, where $h_{\theta_{m}}$, termed the component model, represents a model’s manifestation of specific foundational capabilities.
-This leads to our key result formalizing how real-world training data activates latent capabilities via component models:
-Proposition 4.3.
-Under Assumption 4.1-4.2, the minimizer of the cross-domain average risk
-$\small\min_{\{\theta_{m}\}}\mathbb{E}_{k\sim\mathcal{K}}\left[\mathcal{L}_{\mathcal{D}_{k}}\left(\textstyle\sum_{m=1}^{M}\pi_{km}h_{\theta_{m}}\right)\right]$ (9)
-admits a solution $\{\theta_{m}^{*}\}_{m=1}^{M}$ where each $h_{\theta_{m}^{*}}$ optimizes the corresponding foundational capability $\tilde{\mathcal{D}}_{m}$.
-Remark Proposition 4.3 reveals an underlying fact: optimal or suboptimal solutions can be achieved through data selection to enhance the overall model performance.
-This is further validated by the role of the coefficients $\pi_{km}$, which act as capability selectors to configure domain-specific behavior, demonstrating the impact of data composition.
-4.3 Diversity Influence on the Optimization
-Building upon Proposition 4.3, we analyze the intrinsic relationship between diversity and capability composition.
-The mixture coefficients $\pi_{km}$ inherently govern both inter- and intra-domain diversity formulated in Sec.
-3.2, differing in their geometric properties in the latent capability space.
-Centroids with Coefficients Let $\mathcal{C}_{m}\in\mathbb{R}^{d}$ denote the centroid vector of latent capability $\tilde{\mathcal{D}}_{m}$ in the embedding space.
-The domain centroid $\mathcal{C}_{k}$ can be expressed as:
-$\small\mathcal{C}_{k}=\sum_{m=1}^{M}\pi_{km}\mathcal{C}_{m}.$ (10)
-This linear combination implies that the inter- and intra-diversity are determined by $\pi_{km}$ configurations:
-Proposition 4.4 (Diversity Decomposition).
-The inter-diversity metric decomposes into:
-$\small\Phi_{\text{inter}}=\sum_{m=1}^{M}\sum_{n=1}^{M}\lambda_{m}\lambda_{n}\|\mathcal{C}_{m}-\mathcal{C}_{n}\|_{2},$ (11)
-where $\lambda_{m}=\mathbb{E}_{k}[\pi_{km}]$. And the intra-diversity satisfies:
-$\small\Phi_{\text{intra}}^{(k)}\leq\sum_{m=1}^{M}\pi_{km}\|\mathcal{C}_{m}-\mathcal{C}_{k}\|_{2}^{2}+\mathbb{E}_{m\sim\pi_{k}}[\text{Var}(\tilde{\mathcal{D}}_{m})].$ (12)
-The inter-diversity result follows from substituting $\mathcal{C}_{k}=\sum\pi_{km}\mathcal{C}_{m}$ into $\mathbb{E}_{k\neq l}\|\mathcal{C}_{k}-\mathcal{C}_{l}\|$ and applying Jensen’s inequality.
-The intra-diversity bound combines the variance decomposition within each latent capability.
-Full proof is provided in Appendix C.2.
-4.4 Theoretical Insights
-Collectively, results in this section provide a unified framework to explain the varied patterns and existence of peak performance observed in Sec.
-3.3, providing theoretical insights and guidance for designing our following algorithm dealing with undetermined data.
-Diversity’s Influence on LLMs’ Capability: Based on Proposition 4.4, we conjecture that the Inter-/Intra-Diversity, which can be explicitly modeled by $\pi$, also governs $\pi$-variation control for component model mixing and thus the final overall performance.
-Proposition 4.3 further suggests that improper diversity levels may induce model suboptimality by constraining overall capability, consistent with the distribution-dependent empirical results in Sec.
-3.
-Robust Diversity Selection Strategies: Based on Proposition 4.4, it can be observed that averaging the $\pi_{km}$ helps to mitigate the missing of underlying optimal predictor $\theta_{m}^{*}$.
-From this view, the random baseline can gain relatively stable improvements in expectation.
-A more robust yet simple strategy can be derived: ensemble across candidate models trained on multiple data pools with distribution-varied inter- and intra-diversity, e.g., via techniques like voting or model weights averaging.
-This can be corroborated by results in Table 1, such as the case for Llama3.1-8B.
-Challenges in Undetermined-Domain Scenarios: However, all the selection strategies discussed so far face intractable solutions in undetermined domains with explicit domain labels, as Proposition 4.4 technically revealed: (1) ambiguous domain boundaries, e.g., blended instruction types; (2) partial overlaps in latent capabilities; and (3) absence of centroid priors $\{\mathcal{C}_{m}\}$.
-These intrinsic constraints motivate our model-aware diversity reward in the following Sec.
-5, which harnesses domain-diversity awareness into reward modeling without requiring explicit parameterization for such structural assumptions.
-5 DaaR: Diversity as a Reward
-To address the challenges identified and leverage the insights gained in previous Sec.
-3 and Sec.
-4, we establish a data selection method DaaR guided by diversity-aware reward signals.
-It comprises three key components illustrated in subsequent sections: (1) model-aware centroid synthesis, (2) two-stage training with reward probe, and (3) diversity-driven data selection.
-5.1 Model-Aware Training Data
-Model-Aware Centroid Construction
-The proposed method initiates with centroid self-synthesis through a two-phase generation process to address two fundamental challenges: (1) eliminating dependency on human annotations through automated domain prototyping, and (2) capturing the base model’s intrinsic feature space geometry for model-aware domain separation.
-  • Phase 1 - Seed Generation: For each domain $k$, generate 5 seed samples $\mathcal{S}_{k}^{(0)}$ via zero-shot prompting with domain-specific description templates, establishing initial semantic anchors.
-The prompt details and ablation of choices on sample number are provided in Appendix B.7.
-  • Phase 2 - Diversity Augmentation: Iteratively expand $S_{k}^{(t)}$ through context-aware generation, conditioned on a sliding window buffer with 3 random anchors sampled from the $(t-1)$ iteration $S_{k}^{(t-1)}$.
-The generated sample $\mathbf{x^{\prime}}$ will be admitted during the iteration via geometric rejection sampling:
-    $\max_{\mathbf{x}\in S_{k}^{(t-1)}}\cos(\mathcal{M}_{\text{ebd}}(\mathbf{x^{\prime}}),\mathcal{M}_{\text{ebd}}(\mathbf{x}))<\tau=0.85,$ (13)
-    where $\mathcal{M}_{\text{ebd}}(\cdot)$ indicates the embedding layer outputs of the given LLM.
-This process terminates when $|S_{k}|=30$.
-More analysis and ablation on these choices are provided in the Appendix.
-In Appendix D.3, we found that the adopted number of samples sufficiently leads to a stable convergence and larger data quantity does not impact the final centroids.
-In Appendix D.4, we found that this synthetic data has the ability to produce domain-representative data with clear distinction.
-Interestingly, the generated content consistently exhibits the greatest divergence between common sense, reasoning, and coding domains across model architectures and parameters.
-The domain centroid is then computed from the final augmented set $\mathcal{S}_{k}$ using the model’s knowledge prior:
-$\mathcal{C}_{k}=\frac{1}{|\mathcal{S}_{k}|}\sum_{\mathbf{x}_{i}\in\mathcal{S}_{k}}\mathcal{M}_{\text{ebd}}(\mathbf{x}_{i}).$ (14)
-This captures the LLM’s intrinsic feature space geometry while eliminating dependency on human annotations.
-Domain-Aware Clustering
-We then automatically construct pseudo-labels for the given data samples based on the previously synthesized centroids $\{C_{k}\}_{k=1}^{K}$.
-Take them as fixed prototypes, we perform constrained k-means clustering in the embedding space:
-$\arg\min_{\{S_{k}\}}\sum_{k=1}^{K}\sum_{\tilde{x}\in S_{k}}\|\mathcal{M}_{\text{ebd}}(\tilde{x})-\mathcal{C}_{k}\|^{2}.$ (15)
-This produces the seed dataset $\mathcal{D}_{\text{probe}}$ containing pseudo-labels $\{\tilde{y}_{i}\}_{i=1}$ where $\tilde{y}_{i}\in\{1,\ldots,K\}$, with model-induced and embedding-derived domain label assignments.
-5.2 Training for Self-Rewarding Abilities
-Stage 1: Domain Discrimination
-The proposed DaaR then establishes model-aware domain discrimination abilities through a multi-layer perceptron (MLP) probe module, $\psi_{\text{dom}}$, attached to the layer-3 of the LLMs $\mathcal{M}_{3}(\tilde{x})$.
-The probe will be trained meanwhile all the parameters of the LLM are frozen.
-This achieves a preferable balance between effectiveness and cost, with detailed analysis regarding the choice of Layer-3 presented in Appendix D.5.
-Specifically, with pseudo-label $\tilde{y}$, we can compute domain probabilities as:
-$p_{k}(\tilde{x})=\text{softmax}\left(\psi_{\text{dom}}(\mathcal{M}_{3}(\tilde{x}))\right),\quad\psi_{\text{dom}}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{K},$ (16)
-where $\psi_{\text{dom}}$ is optimized via cross-entropy loss function:
-$\mathcal{L}_{\text{dom}}=-\frac{1}{|\mathcal{D}_{\text{probe}}|}\sum_{(\tilde{x},\tilde{y})\in\mathcal{D}_{\text{probe}}}\sum_{k=1}^{K}\mathbb{I}_{[k=\tilde{y}]}\log p_{k}(\tilde{x}),$ (17)
-where $\mathbb{I}_{[k=\tilde{y}]}$ denotes the indicator function.
-We employ single-sample batches with the AdamW optimizer to prevent gradient averaging across domains.
-Training consistently converges and achieves 92.7% validation accuracy on domain classification, as shown in Fig.
-2 (a).
-Figure 2: Training loss and validation process of the two stages of DaaR on Qwen2-7B, detailed results are in Appendix D.6.
-Stage 2: Diversity Rewarding
-Building on the stabilized domain probe module, we quantify sample-level diversity through predictive entropy:
-$H(\tilde{x})=-\sum_{k=1}^{K}p_{k}(\tilde{x})\log p_{k}(\tilde{x}).$ (18)
-To enable efficient reward computation during data selection, we then train another 5-layer MLP $\psi_{\text{div}}$ to directly estimate $H(\tilde{x})$ from $\mathcal{M}_{3}(\tilde{x})$:
-$\hat{H}(\tilde{x})=\psi_{\text{div}}(\mathcal{M}_{3}(\tilde{x})),\quad\psi_{\text{div}}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{+}.$ (19)
-This diversity probe module $\psi_{\text{div}}$ shares $\psi_{\text{dom}}$’s architecture up to its final layer (replaced with single-output regression head), trained using entropy-scaled MSE:
-$\mathcal{L}_{\text{div}}=\frac{1}{|\mathcal{D}_{\text{probe}}|}\sum_{\tilde{x}\in\mathcal{D}_{\text{probe}}}\left(\hat{H}(\tilde{x})-H(\tilde{x})\right)^{2}.$ (20)
-This module is also well-converged as shown in Fig. 2 (b).
-Table 2: Evaluation results of Llama3.1-8B, Qwen2-7B, and Qwen2.5-7B across various downstream task benchmarks. DaaR demonstrates superiority in average performance (AVG) compared to other baselines.
-Models Distribution Common Sense Reasoning Mathematics Coding Avg
-NQ TriviaQA Hellaswag GSM8K MATH MBPP HumanEval
-Llama3.1-8B Raw 14.13 65.90 74.62 54.80 7.90 5.00 28.66 35.86
-Random 21.99 64.83 74.72 55.70 14.50 5.10 24.09 37.27
-Instruction Len 15.34 63.60 73.73 54.00 15.40 3.60 30.80 36.64
-Alpagasus (Chen et al., 2024d) 21.57 64.37 74.87 55.20 17.65 4.60 16.16 36.34
-Instag-Best (Lu et al., 2024) 18.12 64.96 74.01 55.70 15.50 4.80 37.81 38.70
-SuperFilter (Li et al., 2024a) 22.95 64.99 76.39 57.60 6.05 2.60 40.55 38.73
-Deita-Best (Liu et al., 2024b) 15.58 64.97 74.21 55.00 13.05 4.60 34.46 37.41
-DaaR (Ours) 20.08 64.55 74.88 54.8 15.30 4.70 37.50 38.83
-Qwen2-7B Raw 8.03 59.58 73.00 78.00 5.70 5.00 60.98 41.47
-Random 13.28 58.27 73.00 75.35 35.36 52.20 63.72 53.02
+Full (40K) 15.61 58.75 72.51 73.80 31.30 51.70 67.38 53.01
+Random (8K) 13.28 58.27 73.00 75.35 35.36 52.20 63.72 53.02
 Instruction Len 8.62 58.44 72.86 73.30 27.05 53.10 63.72 51.01
-Alpagasus (Chen et al., 2024d) 13.67 57.94 73.04 73.90 32.30 51.40 63.41 52.24
-Instag-Best (Lu et al., 2024) 9.51 58.50 73.06 74.70 35.35 51.90 64.70 52.53
-SuperFilter (Li et al., 2024a) 19.16 58.98 72.99 73.70 30.10 52.40 58.85 52.31
-Deita-Best (Liu et al., 2024b) 16.41 57.80 72.70 76.10 29.05 52.40 64.63 52.73
+Alpagasus chen2024alpagasus 13.67 57.94 73.04 73.90 32.30 51.40 63.41 52.24
+Instag-Best lu2023instag 9.51 58.50 73.06 74.70 35.35 51.90 64.70 52.53
+SuperFilter li2024super 19.16 58.98 72.99 73.70 30.10 52.40 58.85 52.31
+Deita-Best liu2023deita 16.41 57.80 72.70 76.10 29.05 52.40 64.63 52.73
 DaaR (Ours) 16.88 57.58 73.03 75.40 38.1 52.00 64.94 53.99
 Qwen2.5-7B Raw 8.84 58.14 72.75 78.20 9.10 7.40 78.05 44.64
-Random 11.46 57.85 73.08 78.90 13.15 62.50 71.65 52.65
+Full (40K) 12.88 58.60 72.28 76.80 13.60 62.80 71.04 52.57
+Random (8K) 11.46 57.85 73.08 78.90 13.15 62.50 71.65 52.65
 Instruction Len 11.34 58.01 72.79 78.00 15.80 62.30 68.12 52.34
-Alpagasus (Chen et al., 2024d) 10.40 57.87 72.92 77.20 18.75 61.80 65.55 52.07
-Instag-Best (Lu et al., 2024) 11.08 58.40 72.79 76.40 16.40 62.90 70.43 52.63
-SuperFilter (Li et al., 2024a) 13.54 58.51 72.89 79.30 11.35 39.50 65.25 48.62
-Deita-Best (Liu et al., 2024b) 10.50 58.17 73.14 74.60 16.60 62.00 72.26 52.47
+Alpagasus chen2024alpagasus 10.40 57.87 72.92 77.20 18.75 61.80 65.55 52.07
+Instag-Best lu2023instag 11.08 58.40 72.79 76.40 16.40 62.90 70.43 52.63
+SuperFilter li2024super 13.54 58.51 72.89 79.30 11.35 39.50 65.25 48.62
+Deita-Best liu2023deita 10.50 58.17 73.14 74.60 16.60 62.00 72.26 52.47
 DaaR (Ours) 15.83 58.65 72.48 80.20 16.70 64.20 68.29 53.76
-Data Selection: After training the module $\psi_{\text{div}}$, we can use its output to select data samples.
-Building on the theoretical insights in Sec.
-4.4, data points that are closer to other centroids and more dispersed within their own centroid are more beneficial for enhancing the comprehensive capabilities of the model.
-Therefore, we use the predicted entropy-diversity score as a reward, selecting the top 20% with the highest scores as the final data subset for fine-tuning.
-5.3 Empirical Validation and Main Results
-To validate the efficacy of DaaR, we conduct experiments comparing more SOTA methods on the data pools in Sec.
-3.1 with critical modifications: all domain-specific labels are deliberately stripped.
-This constraint mimics more challenging real-world scenarios and precludes direct comparison with data mixture methods requiring domain label prior.
-Baselines
-We use the following data selection methods for comprehensive evaluation: (1) Random Selection: traditional random sampling; (2) Instruction Len: measuring instruction complexity by token count  (Cao et al., 2023); (3) Alpagasus (Chen et al., 2024d): using ChatGPT for direct quality scoring of instruction pairs; (4-5) Instag (Lu et al., 2024): semantic analysis approach with Instag-C (complexity scoring via tag quantity) and Instag-D (diversity measurement through tag set expansion); (6) SuperFilter (Li et al., 2024a): response-loss-based complexity estimation using compact models; (7-9) Deita (Liu et al., 2024b): model-driven evaluation with Deita-C (complexity scoring), Deita-Q (quality scoring), and Deita-D (diversity-aware selection).
-Detailed configurations for these baselines are in Appendix B.4.
-5.3.1 Overall Performance
-The experimental results are presented in Table 2, where Instag-Best and Deita-Best represent the optimal variants from their respective method families.
-Our experiments clearly demonstrate the effectiveness of DaaR across three major language models and seven challenging benchmarks.
-A detailed analysis is provided below.
-High-Difficulty Scenario: The task of balanced capability enhancement proves particularly challenging for existing methods.
+5.1 Overall Performance
+The main experimental results are presented in Table 2, where Instag-Best and Deita-Best represent the optimal variants from their method families.
+Our experiments clearly demonstrate the effectiveness of DaaR across three major language models and seven challenging benchmarks:
+High-Difficulty Scenario
+: The task of balanced capability enhancement proves particularly challenging for existing methods.
 While some baselines achieve strong performance on specific tasks (e.g., SuperFilter’s 40.55 on HumanEval for Llama3.1), they suffer from catastrophic performance drops in other domains (e.g., SuperFilter’s 6.05 on MATH).
-Only three baseline methods perform better than Random selection, with notably severe degradation applied in Qwen-series models, all of which fall below the Random performance.
-We hypothesize this stems from over-specialization – excessive focus on narrow capability peaks at the expense of broad competence (visualized in Appendix B.5).
-Our method’s robustness stems from preventing extreme distribution shifts through diversity constraints.
-Universal Superiority: DaaR establishes new SOTA averages across all models, surpassing the best baselines by +0.14 (Llama3.1), +0.97 (Qwen2), and +1.11 (Qwen2.5).
+Only 3 baselines perform better than Random on Llama3.1-8B.
+In the Qwen series, even all baselines fall below the Random performance.
+We conjecture this stems from selected distribution-bias, as the baselines are unable to balance the proportions across different domains, a perspective visually supported and analyzed in Appendix D.2.
+Comprehensive Performance
+: DaaR establishes new SOTA averages across all models, surpassing the best baselines by +0.14 (Llama3.1), +0.97 (Qwen2), and +1.11 (Qwen2.5).
 The proposed method uniquely achieves dual optimization in critical capabilities: Mathematical Reasoning: Scores 38.1 MATH (Qwen2) and 16.70 MATH (Qwen2.5), with 7.4% and 27.0% higher than respective random baselines.
 Coding Proficiency: Maintains 64.94 HumanEval (Qwen2) and 64.20 MBPP (Qwen2.5) accuracy with <1% degradation from peak performance.
-This demonstrates DaaR’s ability to enhance challenging STEM capabilities while preserving core competencies, a critical advancement over specialized but unstable baselines.
-Cost-Efficiency and Flexibility: Compared to baseline methods requiring GPT-based evaluators (Alpagasus, Instag) or full LLaMA-7B inference (Deita), DaaR achieves superior efficiency through data-model co-optimizations.
-Our method makes the LLM ability of self-rewarding from dedicated data synthesis, and operates on frozen embeddings from layer 3 (vs.
-full 32-layer inference in comparable methods), reducing computational overhead while maintaining capability integrity.
-The lightweight 5-layer MLP probe module requires only 9GB GPU memory during training (vs.
-18$\sim$24GB for LLM-based evaluators) and adds merely 76M parameters.
-This plug-and-play feature enables seamless integration of DaaR with existing LLMs without additional dependency management.
-6 Conclusion and Future Work
+This demonstrates DaaR’s ability to enhance challenging STEM capabilities while preserving core competencies.
+5.2 Generalizability and Uniqueness of DaaR
+Evaluation on MMLU In the main experiments, the data and benchmarks we employed are meticulously aligned, which validates the In-Distribution capabilities of DaaR.
+To further validate the generalization ability of DaaR, we present MMLU mmlu2021 evaluations in Appendix F.1.
+MMLU encompasses a significantly broader range of content than our dataset, thus serving as an Out-of-Distribution scenario.
+Results demonstrate that DaaR achieved top accuracy on Qwen2 with an improvement of 0.61 points, and secured a competitive second-place performance on Qwen2.5.
+At a granular level, DaaR particularly excels in the STEM domain, outperforming baselines by an average of 7.6% and 1.7% on Qwen2 and Qwen2.5, which aligns the findings above.
+Evaluation on Qwen3 In Appendix F.2, we explore the performance of DaaR on the mainstream RL-based LLM, Qwen3-8B.
+We observe that despite difference in structure, the phenomena (Fig.
+8 & Table 8) and the dynamics of training (Fig.
+9) remain consistent with the main observations.
+Then we compare DaaR with baselines in Table 9, although Qwen3-8B is not directly designed for SFT paradigm, DaaR maintains the top position in Avg performance across baselines, surpasses all baselines up to 4.7% and notably exceeds the avg of baselines by 25.91% on MATH.
+Customized Characteristics Unlike baselines, DaaR enables domain-aware data selection in undetermined scenario, showcasing potential customization on $\lambda_{k}$.
+As detailed in Appendix F.3, by adjusting the selection ratio for specific domains, we achieve SOTA performance in targeted domain without compromising overall ability, promising practical utility in real-world application.
+Table 3: Illustration results on MMLU, Qwen3 and customization, detailed in Table 6, 9, 10
+(a) MMLU Performance
+Qwen2-7B MMLU-Avg
+Raw 28.98
+Rand (8K) 68.81
+Instruction-L 68.55
+Alpagasus 59.34
+Instag-Best 68.65
+SuperFilter 54.27
+Deita-Best 68.08
+DaaR (Ours) 69.42
+(b) Eval on Qwen3
+Qwen3-8B Avg
+Raw 49.41
+Rand (8K) 48.16
+Instruction-L 49.19
+Alpagasus 47.81
+Instag-Best 48.69
+SuperFilter 49.51
+Deita-Best 49.90
+DaaR (Ours) 50.06
+(c) Customized Selection on Domain
+Qwen2.5 NQ Hellaswag Avg
+Raw 8.84 72.75 44.64
+Rand 11.46 73.08 52.66
+Instag 11.08 72.79 52.63
+DaaR 15.83 72.48 53.76
+Common 17.56 72.51 52.85
+Reason 16.56 74.62 53.39
+5.3 Ablation Studies and Discussions
+Distinct Domain Centroid As shown in Appendix G.1, DaaR can generate domain-representative samples with clear distinctions.
+Notably, the diversity is consistently most pronounced between common sense, reasoning, and coding domains across model architectures and parameters.
+Impacts of Hyper-Parameters In Appendix G.2-Appendix G.6, we conduct a series of ablation studies, including (1) selection of layer: the discrimination accuracy derived from layer 1-10 fluctuates in 91.2-93.6%; (2) the number of seed: the similarity between centroids remains highly consistent with variance ranging from 5.4E-7 to 1.8E-5; (3) the number of augmented data: the mean similarity of centroids generated from sizes of 10 and 30 is 0.983 and 0.981 on Qwen2.5 and Llama3.1 respectively; (4) the size of sliding window & similarity threshold: the impact of different hyper-parameters on centroid generation is minimal, ensuring the subsequent training stability of DaaR.
+Cost-Efficiency and Flexibility As shown in Appendix G.7, our method demonstrates significant computational efficiency compared to baselines that rely on GPT-based evaluators or full-LLM inference.
+By operating on frozen hidden layers and generating regression scores instead of full vocabulary distributions, our approach achieves 70% lower GPU usage and 2.5x faster inference.
+6 Conclusions, Limitations & Future Works
 In this paper, we propose a new approach to fine-tuning LLMs with data that lacks clear domain labels, using diversity as a guiding principle.
-Our method allows LLMs to automatically select and benefit from diverse datasets.
 By measuring semantic diversity with entropy, we employ a self-reward mechanism built upon the given LLM, identifying data that best fits the model’s natural tendencies in terms of its underlying knowledge distribution.
-Our experiments with various SOTA LLMs show notable superiority of the method over SOTA methods, highlighting the potential of data diversity to enhance model overall performance.
-This research demonstrates feasibility of reward modeling on LLM data diversity and deepens the understanding of its utilities, especially when domain labels are missing or uncertain.
-Future work could include developing methods to efficiently adjust diversity measures and learning algorithm, thus supporting LLMs in self-evolving towards artificial general intelligence in dynamic environments.
+Our experiments with various SOTA LLMs show notable superiority of the method over competitive baselines, highlighting the potential of data diversity to enhance model overall performance.
+There are several limitations to our study.
+Regarding the implementation of DaaR, the customization of selection ratios holds potential for broader application in real-world scenarios.
+Future work could include developing methods to efficiently adjust diversity measures and learning algorithms, thus supporting LLMs in self-evolving towards artificial general intelligence in dynamic environments.
 Appendix
@@
   • Sec. A: The symbols used in this paper.
+  • Sec. B: The details about experimental setup, including data pool construction (Sec. B.1), implementation of embedding (Sec. B.2), benchmarks (Sec. B.3) and train / eval details (Sec. B.4).
+  • Sec. C: The implementation details of DaaR, including hyper-parameter setting (Sec. C.1) and the process of model-aware training data generation (Sec. C.2).
+  • Sec. D: The details of utilized baselines, including the implementation of baselines (Sec. D.1) and visualization of the data selected by baselines (Sec. D.2).
+  • Sec. E: The detailed derivation of Proposition 3.2.
   • Sec.
-B: The details about experiments and DaaR implementation, including the construction of data pools (Sec.
-B.1), the benchmarks (Sec.
-B.2), the training & evaluation implementation (Sec.
-B.3), the implementation of baselines (Sec.
-B.4), the visualization of baseline’s selected data samples (Sec.
-B.5), the implementation of the synthetic data generation (Sec.
-B.6), the implementation of seed generation & clustering (Sec.
-B.7), and the prompts for the generation stage (Sec.
-B.8).
-  • Sec C: The full proof for theoretical results are presented in Section 4.
-  • Sec D: The more experimental results, including the visualization results of embeddings on Llama3.1-8B & Qwen2.5-7B (Sec D.1), the visualization results of inter- & intra-diversity distributions (Sec D.2), the similarity analysis in generated centroid quantities (Sec.
-D.3), the similarity analysis in generated centroid domains (Sec.
-D.4), the DaaR layer selection protocol (Sec.
-D.5), the comprehensive training dynamics (Sec.
-D.6), the validation of inter-diversity and intra-diversity (Sec.
-D.7) and the complete results of DaaR with baselines on all LLMs (Sec.
-D.8).
+F: More comprehensive experiments about DaaR, including the evaluation on MMLU (Sec.
+F.1), the validation on Qwen3 (Sec.
+F.2), the customized selection (Sec.
+F.3) and the utilization of entropy (Sec.
+F.4).
+  • Sec.
+G: Complete ablation studies and discussions of DaaR, including the centroid similarity (Sec.
+G.1), LLM’s layer selection (Sec.
+G.2).
+As well as several ablation studies of the number of seeds (Sec.
+G.3), the number of diversity augmentation data (Sec.
+G.4), the size of sliding window (Sec.
+G.5) and diversity threshold (Sec.
+G.6).
+Additionally, we verify the cost-efficiency (Sec.
+G.7) and the stability of DaaR (Sec.
+G.8).
+  • Sec H: The comprehensive experimental results, including the visualization results of embeddings on Qwen2-7B, Llama3.1-8B & Qwen2.5-7B (Sec H.1), the visualization of inter- & intra-diversity distributions (Sec H.2), the comprehensive training dynamics (Sec.
+H.3), the validation of inter-diversity and intra-diversity (Sec.
+H.4) and complete results of DaaR with baselines (Sec.
+H.5).
 Appendix A Notation
-For ease of reading and reference, we present all mathematical symbols used in this paper in Table 3.
+For ease of reading and reference, we present the mathematical symbols used in this paper in Table 4.
 Table 4: Symbol Notation
@@
 $\phi_{\text{intra}}$ Sample similarity to its domain centroid $\Phi_{\text{intra}}^{(k)}$ Domain-level variance of domain-$k$
-$\mathcal{K}$ Set of labeled domains $h_{k}$ Independent model of domain $k$
-$l$ Loss function $\mathcal{L}_{\mathcal{D}_{k}}$ Empirical risk
-$\tilde{\mathcal{D}}_{m}$ Latent distribution of foundational capabilities $\pi_{km}$ Mixture weights
-$h_{\theta_{m}}$ Component model $\theta_{m}^{*}$ Optimal predictor
+$l$ Loss function $p,q$ Data distributions
 $\mathcal{S}_{k}^{(0)}$ Generated centroids’ seed sample $\mathcal{S}_{k}^{(t)}$ Generated centroid sample at $t$-th round
-$\mathcal{M}_{\text{ebd}}$ Embedding layer of LLM $\mathcal{M}_{3}$ Layer-3 of LLM
+$\mathcal{M}_{\text{ebd}}$ Embedding layer of LLM $\mathcal{M}$ Hidden layer of LLM
 $\mathcal{D}_{\text{probe}}$ DaaR training set $\tau$ A similarity threshold
 $\tilde{x}_{i}$ Data sample in $\mathcal{D}_{\text{probe}}$ $\tilde{y}_{i}$ Pseudo-label of $\tilde{x}_{i}$
-$\psi_{\text{dom}}$ A MLP-based domain classifier $\psi_{\text{div}}$ A MLP-based entropy predictor
+$\psi_{\text{dom}}$ A MLP-based domain predictor $\psi_{\text{div}}$ A MLP-based entropy predictor
 $p_{k}(\tilde{x})$ Domain predicted probability by $\psi_{\text{dom}}$ $\hat{H}(\mathbf{x})$ Predictive entropy by $\psi_{\text{dom}}$
 $\mathcal{L}_{\text{dom}}$ Cross-entropy loss of $\psi_{\text{dom}}$ $\mathcal{L}_{\text{div}}$ MSE loss of $\psi_{\text{dom}}$
-Appendix B Experiments and Method Details
+Appendix B Experimental Setup Details
 B.1 Construction of Data Pool
@@
 We select the following datasets to evaluate specific abilities:
-  • Common Sense: Dolly-15K (Conover et al., 2023) with 15,011 samples, an open source dataset of instruction-following records generated by thousands of Databricks employees in several of the behavioral categories.
-  • Reasoning: Cot-en (Chung et al., 2024) with 74,771 samples, is created by formatting and combining nine CoT datasets released by FLAN.
-  • Mathematics: Math-Instruct (Yue et al., 2024) with 262,039 samples, is compiled from 13 math rationale datasets, six of which are newly curated by this work.
+  • Common Sense: Dolly-15K [14] with 15,011 samples, an open source dataset of instruction-following records generated by thousands of Databricks employees in several of the behavioral categories.
+  • Reasoning: Cot-en [12] with 74,771 samples, is created by formatting and combining nine CoT datasets released by FLAN.
+  • Mathematics: Math-Instruct [49] with 262,039 samples, is compiled from 13 math rationale datasets, six of which are newly curated by this work.
 It uniquely ensures extensive coverage of diverse mathematical fields.
-  • Coding: Code-Alpaca (Chaudhary, 2023) with 20,016 samples, is constructed from real-world code examples, providing a rich set of tasks designed to guide models in generating accurate and functional code.
+  • Coding: Code-Alpaca [6] with 20,016 samples, is constructed from real-world code examples, providing a rich set of tasks designed to guide models in generating accurate and functional code.
 Each dataset was initially filtered and randomly reduced to 10,000 entries, resulting in a combined data pool of 40,000 entries.
 Specifically, for the Math-Instruct dataset, due to its inclusion of CoT and certain coding capabilities, we extract a highly mathematics-related subset and use regular expressions to filter out the coding-related content (including ’program’, ’python’, ’def’, ’import’, ’print’, ’return’), ensuring it remains within the domain of mathematics.
-B.2 Benchmarks
-To evaluate the models’ true capabilities and performance across different domains, we follow the approach of major open-source LLMs (e.g., the Llama3 series (Dubey et al., 2024) and Qwen2 series (Yang et al., 2024)) and select the following widely used evaluation sets.
+B.2 Implementation of Embedding
+Given the Alpaca-based [38] data format employed in our study, we concatenate the ’instruction’, ’input’, and ’output’ components of each data sample to capture domain-aware semantic representations.
+These concatenated sequences are then fed into the LLM, where we compute the average of all token embeddings extracted from the embedding layer to derive the semantic vector characterizing each sample.
+The embedding layer is deliberately selected due to its shallow architectural position: (1) it preserves precise semantic information capture through raw input representations, and (2) it incurs reduced computational overhead compared to deeper layers.
+As evidenced by the visualization in Fig.
+12 (Appendix G.2), this approach demonstrates robust domain discrimination capabilities across different LLM layers, thereby validating the robustness of our methodology.
+B.3 Benchmarks
+To evaluate the models’ true capabilities and performance across different domains, we follow the approach of major open-source LLMs (e.g., the Llama3 series [17] and Qwen2 series [46]) and select the following widely used evaluation sets.
 All evaluations were conducted on the OpenCompass platform [^1] https://opencompass.org.cn/.
-  • Common Sense: NQ (Kwiatkowski et al., 2019) and TriviaQA (Joshi et al., 2017), which cover factual knowledge-based questions of varying difficulty.
-  • Reasoning: HellaSwag (Zellers et al., 2019), which effectively evaluates the model’s comprehensive reasoning ability.
-  • Mathematics: GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) benchmarks, which encompass problems ranging from elementary to competition-level difficulty.
-  • Coding: MBPP (Austin et al., 2021) and HumanEval (Chen et al., 2021), which include evaluations of basic coding abilities in Python.
+  • Common Sense: NQ [25] and TriviaQA [24], which cover factual knowledge-based questions of varying difficulty.
+  • Reasoning: HellaSwag [50], which effectively evaluates the model’s comprehensive reasoning ability.
+  • Mathematics: GSM8K [13] and MATH [22] benchmarks, which encompass problems ranging from elementary to competition-level difficulty.
+  • Coding: MBPP [4] and HumanEval [11], which include evaluations of basic coding abilities in Python.
 We use the average of various metrics to demonstrate the models’ overall performance across different domains.
@@
 Utilized 3,610 5,000 10,042 500 1,000 500 164
-B.3 Training and Evaluation Details
+B.4 Training and Evaluation Details
 Platform
-We implement our approaches using PyTorch (Paszke et al., 2019) v2.4.1, coupled with PEFT v0.12.0 and the Transformers library (Wolf et al., 2020) v4.45.2.
+We implement our approaches using PyTorch [31] v2.4.1, coupled with PEFT v0.12.0 and the Transformers library [42] v4.45.2.
 Experiments are conducted on a computing platform equipped with four NVIDIA A100 GPUs (40GB), with pre-trained LLMs loaded as 16-bit floating-point numbers.
-The specific data-model development processes are completed in Data-Juicer Sandbox (Chen et al., 2024c), via integration with the ms-swift (Zhao et al., 2024b) training repository, and the OpenCompass (Contributors, 2023) evaluation repository.
+The specific data-model development processes are completed in Data-Juicer Sandbox [9], via integration with the ms-swift [54] training repository, and the OpenCompass [15] evaluation repository.
 Training Details
-In our experimental setup, we employ Low-Rank Adaptation (LoRA) (Hu et al., 2021) adapters for the fine-tuning process, utilizing a LoRA-rank of 8 and a LoRA-alpha of 16.
+In our experimental setup, we employ Low-Rank Adaptation (LoRA) [23] adapters for the fine-tuning process, utilizing a LoRA-rank of 8 and a LoRA-alpha of 16.
 The learning rate was consistently maintained at $5\times 10^{-5}$ across all experiments to ensure uniformity in training dynamics.
 We utilize a batch size of 4 and set the maximum sequence length to 2048 tokens to accommodate the model’s capacity.
-To optimize the training process, a warmup ratio of 0.05 was applied and a validation ratio of 0.03 was used.
+To optimize the training process, a warmup ratio of 0.05 was applied, and a validation ratio of 0.03 was used.
 The training was conducted over a single epoch, balancing computational efficiency with the need for effective model adaptation.
-Following some effective instruction-tuning work (Zhou et al., 2024; Lu et al., 2024), we set the size of our subset to 8,000 entries, which constitutes 20% of the data pool.
+Following some effective instruction-tuning work [55, 30], we set the size of our subset to 8,000 entries, which constitutes 20% of the data pool.
 Evaluation Details
-Following the guidelines provided by OpenCompass (Contributors, 2023), we adhered to the default settings for our evaluation process.
-We select the hf-type as base and utilized a batch size of 16 to ensure efficient processing.
+Following the guidelines provided by OpenCompass [15], we adhered to the default settings for our evaluation process.
+We select the hf-type as base and utilize a batch size of 16 to ensure efficient processing.
 For most tasks, we employ the gen mode, while for the Hellaswag task, we opt for the ppl mode to better assess perplexity.
-B.4 Implementation of Baselines
+Appendix C Implementation of DaaR
+C.1 Implementation of the Synthetic Data Generation
+For the generation of synthetic data, we utilize the model.generate function from the PyTorch library.
+Key parameters included setting max-new-tokens to 2048 to control the length of the output, enabling do-sample to True to allow sampling, and configuring top-p to 0.95 and temperature to 0.9 to ensure diversity in the generated content.
+During the content extraction phase, we employ regular expressions to efficiently extract and structure the desired information.
+For implementation details of DaaR’s hyper-parameters, we set the number of seed samples $\mathcal{S}_{k}^{(0)}$ as 5 (ablation in Appendix G.3), the number of anchors in sliding window as 3 (ablation in Appendix G.5), the similarity threshold $\tau$ as 0.9 for Mathematics and 0.85 for others (ablation in Appendix G.6), the terminated iteration as 30 (ablation in Appendix G.4), the hidden layer $\mathcal{M}(\tilde{x})$ as layer-3 (ablation in Appendix G.2).
+C.2 Implementation of Model-Aware Training Data Generation
+Seed Generation
+This phase aims to provide zero-shot LLM-generated seed data that reflects how LLMs inherently perceive the characteristics of domain-specific data.
+We employ the LLM itself to generate concise domain descriptions across diverse fields illustrated in Prompt C.2, and the prompt shown in Prompt C.2.
+Due to the difficulty of pre-trained LLMs in providing accurate and coherent responses, we utilize their corresponding Instruct versions for centroid data generation.
+Specifically, we apply Llama3.1-8B-Instruct for Llama3.1-8B, Qwen2-7B-Instruct for Qwen2-7B, and Qwen2.5-7B-Instruct for Qwen2.5-7B.
+Notably, these descriptions are not required to be optimal or prescriptive, rather, they serve as representative examples compatible with our experimental scenarios.
+Users retain the flexibility to customize target domain specifications according to specific requirements.
+The associated stability analysis regarding this design choice will be elaborated in Appendix G.8.
+Diversity Augmentation
+While seed data effectively captures the intrinsic domain knowledge of LLMs, its limited quantity and the constrained diversity of single-prompt generation methods necessitate enhancement.
+We address these limitations through two complementary strategies including a sliding window mechanism and a rule-based filter.
+The implementation details are formalized in Prompt C.2 and Eq.
+(7).
+Domain-Aware Clustering
+During reward model training data clustering, we rigorously curate a subset of 5,000 entries that are distribution-matched to the fine-tuning dataset while separate from it.
+[…]
+[…]
+Reasoning: Reasoning involves the ability to think logically about a situation or problem, to draw conclusions from available information, and to apply knowledge in new situations.
+Mathematics: Mathematical skills include the ability to perform calculations, understand mathematical concepts, solve hard and professional math problems, and apply mathematical reasoning.
+Coding: Design and generate specific code programs, or apply algorithms and data structures, with code generation in the Output.
+[…]
+[…]
+Generate 5 different instruction pairs related to this field with various lengths. Maintain the index format: Instruction [1 … 5].
+The response should include three parts:
+1. Instruction: A clear command or question that can be understood by the assistant.
+2. Input: Any information provided to help it understand the instruction. If there is no need to generate, just keep it empty.
+3. Output: The expected answer or action.
+Keep the generated content focused on {selected_domain}. And do not involve {unselected_domains} related knowledge.
+[…]
+[…]
+Generate only an instruction pair related to this field. The response should include three parts:
+Instruction: A clear command or question that can be understood by the assistant.
+Input: Any information provided to help it understand the instruction. If there is no need to generate, just keep empty.
+Output: The expected answer or action.
+Keep the generated content focused on {selected_domain}. Do not involve {unselected_domain} related knowledge.
+Note that you should generate content strongly unrelated and different to these examples to ensure diversity in the generated output:
+Counterexample: {}
+The format of the generated content should be: Instruction: [], Input: [], Output: [].
+Appendix D Details of Baselines
+D.1 Implementation of Baselines
 We select the following representative methods and works related to data selection as our baselines to evaluate their performance in the context of a mixture of downstream tasks.
@@
 The length of the instruction can be considered a measure of input complexity.
-It is widely believed (Cao et al., 2023; Zhao et al., 2023) that more complex data is beneficial for enhancing model capabilities.
+It is widely believed [5, 53] that more complex data is beneficial for enhancing model capabilities.
 Therefore, we select a subset of 8,000 entries with the maximum number of words (based on spaces) in the concatenation of Instruction and Input as part of the filtering process.
-AlpaGasus (Chen et al., 2024d)
+AlpaGasus [10]
 Using prompts to directly score and annotate the quality of the data leverages the powerful cognitive abilities of LLMs for evaluation and selection.
@@
 Figure 3: The distribution of the score by Alpagasus.
-Instag (Lu et al., 2024)
+Instag [30]
 first utilizes ChatGPT to tag the samples based on semantics and intentions, then trains a LLaMA-based tagger on the ChatGPT tags to tag data.
@@
 User Prompt:
-Please provide coarse-grained tags, such as ”Spelling and Grammar Check” and ”Cosplay”, to identify main intentions of above instruction.
+Please provide coarse-grained tags, such as "Spelling and Grammar Check" and "Cosplay", to identify main intentions of above instruction.
 Your answer should be a list including titles of tags and a brief explanation of each tag.
-Your response have to strictly follow this JSON format: [”tag”: str, ”explanation”: str].
+Your response have to strictly follow this JSON format: ["tag": str, "explanation": str].
 Please response in English.
@@
 This approach captures a broader range of unique tags, thereby enhancing the diversity and representativeness of the dataset.
-SuperFilter (Li et al., 2024a)
+SuperFilter [26]
 introduces a method that astonishingly utilizes a small GPT2 model to successfully filter out the high-quality subset from the existing GPT4-generated instruction tuning dataset.
@@
 By meticulously following this methodology, we utilize GPT-2 to select 8,000 data samples.
-Deita (Liu et al., 2024b)
+Deita [29]
 is an open-sourced project designed to facilitate Automatic Data Selection for instruction tuning in LLMs.
@@
 Following its methodology, we gain 8,000 entries as a baseline.
-B.5 Visualization of Baseline’s Selected Data
+D.2 Visualization of Baseline’s Selected Data
 We project all baseline-selected data samples through Llama3.1-8B’s, Qwen2-7B’s and Qwen2.5-7B’s embedding layer and visualize their distributions via t-SNE dimensionality reduction in Fig.
 5, Fig.
-6 and Fig.
-7 respectively.
+6, and Fig.
+7, respectively.
 This visualization provides intuitive insights into how different data selection strategies handle domain-mixed data.
-Three key observations emerge:
-  • 1
-    Certain methods (notably Instruction Length, Deita-C, and Deita-Q baselines) exhibit significant distributional bias against coding-domain samples, correlating with their poor HumanEval performance in Table 10.
-  • 2
-    The visualization reveals inherent challenges in maintaining original data distributions when handling domain-undetermined samples.
-  • 3
-    Despite observable distribution shifts, not all methods show proportional performance degradation, aligning with our theoretical analysis in Section 4 about the link between latent distribution and LLM’s fundamental capabilities.
-Figure 5: t-SNE visualization of data samples selected by different baselines using Llama3.1-8B embeddings.
-While original labels were removed during training, we preserve them in the legend for interpretability.
-Figure 6: t-SNE visualization of data samples selected by different baselines using Qwen2-7B embeddings.
-While original labels were removed during training, we preserve them in the legend for interpretability.
-Figure 7: t-SNE visualization of data samples selected by different baselines using Qwen2.5-7B embeddings.
-While original labels were removed during training, we preserve them in the legend for interpretability.
-B.6 Implementation of the Synthetic Data Generation
-For the generation of synthetic data, we utilize the model.generate function from the PyTorch library.
-Key parameters included setting max-new-tokens to 2048 to control the length of the output, enabling do-sample to True to allow sampling, and configuring top-p to 0.95 and temperature to 0.9 to ensure diversity in the generated content.
-During the content extraction phase, we employ regular expressions to efficiently extract and structure the desired information.
-B.7 Implementation of Seed Generation & Clustering
-Due to the difficulty of pre-trained LLMs in providing accurate and coherent responses, we utilize their corresponding Instruct versions for centroid data generation.
-Specifically, we apply Llama3.1-8B-Instruct for Llama3.1-8B, Qwen2-7B-Instruct for Qwen2-7B, and Qwen2.5-7B-Instruct for Qwen2.5-7B.
-The domain description and the zero-shot prompt used in the Seed Generation phase are as follows.
-When clustering reward training data based on the seed, we select 5,000 entries that match the distribution of the fine-tuning data but do not overlap with it.
-This practice has proven that even a small amount of data can effectively train a reward probe module of DaaR.
-[…]
-[…]
-Reasoning: Reasoning involves the ability to think logically about a situation or problem, to draw conclusions from available information, and to apply knowledge in new situations.
-Mathematics: Mathematical skills include the ability to perform calculations, understand mathematical concepts, solve hard and professional math problems, and apply mathematical reasoning.
-Coding: Design and generate specific code programs, or apply algorithms and data structures, with code generation in the Output.
-[…]
-[…]
-Generate 5 different instruction pairs related to this field with various lengths. Maintain the index format: Instruction [1 … 5].
-The response should include three parts:
-1. Instruction: A clear command or question that can be understood by the assistant.
-2. Input: Any information provided to help it understand the instruction. If there is no need to generate, just keep it empty.
-3. Output: The expected answer or action.
-Keep the generated content focused on {selected_domain}. And do not involve {unselected_domains} related knowledge.
-B.8 Prompts for the Generation Stage
-[…]
-[…]
-Generate only an instruction pair related to this field. The response should include three parts:
-Instruction: A clear command or question that can be understood by the assistant.
-Input: Any information provided to help it understand the instruction. If there is no need to generate, just keep empty.
-Output: The expected answer or action.
-Keep the generated content focused on {selected_domain}. Do not involve {unselected_domain} related knowledge.
-Note that you should generate content strongly unrelated and different to these examples to ensure diversity in the generated output:
-Counterexample: {}
-The format of the generated content should be: Instruction: [], Input: [], Output: [].
-Appendix C Proof of Section 4
-C.1 Proof of Proposition 4.3
-Lemma C.1 (Cross-Domain Risk Decomposition).
-Under Assumption 4.1, the average cross-domain risk becomes:
-$\mathbb{E}_{k\sim\mathcal{K}}\mathcal{L}_{\mathcal{D}_{k}}\left(\textstyle\sum_{m=1}^{M}\pi_{km}h_{\theta_{m}}\right)=\sum_{m=1}^{M}\lambda_{m}\mathbb{E}_{\tilde{\mathcal{D}}_{m}}\left[l\left(\textstyle\sum_{m^{\prime}=1}^{M}\pi_{km^{\prime}}h_{\theta_{m^{\prime}}}(x),y\right)\right],$ (21)
-where $\lambda_{m}=\mathbb{E}_{k}[\pi_{km}]$ represents global activation of capability $m$.
-Proof.
-Under the finite variance condition $\mathbb{E}_{\tilde{\mathcal{D}}_{m}}[\|l\|_{2}^{2}]<\infty$, Fubini’s theorem holds, and results in:
-$\mathbb{E}_{k}\mathcal{L}_{\mathcal{D}_{k}}$ $=\mathbb{E}_{k}\mathbb{E}_{(x,y)\sim\mathcal{D}_{k}}\left[l\left(\textstyle\sum_{m}\pi_{km}h_{\theta_{m}}(x),y\right)\right]$
-$=\mathbb{E}_{k}\mathbb{E}_{m\sim\pi_{k}}\mathbb{E}_{\tilde{\mathcal{D}}_{m}}\left[l\left(\textstyle\sum_{m^{\prime}}\pi_{km^{\prime}}h_{\theta_{m^{\prime}}}(x),y\right)\right]$
-$=\sum_{m=1}^{M}\lambda_{m}\mathbb{E}_{\tilde{\mathcal{D}}_{m}}\left[l\left(\textstyle\sum_{m^{\prime}}\pi_{km^{\prime}}h_{\theta_{m^{\prime}}}(x),y\right)\right]$ (22)
-Theorem C.2 (Global Optimality Condition).
-Under Assumptions 4.1-4.2, the optimal predictors $\{h_{\theta_{m}^{*}}\}$ minimize the cross-domain risk when:
-$\forall m,\ \theta_{m}^{*}=\mathop{\mathrm{argmin}}_{\theta}\mathbb{E}_{\tilde{\mathcal{D}}_{m}}[l(h_{\theta}(x),y)].$ (23)
-Proof.
-Taking derivatives of Lemma C.1’s right-hand side:
-$\frac{\partial}{\partial\theta_{m}}\sum_{m^{\prime}=1}^{M}\lambda_{m^{\prime}}\mathbb{E}_{\tilde{\mathcal{D}}_{m^{\prime}}}[\cdot]=\lambda_{m}\mathbb{E}_{\tilde{\mathcal{D}}_{m}}\left[\frac{\partial l}{\partial h}\frac{\partial h_{\theta_{m}}}{\partial\theta_{m}}\right]+\sum_{m^{\prime}\neq m}\lambda_{m^{\prime}}\mathbb{E}_{\tilde{\mathcal{D}}_{m^{\prime}}}\left[\frac{\partial l}{\partial h}\frac{\partial h_{\theta_{m}}}{\partial\theta_{m}}\right].$
-By strong convexity and optimality of $\theta_{m}^{*}$, all terms vanish when $\theta_{m}=\theta_{m}^{*}$. ∎
-C.2 Proof of Proposition 4.4
-Proof of Inter-Diversity
-To prove the inter-diversity decomposition as given in Proposition 4.4, we begin by considering the inter-diversity metric $\Phi_{\text{inter}}$, defined as the expected pairwise distance between domain centroids from Eq (3):
-$\Phi_{\text{inter}}=\mathbb{E}_{k\neq l}\left[\|\mathcal{C}_{k}-\mathcal{C}_{l}\|_{2}\right].$ (24)
-Substituting the expression for each domain centroid (Eq (10)) $\mathcal{C}_{k}=\sum_{m=1}^{M}\pi_{km}\mathcal{C}_{m}$, we have:
-$\|\mathcal{C}_{k}-\mathcal{C}_{l}\|_{2}=\left\|\sum_{m=1}^{M}\pi_{km}\mathcal{C}_{m}-\sum_{n=1}^{M}\pi_{ln}\mathcal{C}_{n}\right\|_{2}.$ (25)
-Applying the triangle inequality and expanding, the distance becomes:
-$=\left\|\sum_{m=1}^{M}\sum_{n=1}^{M}(\pi_{km}-\pi_{ln})\mathcal{C}_{m}+\pi_{ln}\mathcal{C}_{n}\right\|_{2}.$ (26)
-Taking the expectation over all domain pairs $k\neq l$ and using linearity, we have:
-$\Phi_{\text{inter}}=\mathbb{E}_{k\neq l}\left[\sum_{m=1}^{M}\sum_{n=1}^{M}(\pi_{km}\pi_{ln})\|\mathcal{C}_{m}-\mathcal{C}_{n}\|_{2}\right].$ (27)
-Recognizing that $\lambda_{m}=\mathbb{E}_{k}[\pi_{km}]$ is the expected activation of capability $m$, we can simplify further:
-$=\sum_{m=1}^{M}\sum_{n=1}^{M}\lambda_{m}\lambda_{n}\|\mathcal{C}_{m}-\mathcal{C}_{n}\|_{2}.$ (28)
-This completes the proof, showing that the inter-diversity metric is determined by the expected pairwise distances between latent capability centroids, weighted by their activation probabilities.
-Proof of Intra-Diversity
-The intra-diversity metric for domain $k$, denoted as $\Phi_{\text{intra}}^{(k)}$, measures the variance of data samples within the domain relative to its centroid $\mathcal{C}_{k}$.
-We aim to show that:
-$\Phi_{\text{intra}}^{(k)}\leq\sum_{m=1}^{M}\pi_{km}\|\mathcal{C}_{m}-\mathcal{C}_{k}\|_{2}^{2}+\mathbb{E}_{m\sim\pi_{k}}[\text{Var}(\tilde{\mathcal{D}}_{m})].$ (29)
-Proof.
-To prove this bound, we start by expressing the intra-diversity metric $\Phi_{\text{intra}}^{(k)}$ as the expected squared distance from each sample $\mathbf{x}_{i}^{(k)}$ to the domain centroid $\mathcal{C}_{k}$ from Eq (5):
-$\Phi_{\text{intra}}^{(k)}=\frac{1}{N_{k}}\sum_{i=1}^{N_{k}}\|\mathbf{x}_{i}^{(k)}-\mathcal{C}_{k}\|_{2}^{2}.$ (30)
-Using the definition of the domain centroid (Eq (10)) $\mathcal{C}_{k}=\sum_{m=1}^{M}\pi_{km}\mathcal{C}_{m}$ in , we can decompose the squared distance for each sample:
-$\|\mathbf{x}_{i}^{(k)}-\mathcal{C}_{k}\|_{2}^{2}=\|\mathbf{x}_{i}^{(k)}-\sum_{m=1}^{M}\pi_{km}\mathcal{C}_{m}\|_{2}^{2}.$ (31)
-Applying the variance decomposition, we have:
-$\|\mathbf{x}_{i}^{(k)}-\mathcal{C}_{k}\|_{2}^{2}=\|\mathbf{x}_{i}^{(k)}-\sum_{m=1}^{M}\pi_{km}\mathcal{C}_{m}\|_{2}^{2}=\sum_{m=1}^{M}\pi_{km}\|\mathbf{x}_{i}^{(k)}-\mathcal{C}_{m}\|_{2}^{2}+\sum_{m=1}^{M}\pi_{km}\|\mathcal{C}_{m}-\mathcal{C}_{k}\|_{2}^{2},$ (32)
-where the first term represents the variance within each latent capability, and the second term accounts for the distance between latent capability centroids and the domain centroid.
-By taking the expectation over the samples in domain $k$, we get:
-$\Phi_{\text{intra}}^{(k)}\leq\sum_{m=1}^{M}\pi_{km}\|\mathcal{C}_{m}-\mathcal{C}_{k}\|_{2}^{2}+\mathbb{E}_{m\sim\pi_{k}}[\text{Var}(\tilde{\mathcal{D}}_{m})],$ (33)
-where $\text{Var}(\tilde{\mathcal{D}}_{m})$ represents the variance within the $m$-th latent capability. This completes the proof of the intra-diversity bound. ∎
-Appendix D More Experimental Results
-D.1 Visualization of Embeddings on Llama3.1-8B & Qwen2.5-7B
-We process concatenated data samples with “instruction”+“input”+“output” pairs through each LLM’s embedding layer, computing mean token embeddings for visualization.
-Beyond Qwen2-7B results in Fig.
-1, we present the t-SNE visualization of data samples from Llama3.1-8 and Qwen2.5-7B with different distributions in Fig.
-8 and Fig.
+Three key observations and conjectures emerge:
+  • Certain methods (particularly Instruction Length, Deita-C, and Deita-Q) demonstrate pronounced distributional skew toward coding-domain samples, which may partially explain their suboptimal HumanEval performance shown in Table 20.
+  • The domain-agnostic nature of baseline approaches reveals fundamental limitations in preserving original data distributions when processing domain-unspecified samples, as evidenced by the visualization.
+  • Although distributional shifts are evident, the absence of commensurate performance deterioration across all methods empirically validates our theoretical framework (Section 3.4), which establishes connections between data distribution shift.
+Figure 5: t-SNE visualization of data samples selected by different baselines using Llama3.1-8B embeddings. While original labels were removed during training, we preserve them for interpretability.
+Figure 6: t-SNE visualization of data samples selected by different baselines using Qwen2-7B embeddings. While original labels were removed during training, we preserve them for interpretability.
+Figure 7: t-SNE visualization of data samples selected by different baselines using Qwen2.5-7B embeddings. While original labels were removed during training, we preserve them for interpretability.
+Appendix E Derivation of Proposition 3.2
+We derive the density ratio by systematically expanding and substituting terms:
+$\frac{q(\mathbf{x})}{p(\mathbf{x})}=\frac{\sum_{k=1}^{K}q(c=k)q(\mathbf{x}|c=k)}{p(\mathbf{x})}$ (10)
+Expand $q(\mathbf{x})$ via the law of total probability over domains.
+$=\frac{\sum_{k=1}^{K}q(c=k)p(\mathbf{x}|c=k)}{p(\mathbf{x})}$ (11)
+Apply Assumption 3.1 ($q(\mathbf{x}|c=k)=p(\mathbf{x}|c=k)$) to unify domain-conditional densities.
+$=\frac{\sum_{k=1}^{K}q(c=k)\frac{p(c=k|\mathbf{x})p(\mathbf{x})}{p(c=k)}}{p(\mathbf{x})}$ (12)
+Express $p(\mathbf{x}|c=k)$ via Bayes’ rule inversion: $p(\mathbf{x}|c=k)=\frac{p(c=k|\mathbf{x})p(\mathbf{x})}{p(c=k)}$.
+$=\sum_{k=1}^{K}\frac{q(c=k)}{p(c=k)}p(c=k|\mathbf{x})\cdot\frac{p(\mathbf{x})}{p(\mathbf{x})}$ (13)
+Factor out $p(\mathbf{x})$ from the numerator and the denominator.
+$=\sum_{k=1}^{K}\underbrace{\frac{q(c=k)}{p(c=k)}}_{\lambda_{k}}p(c=k|\mathbf{x})$ (14)
+Define domain proportion ratio $\lambda_{k}:=\frac{q(c=k)}{p(c=k)}$.
+This demonstrates that the density ratio decomposes linearly into domain-specific components weighted by their proportion shifts ($\lambda_{k}$) and posterior probabilities ($p(c=k|\mathbf{x})$).
+Appendix F Comprehensive Experiments
+F.1 Performance on Comprehensive Benchmark MMLU
+The benchmarks selected in Section 3.1 are specifically tailored to correspond with the domain-specific capabilities of the data sources.
+To comprehensively evaluate the performance of DaaR across broader assessment metrics and more extensive benchmarking frameworks, we adopted MMLU as an out-of-distribution (OOD) evaluation protocol, detailed description as:
+  • MMLU is a benchmark designed to evaluate the multitask capabilities of language models across diverse subjects, covering 57 tasks spanning topics with questions ranging from elementary to professional levels.
+Table 6: Evaluation results of Qwen2-7B and Qwen2.5-7B on MMLU.
+Models Methods Humanities STEM Social Science Other MMLU-Avg
+Qwen2-7B Raw 38.58 25.77 24.94 27.82 28.98
+Random 70.02 58.54 80.00 72.26 68.81
+Instruction Len 71.34 58.09 79.77 70.71 68.55
+Alpagasus [10] 56.55 49.10 71.88 65.54 59.34
+Instag-C [30] 70.28 58.69 79.92 71.17 68.65
+Instag-D [30] 67.87 58.35 79.82 71.61 68.07
+SuperFilter [26] 52.55 49.16 57.99 60.01 54.27
+Deita-C [29] 68.05 58.53 80.12 70.97 68.08
+Deita-D [29] 67.11 58.22 79.17 72.13 67.83
+Deita-Q [29] 56.79 52.82 69.01 63.06 59.47
+DaaR (Ours) 70.49 59.94 79.67 72.74 69.42
+Qwen2.5-7B Raw 73.41 48.96 72.79 58.46 61.72
+Random 75.40 64.47 81.46 72.71 72.42
+Instruction Len 75.08 63.74 81.40 72.96 72.15
+Alpagasus [10] 75.39 64.85 81.73 73.13 72.70
+Instag-C [30] 75.26 64.35 81.64 73.03 72.46
+Instag-D [30] 74.83 65.05 81.48 73.16 72.59
+SuperFilter [26] 75.24 66.34 82.19 73.99 73.45
+Deita-C [29] 75.76 64.18 81.59 72.83 72.46
+Deita-D [29] 75.27 65.60 81.55 72.99 72.85
+Deita-Q [29] 75.81 64.14 81.41 73.65 72.61
+DaaR (Ours) 74.91 65.84 81.85 73.68 73.07
+As demonstrated in Table 6, despite not being explicitly optimized for broader capabilities during its design phase, DaaR still demonstrates competitive performance in comprehensive capability assessments.
+The method achieved the highest score on Qwen2-7B, surpassing the second-place Random selection by 0.61 accuracy points.
+On Qwen2.5-7B, it exhibited performance closely following SuperFilter, thereby confirming the generalization potential of DaaR to a certain extent.
+Notably, Llama3.1-8B was excluded from reference comparisons due to significant result fluctuations observed during Opencompass evaluations (as shown in Table 7 that the results range from 0.04 to 42.52).
+We further identified inconsistent failures in model prediction generation during the evaluation process, which compromised the reliability of performance measurement.
+These implementation challenges prevented valid comparative analysis with Llama3.1-8B in our experiments.
+Table 7: Invalid evaluation results of Llama3.1-8B on MMLU.
+Models Methods Humanities STEM Social Science Other MMLU-Avg
+Llama3.1-8B Raw 0.02 0.01 0.08 0.07 0.04
+Random 30.13 15.34 30.73 27.51 24.73
+Instruction Len 23.38 11.65 20.43 18.26 17.68
+Alpagasus [10] 35.99 23.28 36.80 34.37 31.56
+Instag-C [30] 25.87 12.58 33.73 23.71 22.60
+Instag-D [30] 42.50 19.44 47.70 37.19 34.70
+SuperFilter [26] 11.61 9.40 12.11 11.71 11.00
+Deita-C [29] 8.79 0.97 0.60 2.27 2.97
+Deita-D [29] 50.79 29.76 53.04 43.21 42.52
+Deita-Q [29] 33.99 23.36 26.31 24.27 26.61
+DaaR (Ours) 45.18 20.94 46.47 34.20 34.87
+F.2 Validation on Qwen3-8B
+As a prevailing trend in the development of current LLMs, the RL-based (reinforcement learning) thinking model has demonstrated significant capabilities and potential (e.g.
+DeepSeek series [28], Llama4 series and Qwen3 series).
+To validate the efficacy of DaaR on the most up-to-date and advanced LLMs, we conducted a series of verification experiments using Qwen3-8B  [^2] https://github.com/QwenLM/Qwen3.
+It should be noted that, since Qwen3 does not adhere to conventional SFT paradigms, as well as the comprehensive support for training & evaluation of the Qwen3 series remains incomplete, the experiments presented in this section primarily serve as exploratory investigations.
+A comprehensive analysis of data diversity across various training methodologies and LLM architecture will be reserved for future research endeavors.
+Observation Studies on Qwen3
+Following the methodology delineated in Section 3, we employed Qwen3’s embedding layer to process the identical data pool.
+The t-SNE visualization of Qwen3’s embeddings (Fig.
+8) demonstrates comparable domain discrimination capabilities.
+Figure 8: The t-SNE visualization of embeddings for data samples with different distributions on Qwen3-8B.
+(a) The data pool of all 40K samples, (b) Distribution of data farthest from other domain centroids on Inter-Diversity, (c) Distribution of data closest to other domain centroids on Inter-Diversity, (d) Distribution of data closest to its own domain centroid on Inter-Diversity, (e) Distribution of data farthest from its own domain centroid on Inter-Diversity.
+Subsequently, we processed the curated dataset through inter-diversity and intra-diversity filters, followed by training Qwen3 on artificially constructed data distributions.
+As evidenced in Table 8, the experimental outcomes reveal that while the performance patterns diverge from those observed in Llama 3.1, Qwen2, and Qwen2.5, varying diversity levels demonstrably influence comprehensive model performance.
+Notably, the absolute value of the comprehensive performance of Qwen3 is even inferior to that of Qwen2.5 and Qwen2.
+We posit that this phenomenon is primarily due to the incomplete support of the evaluation platform OpenCompass (as of the submission date), where the distinctive thinking and reasoning capabilities of Qwen3 have not been fully leveraged.
+However, the relative comparison of different processes within the same data pool remains highly informative.
+Table 8: Validation results of Inter-Diversity and Intra-Diversity on Qwen3-8B across benchmarks.
+Qwen3-8B Common Sense Reasoning Mathematics Coding Avg
+NQ TriviaQA Hellaswag GSM8K MATH MBPP HumanEval
+Raw 9.64 59.90 71.56 85.60 23.60 22.40 73.17 49.41
+Rand (8K) 12.24 59.22 71.71 83.00 25.2 12.6 73.17 48.16
+Inter-Diversity (0-20) 12.94 60.28 71.79 86.20 23.80 11.80 75.61 48.92
+Inter-Diversity (40-60) 12.71 59.76 71.71 83.20 28.40 14.40 73.78 49.14
+Inter-Diversity (80-100) 11.36 59.52 71.61 83.00 25.40 14.20 68.29 47.63
+Intra-Diversity (0-20) 12.33 59.92 71.58 84.00 28.40 19.40 71.34 49.57
+Intra-Diversity (40-60) 13.10 59.42 71.64 82.40 21.00 13.20 72.56 47.62
+Intra-Diversity (80-100) 12.58 60.06 71.88 84.60 34.90 9.40 73.17 49.51
+DaaR Training on Qwen3
+Subsequently, we employed DaaR to conduct diversity probe training on Qwen3.
+Consistent with observations in other models, the two-stage MLP training framework demonstrated robust stability and convergence, shown in Fig.
 9.
-It is evident that, despite differences in architecture or model, the method of filtering data through Inter-Diversity and Intra-Diversity is effective.
-Notably, although the embedding distributions of different models are not identical, they exhibit similar behavior, indicating that the representations learned by the embedding layers are comparable.
-Figure 8: The t-SNE visualization of embeddings for data samples with different distributions on Llama3.1-8B.
-(a) The data pool of all 40,000 samples, (b) Randomly selected subset, (c) Distribution of data farthest from other domain centroids based on Inter-Diversity, (d) Distribution of data closest to other domain centroids based on Inter-Diversity, (e) Distribution of data closest to its own domain centroid based on Inter-Diversity, (f) Distribution of data farthest from its own domain centroid based on Inter-Diversity.
-Figure 9: The t-SNE visualization of embeddings for data samples with different distributions on Qwen2.5-7B.
-(a) The data pool of all 40,000 samples, (b) Randomly selected subset, (c) Distribution of data farthest from other domain centroids based on Inter-Diversity, (d) Distribution of data closest to other domain centroids based on Inter-Diversity, (e) Distribution of data closest to its own domain centroid based on Inter-Diversity, (f) Distribution of data farthest from its own domain centroid based on Inter-Diversity.
-D.2 The Visualization of Inter- & Intra-Diversity Distribution Data
-Using the Qwen2-7B model as an example, we construct data based on the Inter-Diversity and Intra-Diversity distribution methods, selecting a batch of data every 20%.
-The visualization process is shown in Fig.
-10 and Fig.
-11.
-As seen in the figures, the data gradually transitions from domain-aware diverse to domain-aware closed, indicating that our data construction method effectively controls the distribution of different data.
-However, it is important to note that t-SNE dimensionality reduction is used only for visualization purposes; during the reduction process, positional information is lost, and it cannot fully capture the true high-dimensional relationships between the data.
-Figure 10: Data visualization (t-SNE) based on different Inter-Diversity distributions on Qwen2-7B.
-Figure 11: Data visualization (t-SNE) based on different Intra-Diversity distributions on Qwen2-7B.
-D.3 Similarity in Generated Centroid Quantities
-To assess the impact of the amount of generated data on the accuracy of Domain Semantic Centroids, we select quantities of 10 and 30 samples for generation.
-Subsequently, we calculate the semantic cosine similarity between these two sets, with results presented in Table 5 (using Qwen2-7B and Llama3.1-8B as examples).
-The observed differences between varying data quantities were not significant, indicating that the generalization of the generated data is sufficient and further increasing the data quantity does not significantly enhance the accuracy of the centroids.
-Table 5: Semantic cosine similarity between generated samples of size 10 and 30 across different domains.
-Similarity of 10 & 30 Common Sense Reasoning Mathematics Coding
-Qwen2-7B 0.9686 0.9866 0.9919 0.9881
-Llama3.1-8B 0.9851 0.9795 0.9685 0.9895
-D.4 Similarity in Generated Centroid Domains
+Figure 9: Training loss and validation process of the two training stages of DaaR on Qwen3-8B, showing the model gradually converging.
+Evaluation Results of DaaR with Baselines
+Following the identical experimental setup outlined in Appendix B, we conducted experiments on Qwen3, with the results presented in Table 9.
+As demonstrated, DaaR achieved the top comprehensive score on Qwen3, surpassing all baselines by up to 4.7%.
+Notably, its performance on the MATH benchmark was particularly outstanding, exceeding the average of baselines by 25.91%.
+The comparative experiments affirm that DaaR maintains competitive effectiveness even on the SOTA LLM Qwen3-8B.
+Table 9: Performance of DaaR with baselines on Qwen3-8B across various benchmarks.
+Qwen3-8B Common Sense Reasoning Mathematics Coding Avg
+NQ TriviaQA Hellaswag GSM8K MATH MBPP HumanEval
+Raw 9.64 59.90 71.56 85.60 23.60 22.40 73.17 49.41
+Rand (8K) 12.24 59.22 71.71 83.00 25.20 12.60 73.17 48.16
+Instruction Len 11.66 59.56 71.67 83.80 26.30 18.20 73.17 49.19
+Alpagasus [10] 12.52 59.34 71.69 83.80 26.60 11.20 69.51 47.81
+Instag-C [30] 11.91 59.62 71.55 83.00 26.40 14.60 73.78 48.69
+Instag-D [30] 11.99 59.48 71.59 87.20 25.30 13.60 71.34 48.64
+SuperFilter [26] 15.73 60.18 71.69 84.80 30.90 16.20 67.07 49.51
+Deita-C [29] 12.55 59.62 71.78 83.60 28.60 13.20 77.44 49.54
+Deita-Q [29] 12.88 59.72 71.64 85.40 32.90 13.00 73.78 49.90
+Deita-D [29] 13.88 59.98 71.50 83.60 26.30 14.00 70.73 48.57
+DaaR (Ours) 13.68 60.14 71.56 83.40 34.50 14.60 72.56 50.06
+F.3 Customized Data Selection Capability of DaaR
+A distinctive characteristic of our method stems from its inherent capability to perform domain-specific data selection under a domain-undetermined scenario, thereby naturally supporting customized data curation.
+In practical scenarios beyond comprehensive LLM improvement, mission-critical applications (e.g., developing specialized LLM variants for coding or mathematical reasoning) often demand focused enhancement on target capabilities.
+Conventional baseline selection approaches, which lack domain-aware mechanisms, fundamentally preclude such fine-grained control.
+To validate this customized selection capability, we conduct a validated experiment on Qwen2.5-7B.
+By strategically allocating higher selection ratios (50%) to specific target domains while reducing others to 10%, while maintaining the total training corpus size at 8,000 instances, we investigate whether such curated data distribution induces domain-specific performance gains.
+As presented in Table 10, our analysis reveals that compared to the strongest baselines (Random and original DaaR), customized DaaR achieves SOTA enhancement in focused domains, despite observing marginal performance degradation in non-target domains that leads to slightly inferior overall performance compared to the original DaaR.
+Interestingly, the performance on Reasoning-Major takes first place on Avg score, highlighting that the reasoning capability may be the crucial factor of LLMs.
+This validation promisingly confirms the operational feasibility of our customization paradigm and demonstrates DaaR’ unique superiority over existing baselines in enabling domain-prioritized LLM specialization.
+Table 10: Validation results of customized selected ratio on Qwen2.5-7B across benchmarks.
+Qwen2.5-7B Common Sense Reasoning Mathematics Coding Avg
+NQ TriviaQA Hellaswag GSM8K MATH MBPP HumanEval
+Raw 8.84 58.14 72.75 78.20 9.10 7.40 78.05 44.64
+Rand (8K) 11.46 57.85 73.08 78.90 13.15 62.50 71.65 52.66
+DaaR (Original) 15.83 58.65 72.48 80.20 16.70 64.20 68.29 53.76
+Common-Major 17.56 59.58 72.51 77.80 12.85 62.10 67.53 52.85
+Reasoning-Major 16.56 58.22 74.62 77.30 17.20 63.40 66.42 53.39
+Math-Major 13.20 57.96 74.16 81.40 16.05 61.80 63.45 52.75
+Coding-Major 10.76 57.02 72.78 78.20 12.80 65.20 73.22 52.85
+F.4 Rationale for Entropy over Diversity as the Reward Signal
+As discussed in Section 4, the explicit diversity metric defined in Eq.
+(2) & (3) inherently depends on cross-sample relationships that cannot be captured through conventional sample-level training paradigms.
+To empirically validate this limitation, we conducted an ablation study using 500 representative samples.
+The inter-diversity scores were calculated using Eq.
+(2) and partitioned into ascending quartiles, each representing distinct diversity levels.
+A 5-layer MLP classifier was subsequently trained with cross-entropy loss as a more tractable verification approach compared to regression.
+Experimental result (Fig.
+10) demonstrates persistent convergence challenges, evidenced by consistently high training loss and subpar classification accuracy capped at 34%.
+These findings substantiate our hypothesis that LLMs fail to learn cross-sample characteristics, thereby necessitating entropy-based proxy metrics for effective reward formulation.
+Figure 10: Validation loss and accuracy on directly use diversity instead of entropy as a reward signal.
+Appendix G Complete Ablation Studies and Discussion
+G.1 Similarity in Generated Domains’ Centroid
 To investigate whether there is a distinct separation between different domains generated, we compute the cosine similarity of semantic centroids across different domains.
@@
 Figure 11: Semantic cosine similarity across different domains for generated samples of size 10.
-D.5 DaaR Layer Selection Protocol
-To determine an appropriate layer for embedding DaaR into a large language model (LLM) that balances performance and computational cost, we first evaluate the capacity of different layers for domain awareness.
+G.2 DaaR Layer Selection Protocol
+We choose layer-3 as the hidden layer attached to $\psi_{dom}$.
+To determine an appropriate layer for embedding DaaR into an LLM that balances performance and computational cost, we first evaluate the capacity of different layers for domain awareness.
 Using Qwen2-7B as an example, we conduct t-SNE visualizations of Layer-5, Layer-10, Layer-15, Layer-20, Layer-25, and the last hidden layer, based on the average token vectors, as shown in Fig.
-13.
+12.
 Figure 12: t-SNE visualization of mean token embeddings across Layer-5, Layer-10, Layer-15, Layer-20, Layer-25, and the last hidden layer of Qwen2-7B.
-Our observations indicate that even at the early Layer-5, the model begins to exhibit classification capabilities across different datasets.
-This prompted us to focus on the initial five layers to enhance learning outcomes.
-Considering that data clustering is inherently influenced by the embedding layer, a natural approach was to directly connect to the embedding layer for training.
-However, the training results showed persistently high training loss and a validation accuracy of only about 0.6, indicating suboptimal learning performance.
-Furthermore, related studies indicate that LLMs can learn semantic understanding and perception in the early layers (Wei et al., 2022), consequently, we select Layer-3 for DaaR.
-Figure 14: Training loss and validation process of Embedding Layer on Qwen2-7B.
-D.6 Comprehensive Training Dynamics
+The observations demonstrate that LLMs maintain cross-dataset classification capabilities across all architectural depths.
+Considering the diminishing marginal utility of computational resources in deeper layers, we prioritize initial layers to optimize learning efficiency.
+The training dynamics across the first ten layers are visualized in Fig.
+13, demonstrating consistent accuracy within the 91.2–93.6% range, verifying that various layers can effectively support DaaR first-stage training.
+Figure 13: Validation loss and accuracy over steps on different layers.
+Given that data clustering is based on the embedding layer, a natural approach was to directly connect to the embedding layer for training.
+However, the training results showed persistently high training loss and a validation accuracy of only about 0.6 (shown in Fig.
+14), indicating suboptimal learning performance.
+Furthermore, related studies indicate that LLMs can learn semantic understanding and perception in the early layers [40], consequently, we select Layer-3 for DaaR.
+Figure 14: Training loss and validation process of Embedding Layer on Qwen2.5-7B.
+G.3 Ablation of the Number of Seed Samples
+To investigate the impact of the number of seed samples $\mathcal{S}_{k}^{(0)}$ on domain centroid construction, we systematically selected 3, 5, and 10 as the number of seed samples while maintaining all other parameters at their default settings on Qwen2.5-7B.
+For each configuration, pairwise cosine similarity between domains was computed to assess whether significant discrepancies emerged across different seed sizes.
+Experimental results summarized in Table 11, demonstrate that similarities remain highly consistent across all domain pairs.
+Except for the common sense domain (with similarity approaching 0.9), varying seed sample counts do not yield substantial differences in the final centroids, with variance ranging from 5.4e-7 to 1.8e-5, demonstrating the stability of this hyperparameter.
+Intuitively, we posit that excessively small seed sizes may compromise domain characterization due to generation randomness, while larger sizes risk introducing redundant patterns.
+We conjecture that the relatively lower absolute similarity in common sense domains stems from inherent sentence brevity, which increases vulnerability to lexical biases.
+However, this does not impair the anchoring effectiveness, as evidenced by the consistent performance across configurations.
+Table 11: Semantic similarity between the number of seed samples among 3, 5 and 10 across domains on Qwen2.5-7B.
+Domains Similarity of 3 & 5 Similarity of 3 & 10 Similarity of 5 & 10
+Common Sense 0.8984 0.8913 0.9074
+Reasoning 0.9713 0.9752 0.9815
+Mathematics 0.9776 0.9761 0.9722
+Coding 0.9775 0.9759 0.9760
+G.4 Ablation of the Number of Diversity Augmentation Data
+To assess the impact of the amount of generated data on the accuracy of domain semantic centroids, we select quantities of 10 and 30 samples for generation.
+Subsequently, we calculate the semantic cosine similarity between these two sets, with results presented in Table 12 (using Qwen2.5-7B and Llama3.1-8B as examples).
+The observed differences between varying data quantities were not significant, indicating that the generalization of the generated data is sufficient, and further increasing the data quantity does not significantly enhance the accuracy of the centroids.
+Table 12: Semantic similarity between generated samples of size 10 and 30 across different domains.
+Similarity of 10 & 30 Common Sense Reasoning Mathematics Coding
+Qwen2.5-7B 0.9686 0.9866 0.9919 0.9881
+Llama3.1-8B 0.9851 0.9795 0.9685 0.9895
+G.5 Ablation of the Size of Sliding Window
+To investigate the impact of the sliding window size in diversity augmentation, we kept all other hyper-parameters constant, setting the sliding window size to 1, 3, and 5 examples, respectively, and compared the similarity between them.
+The experimental results on Qwen2.5-7B, presented in Table 13, indicate that different window sizes maintain a high degree of similarity, suggesting that DaaR does not heavily depend on this hyperparameter choice.
+Considering the prompt length and stability, we opted for a window size of 3 in the implementation.
+Table 13: Semantic similarity between the size of sliding window among 1, 3 and 5 across domains on Qwen2.5-7B.
+Domains Similarity of 1 & 3 Similarity of 1 & 5 Similarity of 3 & 5
+Common Sense 0.9205 0.9094 0.9022
+Reasoning 0.9746 0.9760 0.9772
+Mathematics 0.9736 0.9618 0.9847
+Coding 0.9849 0.9647 0.9716
+G.6 Ablation of the Diversity Threshold
+The threshold $\tau$ design needs to satisfy two objectives: (1) mitigate redundancy in LLM-generated outputs by preventing exact duplicates in synthetic data, and (2) ensure sufficient similarity among instances within the same domain embedding space.
+As illustrated in Table 14, we conduct an experiment using different similarity thresholds across various domains on Qwen2.5-7B, observing the number of attempts required for successful generation.
+The results indicate that distinct domains exhibit varying sensitivity to similarity metrics, with the number of attempts increasing sharply as the threshold decreases.
+Based on these observations, we empirically determined a threshold value of 0.9 for Mathematics and 0.85 for others as a balanced compromise.
+Table 14: Generation attempt frequencies for 10 seed samples across different domains and similarity levels on Qwen2.5-7B.
+Domains 0.8 0.85 0.9 0.95
+Common Sense 153 31 21 12
+Reasoning 135 61 44 15
+Mathematics $>1000$ 804 236 25
+Coding 74 21 16 13
+We further conducted ablation experiments by comparing the similarity of embedding centroids generated under different thresholds to systematically evaluate the impact of threshold variations on downstream performance.
+The results presented in Table 15 demonstrate, with the exception of common sense data (similar to the results in Table 11), that varying thresholds do not significantly affect the role of these centroids as initial points for clustering.
+We conjecture that lower thresholds ensure greater diversity within domain, whereas higher thresholds tend to result in repetitive data.
+Table 15: Semantic similarity between the similarity threshold among 0.85, 0.9, and 0.95 across domains on Qwen2.5-7B.
+Domains Similarity of 0.85 & 0.9 Similarity of 0.85 & 0.95 Similarity of 0.85 & 0.95
+Common Sense 0.8869 0.8963 0.9502
+Reasoning 0.9791 0.9785 0.9795
+Mathematics 0.9730 0.9754 0.9703
+Coding 0.9646 0.9740 0.9691
+G.7 Cost-Efficiency and Flexibility
+Compared to baseline methods requiring GPT-based evaluators (Alpagasus, Instag) or full-LLM inference (Deita), our approach achieves superior efficiency through data-model co-optimizations.
+Our method demonstrates computational efficiency through two key aspects enabled by the LLM’s self-rewarding capability during dedicated data synthesis:
+  • Our framework operates on frozen embeddings extracted from layer 3 while truncating subsequent layers, which constitutes a significantly shallower architecture compared to the conventional 32-layer structure in 7B-scale LLMs.
+Taking Qwen2 as an example, full model inference requires loading complete parameters alongside cache management components, consuming 18–24 GB GPU memory.
+In contrast, our method accomplishes data filtering with merely 5–6 GB GPU memory footprint.
+  • The lightweight 5-layer MLP probe module introduces only 76 million additional parameters.
+As a regression model producing single scalar outputs per instance, it achieves substantial speed improvements over LLM-inference-based approaches.
+Experimental results on 5,000 samples demonstrate this efficiency advantage: our method requires approximately 40 minutes for sample-level inference, whereas conventional LLM inference approaches demand nearly two hours.
+G.8 Stability of DaaR
+Our dual-MLP design of DaaR strategically separates entropy regression (second MLP) from neural softmax-based measurement (first MLP), both showing stable convergence in Fig 2.
+While precise analysis of compounding errors remains a theoretical concern, our framework inherently mitigates these errors through these aspects: (1) Seed Generation: LLM-generated descriptions exhibit minimal variation across iterations, and repeated seeding mitigates uncertainty.
+(2) Layer Selection: Semantic feature extraction across layers is demonstrated to effectively capture domain-specific patterns, thereby establishing methodological robustness, detailed in Appendix G.2.
+(3) Clustering: Data seeding as initialization for K-means clustering, where minimal perturbation of initial points ensures stable training of the first MLP.
+(4) Selection Strategy: Percentage-based data selection prioritizing sample distribution robustness over fine-grained entropy value utilization.
+Appendix H Comprehensive Visualization and Experimental Results
+H.1 Visualization of Embeddings on Llama3.1-8B , Qwen2-7B & Qwen2.5-7B
+We process concatenated data samples with “instruction”+“input”+“output” pairs through each LLM’s embedding layer, computing mean token embeddings for visualization.
+We present the t-SNE visualization of data samples from Llama3.1-8, Qwen2-7B and Qwen2.5-7B with different distributions in Fig.
+15, Fig.
+16 and Fig.
+17.
+It is evident that, despite differences in architecture or model, the method of filtering data through Inter-Diversity and Intra-Diversity is effective.
+Notably, although the embedding distributions of different models are not identical, they exhibit similar behavior, indicating that the representations learned by the embedding layers are comparable.
+Figure 15: The t-SNE visualization of embeddings for data samples with different distributions on Llama3.1-8B.
+(a) The data pool of all 40K samples, (b) Randomly selected subset, (c) Distribution of data farthest from other domain centroids on Inter-Diversity, (d) Distribution of data closest to other domain centroids on Inter-Diversity, (e) Distribution of data closest to its own domain centroid on Inter-Diversity, (f) Distribution of data farthest from its own domain centroid on Inter-Diversity.
+Figure 16: The t-SNE visualization of embeddings for data samples with different distributions on Qwen2-7B.
+(a) The data pool of all 40K samples, (b) Randomly selected subset, (c) Distribution of data farthest from other domain centroids on Inter-Diversity, (d) Distribution of data closest to other domain centroids on Inter-Diversity, (e) Distribution of data closest to its own domain centroid on Inter-Diversity, (f) Distribution of data farthest from its own domain centroid on Inter-Diversity.
+Figure 17: The t-SNE visualization of embeddings for data samples with different distributions on Qwen2.5-7B.
+(a) The data pool of all 40K samples, (b) Randomly selected subset, (c) Distribution of data farthest from other domain centroids on Inter-Diversity, (d) Distribution of data closest to other domain centroids on Inter-Diversity, (e) Distribution of data closest to its own domain centroid on Inter-Diversity, (f) Distribution of data farthest from its own domain centroid on Inter-Diversity.
+H.2 Visualization of Inter- & Intra-Diversity Distribution Data
+Using the Qwen2-7B model as an example, we construct data based on the Inter-Diversity and Intra-Diversity distribution methods, selecting a batch of data every 20%.
+The visualization process is shown in Fig.
+18 and Fig.
+19.
+As seen in the figures, the data gradually transitions from domain-aware diverse to domain-aware closed, indicating that our data construction method effectively controls the distribution of different data.
+However, it is important to note that t-SNE dimensionality reduction is used only for visualization purposes; during the reduction process, positional information is lost, and it cannot fully capture the true high-dimensional relationships between the data.
+Figure 18: Data visualization (t-SNE) based on different Inter-Diversity distributions on Qwen2-7B.
+Figure 19: Data visualization (t-SNE) based on different Intra-Diversity distributions on Qwen2-7B.
+H.3 Comprehensive Training Dynamics
 We present the training loss and validation results for two distinct models, Llama3.1-8B and Qwen2.5-7B, in Fig 20.
@@
 Figure 20: Training loss and validation process of the two training stages of DaaR on Llama3.1-8B and Qwen2.5-7B, showing the model gradually converging.
-D.7 Complete Validation Results of Inter-Diversity and Intra-Diversity
+H.4 Complete Validation Results of Inter-Diversity and Intra-Diversity
 We present the complete experimental results of Llama3.1-8B, Qwen2-7B, and Qwen2.5-7B in the validation experiments in Tables 16, Tables 17, and Tables 18, respectively.
 It can be observed that for any complete dataset, the conclusions from Section 3.3 remain valid, specifically that the peak distribution of results is uneven, with significant differences among them.
-Table 6: Validation Results of Inter-Diversity and Intra-Diversity on Llama3.1-8B across various benchmarks.
+Table 16: Validation results of Inter-Diversity and Intra-Diversity on Llama3.1-8B across benchmarks.
 Llama3.1-8B Common Sense Reasoning Mathematics Coding Avg
@@
 Raw 14.13 65.90 74.62 54.80 7.90 5.00 28.66 35.86
-Rand 21.99 64.83 74.72 55.70 14.50 5.10 24.09 37.27
+Full (40K) 21.92 65.11 73.62 51.70 7.60 4.10 36.50 37.22
+Rand (8K) 21.99 64.83 74.72 55.70 14.50 5.10 24.09 37.27
 Inter-Diversity (0-20) 19.28 65.79 74.44 54.90 6.50 4.30 35.06 37.18
@@
 Intra-Diversity (80-100) 19.78 64.77 74.51 56.50 13.00 5.20 37.50 38.75
-Table 7: Validation Results of Inter-Diversity and Intra-Diversity on Qwen2-7B across various benchmarks.
+Table 17: Validation results of Inter-Diversity and Intra-Diversity on Qwen2-7B across benchmarks.
 Qwen2-7B Common Sense Reasoning Mathematics Coding Avg
@@
 Raw 8.03 59.58 73.00 78.00 5.70 5.00 60.98 41.47
-Rand 13.28 58.27 73.00 75.35 35.36 52.20 63.72 53.02
+Full (40K) 15.61 58.75 72.51 73.80 31.30 51.70 67.38 53.01
+Rand (8K) 13.28 58.27 73.00 75.35 35.36 52.20 63.72 53.02
 Inter-Diversity (0-20) 15.18 59.28 73.34 74.50 34.94 53.10 68.60 54.13
@@
 Intra-Diversity (80-100) 11.91 57.88 73.29 75.00 36.05 52.50 66.16 53.25
-Table 8: Validation Results of Inter-Diversity and Intra-Diversity on Qwen2.5-7B across various benchmarks.
+Table 18: Validation results of Inter-Diversity and Intra-Diversity on Qwen2.5-7B across benchmarks.
 Qwen2.5-7B Common Sense Reasoning Mathematics Coding Avg
@@
 Raw 8.84 58.14 72.75 78.20 9.10 7.40 78.05 44.64
-Rand 11.46 57.85 73.08 78.90 13.15 62.50 71.65 52.65
+Full (40K) 12.88 58.60 72.28 76.80 13.60 62.80 71.04 52.57
+Rand (8K) 11.46 57.85 73.08 78.90 13.15 62.50 71.65 52.65
 Inter-Diversity (0-20) 13.23 58.15 73.27 78.70 11.45 62.30 69.21 52.33
@@
 Intra-Diversity (80-100) 12.79 58.09 73.21 75.40 16.05 62.50 49.39 49.63
-D.8 Complete Results of DaaR with Baselines
+H.5 Complete Results of DaaR with Baselines
 Additionally, we include the complete results of DaaR and the comparative baselines in the three tables: Table 19, Table 20 and Table 21.
@@
 Raw 14.13 65.90 74.62 54.80 7.90 5.00 28.66 35.86
-Rand 21.99 64.83 74.72 55.70 14.50 5.10 24.09 37.27
+Full (40K) 21.92 65.11 73.62 51.70 7.60 4.10 36.50 37.22
+Rand (8K) 21.99 64.83 74.72 55.70 14.50 5.10 24.09 37.27
 Instruction Len 15.34 63.60 73.73 54.00 15.40 3.60 30.80 36.64
-Alpagasus (Chen et al., 2024d) 21.57 64.37 74.87 55.20 17.65 4.60 16.16 36.34
-Instag-C (Lu et al., 2024) 18.12 64.96 74.01 55.70 15.50 4.80 37.81 38.70
-Instag-D (Lu et al., 2024) 21.94 64.69 74.87 54.80 12.80 4.10 9.76 34.71
-SuperFilter (Li et al., 2024a) 22.95 64.99 76.39 57.60 6.05 2.60 40.55 38.73
-Deita-C (Liu et al., 2024b) 15.58 64.97 74.21 55.00 13.05 4.60 34.46 37.41
-Deita-Q (Liu et al., 2024b) 19.57 64.22 75.15 54.00 7.20 4.20 28.35 36.10
-Deita-D (Liu et al., 2024b) 20.97 63.32 75.10 54.90 7.00 4.00 31.71 36.71
+Alpagasus [10] 21.57 64.37 74.87 55.20 17.65 4.60 16.16 36.34
+Instag-C [30] 18.12 64.96 74.01 55.70 15.50 4.80 37.81 38.70
+Instag-D [30] 21.94 64.69 74.87 54.80 12.80 4.10 9.76 34.71
+SuperFilter [26] 22.95 64.99 76.39 57.60 6.05 2.60 40.55 38.73
+Deita-C [29] 15.58 64.97 74.21 55.00 13.05 4.60 34.46 37.41
+Deita-Q [29] 19.57 64.22 75.15 54.00 7.20 4.20 28.35 36.10
+Deita-D [29] 20.97 63.32 75.10 54.90 7.00 4.00 31.71 36.71
 DaaR (Ours) 20.08 64.55 74.88 54.8 15.30 4.70 37.50 38.83
@@
 Raw 8.03 59.58 73.00 78.00 5.70 5.00 60.98 41.47
-Rand 13.28 58.27 73.00 75.35 35.36 52.20 63.72 53.02
+Full (40K) 15.61 58.75 72.51 73.80 31.30 51.70 67.38 53.01
+Rand (8K) 13.28 58.27 73.00 75.35 35.36 52.20 63.72 53.02
 Instruction Len 8.62 58.44 72.86 73.30 27.05 53.10 63.72 51.01
-Alpagasus (Chen et al., 2024d) 13.67 57.94 73.04 73.90 32.30 51.40 63.41 52.24
-Instag-C (Lu et al., 2024) 9.51 58.50 73.06 74.70 35.35 51.90 64.70 52.53
-Instag-D (Lu et al., 2024) 12.87 57.48 72.80 74.40 33.75 51.80 64.02 52.45
-SuperFilter (Li et al., 2024a) 19.16 58.98 72.99 73.70 30.10 52.40 58.85 52.31
-Deita-C (Liu et al., 2024b) 8.94 58.07 73.06 73.90 35.55 52.90 62.20 52.09
-Deita-Q (Liu et al., 2024b) 14.06 59.07 73.16 75.80 35.50 23.00 58.24 48.40
-Deita-D (Liu et al., 2024b) 16.41 57.80 72.70 76.10 29.05 52.40 64.63 52.73
+Alpagasus [10] 13.67 57.94 73.04 73.90 32.30 51.40 63.41 52.24
+Instag-C [30] 9.51 58.50 73.06 74.70 35.35 51.90 64.70 52.53
+Instag-D [30] 12.87 57.48 72.80 74.40 33.75 51.80 64.02 52.45
+SuperFilter [26] 19.16 58.98 72.99 73.70 30.10 52.40 58.85 52.31
+Deita-C [29] 8.94 58.07 73.06 73.90 35.55 52.90 62.20 52.09
+Deita-Q [29] 14.06 59.07 73.16 75.80 35.50 23.00 58.24 48.40
+Deita-D [29] 16.41 57.80 72.70 76.10 29.05 52.40 64.63 52.73
 DaaR (Ours) 16.88 57.58 73.03 75.40 38.1 52.00 64.94 53.99
@@
 Raw 8.84 58.14 72.75 78.20 9.10 7.40 78.05 44.64
-Rand 11.46 57.85 73.08 78.90 13.15 62.50 71.65 52.65
+Full (40K) 12.88 58.60 72.28 76.80 13.60 62.80 71.04 52.57
+Rand (8K) 11.46 57.85 73.08 78.90 13.15 62.50 71.65 52.65
 Instruction Len 11.34 58.01 72.79 78.00 15.80 62.30 68.12 52.34
-Alpagasus (Chen et al., 2024d) 10.40 57.87 72.92 77.20 18.75 61.80 65.55 52.07
-Instag-C (Lu et al., 2024) 10.81 58.45 73.27 76.00 13.30 61.80 68.29 51.70
-Instag-D (Lu et al., 2024) 11.08 58.40 72.79 76.40 16.40 62.90 70.43 52.63
-SuperFilter (Li et al., 2024a) 13.54 58.51 72.89 79.30 11.35 39.50 65.25 48.62
-Deita-C (Liu et al., 2024b) 10.50 58.17 73.14 74.60 16.60 62.00 72.26 52.47
-Deita-Q (Liu et al., 2024b) 11.24 57.83 72.97 78.50 12.95 38.10 67.68 48.47
-Deita-D (Liu et al., 2024b) 10.48 57.81 73.05 77.20 15.25 52.90 69.21 50.84
+Alpagasus [10] 10.40 57.87 72.92 77.20 18.75 61.80 65.55 52.07
+Instag-C [30] 10.81 58.45 73.27 76.00 13.30 61.80 68.29 51.70
+Instag-D [30] 11.08 58.40 72.79 76.40 16.40 62.90 70.43 52.63
+SuperFilter [26] 13.54 58.51 72.89 79.30 11.35 39.50 65.25 48.62
+Deita-C [29] 10.50 58.17 73.14 74.60 16.60 62.00 72.26 52.47
+Deita-Q [29] 11.24 57.83 72.97 78.50 12.95 38.10 67.68 48.47
+Deita-D [29] 10.48 57.81 73.05 77.20 15.25 52.90 69.21 50.84
 DaaR (Ours) 15.83 58.65 72.48 80.20 16.70 64.20 68.29 53.76

### v3

@@
 We release our code and hope this study can shed light on the understanding of data diversity and advance feedback-driven data-model co-design for LLMs.
-11footnotetext: Equal contribution.
+11footnotetext: Equal contribution.22footnotetext: Corresponding author.33footnotetext: FSIETP: Guangdong Provincial Key Laboratory of Fire Science and Intelligent Emergency Technology.
 1 Introduction
@@
 Fine-tuning these models further improves their usability by aligning their behaviors with specific instructions and human preferences alpaca2023 ; dpo .
-To cultivate comprehensive capabilities in LLMs, prior studies have explored preferable trade-offs between quality, quantity, and diversity of their training data  li2024quantity ; djv2 ; qin2024surveycodev ; zhao2024beyond .
+To cultivate comprehensive capabilities in LLMs, prior studies have explored preferable trade-offs between quality, quantity, and diversity of their training data  li2024quantity ; djv1 ; qin2024surveycodev ; zhao2024beyond .
 For example, methods focusing on data selection li2024super ; xialess ; wang2023far and mixture ge2024bimix demonstrate promising capabilities to enhance model performance, particularly through semantic diversity  lu2023instag ; liu2023deita .
@@
 These findings motivate our theoretical analysis, which establishes that traditional data mixing strategies assuming known domain weights ye2024data cannot achieve Pareto-optimal performance in label-free settings.
-We prove that when domain labels are unavailable, the optimal data selection strategy must simultaneously maximize the mutual information between selected samples and preserve the base model’s latent feature geometry.
+Our analysis further reveals that when domain labels are unavailable, an effective data selection strategy should prioritize samples that enhance sample-level diversity.
+This strategy involves balancing the goal of enhancing sample-level diversity to approximate the ideal importance weights, with the need to preserve the base model’s latent feature geometry.
 We operationalize these insights through DaaR, a self-supervised framework that learns diversity as a reward signal through three technical innovations: (1) Automatic synthesis of model-aware domain centroids via iterative embedding-space generation, (2) A lightweight MLP probe trained to predict semantic entropy using only the LLM’s frozen embeddings, and (3) Closed-loop fine-tuning where the model’s own diversity estimates guide subsequent data selection.
@@
 Crucially, our method requires no domain labels or external models - the LLM self-supervises its diversity estimation through geometric constraints in its native embedding space.
-Our core contributions for LLM data optimization are three-fold:
+Our core contributions for LLM data optimization are:
   • Formal characterization of diversity’s dual role in LLM fine-tuning through controlled experiments with contrastive data pools.
-  • A new method with label-free diversity reward mechanism using embedding-space entropy prediction, theoretically grounded in mutual information maximization.
+  • A new method with label-free diversity reward mechanism using embedding-space entropy prediction, theoretically grounded in importance sampling.
   • Extensive empirical validations showing consistent gains across model families and tasks, with notable improvements in mathematical reasoning (+27%) and coding (+7.4%), whereas other baselines struggle in such challenging scenarios.
@@
 Fine-tuning is a pivotal training paradigm for enhancing LLMs’ domain-specific capabilities.
-Research has shown that a small set of instruction pairs can enable LLMs to follow major instructions effectively zhou2024lima ; djv1 .
+Research has shown that a small set of instruction pairs can enable LLMs to follow major instructions effectively zhou2024lima ; mindgym .
 This fine-tuning can be achieved through rule-based methods, which focus on attributes such as Error L2-Norm paul2021el2n and token length raffel2020exploring .
@@
 Diversity & Data Mixing
-A fundamental principle for LLMs is being able to handle diverse human requests, underscoring data diversity as essential for effective fine-tuning yu2022can ; ding2023enhancing .
+A fundamental principle for LLMs is being able to handle diverse human requests, underscoring data diversity as essential for effective pre-training liu2024regmix ; chen2024aioli ; xie2025chameleon ; fan2024dynamic and fine-tuning yu2022can ; ding2023enhancing .
 Data diversity encompasses aspects such as data deduplication abbas2023semdedup , the coverage scope of tags lu2023instag , model-based diversity evaluation liu2023deita ; zhang2024harnessing , and scaling properties song2024scaling .
-Typically, data mixing approaches albalak2023efficient ; ye2024data ; ge2024bimix focus on adjusting the proportional weights of different domains to enhance model capabilities.
+Typically, data mixing approaches albalak2023efficient ; ye2024data ; ge2024bimix ; liu2024regmix focus on adjusting the proportional weights of different domains to enhance model capabilities.
 Our Position
@@
 $\small\min_{\alpha}\sum_{k=1}^{K}\lambda_{k}\mathbb{E}_{x\sim\mathcal{D}_{k}}[\ell(x;\theta^{*}(\alpha))],\quad\text{s.t.}\ \ \theta^{*}(\alpha)=\arg\min_{\theta}\sum_{i=1}^{N}\alpha_{i}\ell(x_{i};\theta).$ (1)
+All notations used are summarized in Appendix A.
 3 Diversity: A Critical Factor for LLM Fine-Tuning
@@
 Models & Implementation We employ the Qwen2 series (Qwen2-7B & Qwen2.5-7B) qwen2series and the Llama3.1-8B llama3series as representative SOTA base models to be fine-tuned.
-All experiments are conducted under identical training and evaluation protocols with two independent repetitions. Full platform, training and evaluation details are provided in Appendix B.
+All experiments are conducted under identical training and evaluation protocols with two independent repetitions. Full platform, training and evaluation details are provided in Appendix D.
 3.2 Data Pools with Contrastive Distributions
@@
 1 (full-size version is in Appendix, Fig.
-16) demonstrates this continuum through t-SNE projections – high $\Phi_{\text{inter}}$ manifests as distinct cluster separation with clear margins (Fig.
-16.(c)), whereas low values produce entangled distributions (Fig.
-16.(d)).
+17) demonstrates this continuum through t-SNE projections – high $\Phi_{\text{inter}}$ manifests as distinct cluster separation with clear margins (Fig.
+17.(c)), whereas low values produce entangled distributions (Fig.
+17.(d)).
 Full analysis is detailed in Appendix H.1.
@@
 19 and Fig.
-19), this quantization process induces measurable distribution shifts.
+20), this quantization process induces measurable distribution shifts.
 3.3 Experimental Observations
-Table 1 presents comprehensive evaluations across seven benchmarks, where the notation Inter-Diversity (X-Y) indicates samples ranked in the top (100-Y)to (100-X)of cross-domain similarity scores.
+Table 1 presents comprehensive evaluations across seven benchmarks, where the notation Inter-Diversity (X-Y) indicates samples ranked in the top (100-Y)% to (100-X)% of cross-domain similarity scores.
 Due to space constraints, we present only the results for the top 20%, middle 20%, and bottom 20%.
@@
 To bridge this distribution shift, we adopt an importance sampling perspective xie2023data ; xie2023doremi ; xialess where the optimal subset corresponds to reweighting samples by the density ratio $\frac{q(x)}{p(x)}$:
-[…]
+$\small\mathbb{E}_{x\sim q(\cdot)}\ell(x)=\int_{x}q(x)\ell_{\theta}(x)dx=\int_{x}p(x)\frac{q(x)}{p(x)}\ell_{\theta}(x)dx=\mathbb{E}_{x\sim p(\cdot)}\frac{q(x)}{p(x)}\ell(x;\theta)=\sum_{x\sim\mathcal{D}}\frac{q(x)}{p(x)}\ell(x;\theta).$ (4)
 The domains of LLM’s data are commonly defined by textual descriptions, which inherently constrain the semantic scope. Thus we make the following assumption regarding distributions $p$ and $q$:
@@
 $\frac{q(x)}{p(x)}=\sum_{k=1}^{K}\lambda_{k}p(c{=}k|x),\lambda_{k}:=\frac{q(c=k)}{p(c=k)}$ (5)
-The derivation is detailed in Appendix E. This decomposition reveals that the Inter-Diversity and Intra-Diversity are governed by $\lambda_{k}$ and $p(c{=}k|x)$, leading to the following analysis.
+The derivation is detailed in Appendix C.1.
+This decomposition reveals that the Inter-Diversity and Intra-Diversity are governed by $\lambda_{k}$ and p(c=k|x)p(c{=}k|x), leading to the following analysis.
 Why Overlook Diversity is Suboptimal
@@
 This formulation is commonly used when domain labels are available, under the implicit assumption that domains are mutually exclusive.
-However, such an assumption rarely holds in the context of LLM training, as cross-domain data may either synergize or conflict zhao2024beyond ; xie2023doremi (evidence in Table 1).
+However, such an assumption rarely holds in the context of LLM training, as cross-domain data may either synergize or conflict zhao2024beyond ; xie2023doremi ; kuang2025atomic (evidence in Table 1).
 Hence, neglecting diversity in data selection potentially leads to suboptimal overall performance.
+We further formalize this suboptimality by analyzing the approximation error incurred by such a deterministic assignment, which provides a supplementary justification for using predictive entropy as a selection criterion.
+For a detailed analysis, please refer to Appendix C.2.
 Guidance to Our Method
-Under the domain-undetermined scenario, our proposed method DaaR operationalizes the gained insights by (1) constructing pseudo-labels (Section 4.1) align with LLM-aware generated seed, (2) explicitly modeling the classifier $p(c{=}k|x)$ through a domain discrimination probe (Section 4.2), and (3) using the probe’s predictive entropy (Section 4.2) as a proxy for diversity, estimating the ability of classifier $p(c{=}k|x)$ to select data toward the target distribution $q$.
+Under the domain-undetermined scenario, our proposed method DaaR operationalizes the gained insights by (1) constructing pseudo-labels (Section 4.1) align with LLM-aware generated seed, (2) explicitly modeling the classifier p(c=k|x)p(c{=}k|x) through a domain discrimination probe (Section 4.2), and (3) using the probe’s predictive entropy (Section 4.2) as a proxy for diversity, estimating the ability of classifier p(c=k|x)p(c{=}k|x) to select data toward the target distribution $q$.
 4 DaaR: Diversity as a Reward
-To address the challenges identified in Sec.
-3.3 and leverage the insights gained in Sec.
-3.4, we establish a data selection method DaaR guided by diversity-aware reward signals.
+To address the challenges identified in Sec. 3.3 and leverage the insights gained in Sec. 3.4, we establish a data selection method DaaR guided by diversity-aware reward signals.
 It comprises three key components illustrated in Fig.
-1: (1) model-aware centroid synthesis, (2) two-stage training with reward probe, and (3) diversity-driven data selection.
+1: (1) model-aware centroid synthesis, which generates domain-representative centroids and seed data capturing the LLM’s intrinsic feature space, (2) two-stage training with reward probe, which yields a probe module capable of predicting sample-level diversity rewards accurately, and (3) diversity-driven data selection to obtain a data subset that effectively boosts LLM’s cross-domain performance.
 4.1 Model-Aware Training Data
@@
   • Phase 1 - Seed Generation: For each domain $k$, generate seed samples $\mathcal{S}_{k}^{(0)}$ via zero-shot prompting with domain-specific description templates generated by LLM itself, alongside with minor injection from downstream task samples, establishing initial semantic anchors.
-  • Phase 2 - Diversity Augmentation: Iteratively expand $S_{k}^{(t)}$ through context-aware generation, conditioned on a sliding window buffer with 3 random anchors sampled from the $(t-1)$ iteration $S_{k}^{(t-1)}$.
+We show that removing the minor injection of downstream samples negligibly impacts final performance but significantly reduces data generation efficiency in Appendix G.8.
+  • Phase 2 - Diversity Augmentation: Iteratively expand $S_{k}^{(t)}$ through context-aware generation, conditioned on a sliding window buffer with 3 random anchors sampled from the (t1)(t-1) iteration $S_{k}^{(t-1)}$.
 The generated sample $\mathbf{x^{\prime}}$ is retained through rejection sampling to enhance diversity.:
@@
 This process terminates when it reaches the predetermined iteration.
-The implementation of all parameters, along with their corresponding ablation studies, is provided in Appendix C.1.
 The domain centroid is then computed from the final augmented set $\mathcal{S}_{k}$ using the model’s embedding, formalized as $\mathcal{C}_{k}=1/|\mathcal{S}_{k}|\cdot\sum_{\mathbf{x}_{i}\in\mathcal{S}_{k}}\mathcal{M}_{\text{ebd}}(\mathbf{x}_{i})$.
 This captures the LLM’s intrinsic feature space geometry while eliminating dependency on human annotations.
+More implementation details including hyper-parameters and the data generation process, along with their corresponding ablation studies, are provided in Appendix B.
 Domain-Aware Clustering
@@
 Stage 1: Domain Predictor
-The proposed DaaR establishes model-aware domain discrimination abilities through a multi-layer perceptron (MLP) probe module, $\psi_{\text{dom}}$, attached to the hidden layer $\mathcal{M}(\tilde{x})$ of the LLMs.
-The probe will be trained meanwhile all the parameters of the LLM are frozen, achieving a preferable balance between effectiveness and cost, with detailed analysis regarding the choice of layers presented in Appendix G.2.
-Specifically, with pseudo-label $\tilde{y}$, we can compute domain predicted probabilities as:
+[…]
 $p_{k}(\tilde{x})=\text{softmax}\left(\psi_{\text{dom}}(\mathcal{M}(\tilde{x}))\right),\quad\psi_{\text{dom}}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{K},$ (8)
-$\psi_{\text{dom}}$ is optimized via cross-entropy loss $\mathcal{L}_{\text{dom}}=-\frac{1}{|\mathcal{D}_{\text{probe}}|}\sum_{(\tilde{x},\tilde{y})\in\mathcal{D}_{\text{probe}}}\sum_{k=1}^{K}\mathbb{I}_{[k=\tilde{y}]}\log p_{k}(\tilde{x})$, where $\mathbb{I}_{[k=\tilde{y}]}$ denotes the indicator function.
-We employ single-sample batches with the AdamW optimizer to prevent gradient averaging across domains.
-Training consistently converges and achieves 92.7% validation accuracy on domain classification, as shown in Fig.
-2 (a).
+[…]
 Stage 2: Diversity Rewarding
-Building on the stabilized domain discrimination probe, we quantify sample-level diversity through predictive entropy $H(\tilde{x})$.
-And to enable efficient reward computation during data selection, we then train another 5-layer MLP $\psi_{\text{div}}$ to directly estimate $H(\tilde{x})$ from $\mathcal{M}(\tilde{x})$:
+[…]
 $H(\tilde{x})=-\sum_{k=1}^{K}p_{k}(\tilde{x})\log p_{k}(\tilde{x}).\quad\hat{H}(\tilde{x})=\psi_{\text{div}}(\mathcal{M}(\tilde{x})),\psi_{\text{div}}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{+}.$ (9)
@@
 We use the following data selection methods for comprehensive evaluation: (1) Random Selection: traditional random sampling; (2) Instruction Len: measuring instruction complexity by token count  cao2023instruction ; (3) Alpagasus chen2024alpagasus : using ChatGPT for direct quality scoring of instruction pairs; (4-5) Instag lu2023instag : semantic analysis approach with Instag-C (complexity scoring via tag quantity) and Instag-D (diversity measurement through tag set expansion); (6) SuperFilter li2024super : response-loss-based complexity estimation using compact models; (7-9) Deita liu2023deita : model-driven evaluation with Deita-C (complexity scoring), Deita-Q (quality scoring), and Deita-D (diversity-aware selection).
-Detailed implementations for these baselines are in Appendix D.1.
+Detailed implementations for these baselines are in Appendix E.1.
 Table 2: Evaluation results of Llama3.1-8B, Qwen2-7B, and Qwen2.5-7B across various downstream task benchmarks. DaaR demonstrates superiority in AVG compared to other baselines.
@@
 In the Qwen series, even all baselines fall below the Random performance.
-We conjecture this stems from selected distribution-bias, as the baselines are unable to balance the proportions across different domains, a perspective visually supported and analyzed in Appendix D.2.
-Comprehensive Performance
-: DaaR establishes new SOTA averages across all models, surpassing the best baselines by +0.14 (Llama3.1), +0.97 (Qwen2), and +1.11 (Qwen2.5).
-The proposed method uniquely achieves dual optimization in critical capabilities: Mathematical Reasoning: Scores 38.1 MATH (Qwen2) and 16.70 MATH (Qwen2.5), with 7.4% and 27.0% higher than respective random baselines.
-Coding Proficiency: Maintains 64.94 HumanEval (Qwen2) and 64.20 MBPP (Qwen2.5) accuracy with <1% degradation from peak performance.
-This demonstrates DaaR’s ability to enhance challenging STEM capabilities while preserving core competencies.
-5.2 Generalizability and Uniqueness of DaaR
-Evaluation on MMLU In the main experiments, the data and benchmarks we employed are meticulously aligned, which validates the In-Distribution capabilities of DaaR.
-To further validate the generalization ability of DaaR, we present MMLU mmlu2021 evaluations in Appendix F.1.
-MMLU encompasses a significantly broader range of content than our dataset, thus serving as an Out-of-Distribution scenario.
-Results demonstrate that DaaR achieved top accuracy on Qwen2 with an improvement of 0.61 points, and secured a competitive second-place performance on Qwen2.5.
-At a granular level, DaaR particularly excels in the STEM domain, outperforming baselines by an average of 7.6% and 1.7% on Qwen2 and Qwen2.5, which aligns the findings above.
-Evaluation on Qwen3 In Appendix F.2, we explore the performance of DaaR on the mainstream RL-based LLM, Qwen3-8B.
-We observe that despite difference in structure, the phenomena (Fig.
-8 & Table 8) and the dynamics of training (Fig.
-9) remain consistent with the main observations.
-Then we compare DaaR with baselines in Table 9, although Qwen3-8B is not directly designed for SFT paradigm, DaaR maintains the top position in Avg performance across baselines, surpasses all baselines up to 4.7% and notably exceeds the avg of baselines by 25.91% on MATH.
-Customized Characteristics Unlike baselines, DaaR enables domain-aware data selection in undetermined scenario, showcasing potential customization on $\lambda_{k}$.
-As detailed in Appendix F.3, by adjusting the selection ratio for specific domains, we achieve SOTA performance in targeted domain without compromising overall ability, promising practical utility in real-world application.
-Table 3: Illustration results on MMLU, Qwen3 and customization, detailed in Table 6, 9, 10
+We conjecture this stems from selected distribution-bias, as the baselines are unable to balance the proportions across different domains, a perspective visually supported and analyzed in Appendix E.2.
+Table 3: Illustration results on MMLU, Qwen3 and customization, detailed in Table 7, 10, 11
 (a) MMLU Performance
@@
 Reason 16.56 74.62 53.39
+Comprehensive Performance
+: DaaR establishes new SOTA averages across all models, surpassing the best baselines by +0.14 (Llama3.1), +0.97 (Qwen2), and +1.11 (Qwen2.5).
+The proposed method uniquely achieves dual optimization in critical capabilities: Mathematical Reasoning: Scores 38.1 MATH (Qwen2) and 16.70 MATH (Qwen2.5), with 7.4% and 27.0% higher than respective random baselines.
+Coding Proficiency: Maintains 64.94 HumanEval (Qwen2) and 64.20 MBPP (Qwen2.5) accuracy with <1% degradation from peak performance.
+This demonstrates DaaR’s ability to enhance challenging STEM capabilities while preserving core competencies.
+5.2 Generalizability, Uniqueness and Stability of DaaR
+Evaluation on MMLU
+In the main experiments, the data and benchmarks we employed are meticulously aligned, which validates the In-Distribution capabilities of DaaR.
+To further validate the generalization ability of DaaR, we present MMLU mmlu2021 evaluations in Appendix F.1.
+MMLU encompasses a significantly broader range of content than our dataset, thus serving as an Out-of-Distribution scenario.
+Results demonstrate that DaaR achieved top accuracy on Qwen2 with an improvement of 0.61 points, and secured a competitive second-place performance on Qwen2.5.
+At a granular level, DaaR particularly excels in the STEM domain, outperforming baselines by an average of 7.6% and 1.7% on Qwen2 and Qwen2.5, which aligns the findings above.
+Evaluation on Qwen3
+In Appendix F.2, we explore the performance of DaaR on the mainstream RL-based LLM, Qwen3-8B.
+We observe that despite difference in structure, the phenomena (Fig.
+9 & Table 9) and the dynamics of training (Fig.
+10) remain consistent with the main observations.
+Then we compare DaaR with baselines in Table 10, although Qwen3-8B is not directly designed for SFT paradigm, DaaR maintains the top position in Avg performance across baselines, surpasses all baselines up to 4.7% and notably exceeds the avg of baselines by 25.91% on MATH.
+Customized Characteristics
+Unlike baselines, DaaR enables domain-aware data selection in undetermined scenario, showcasing potential customization on $\lambda_{k}$.
+As detailed in Appendix F.3, by adjusting the selection ratio for specific domains, we achieve SOTA performance in targeted domain without compromising overall ability, promising practical utility in real-world application.
+Table 4: DaaR using model-aware pseudo-labels vs. Ground-Truth (GT) labels.
+Model Setting NQ TriviaQA Hellaswag GSM8k MATH MBPP HumanEval Avg
+Llama3.1-8B DaaR w/ Pseudo 20.08 64.55 74.88 54.80 15.30 4.70 37.50 38.83
+DaaR w/ GT 22.54 65.52 73.43 53.80 14.35 4.40 36.50 38.65
+Diff +2.46 +0.97 -1.45 -1.00 -0.95 -0.30 -1.00 -0.18
+Qwen2-7B DaaR w/ Pseudo 16.88 57.58 73.03 75.40 38.10 52.00 64.94 53.99
+DaaR w/ GT 18.75 58.35 72.02 73.40 35.75 51.50 64.01 53.40
+Diff +1.87 +0.77 -1.01 -2.00 -2.35 -0.50 -0.93 -0.59
+Qwen2.5-7B DaaR w/ Pseudo 15.83 58.65 72.48 80.20 16.70 64.20 68.29 53.76
+DaaR w/ GT 16.04 58.75 71.87 79.40 16.30 64.25 67.76 53.48
+Diff +0.21 +0.10 -0.61 -0.80 -0.40 +0.05 -0.53 -0.28
+Robustness and Stability
+We validate our model-aware design choices that relying on external ground-truth labels can degrade performance by up to 0.59 points in Table 4 and Appendix F.4.
+Furthermore, we demonstrate high end-to-end robustness, with independent runs yielding inter-centroid similarities of >0.98 and a final performance standard deviation of only 0.08 in Appendix F.5
 5.3 Ablation Studies and Discussions
-Distinct Domain Centroid As shown in Appendix G.1, DaaR can generate domain-representative samples with clear distinctions.
-Notably, the diversity is consistently most pronounced between common sense, reasoning, and coding domains across model architectures and parameters.
-Impacts of Hyper-Parameters In Appendix G.2-Appendix G.6, we conduct a series of ablation studies, including (1) selection of layer: the discrimination accuracy derived from layer 1-10 fluctuates in 91.2-93.6%; (2) the number of seed: the similarity between centroids remains highly consistent with variance ranging from 5.4E-7 to 1.8E-5; (3) the number of augmented data: the mean similarity of centroids generated from sizes of 10 and 30 is 0.983 and 0.981 on Qwen2.5 and Llama3.1 respectively; (4) the size of sliding window & similarity threshold: the impact of different hyper-parameters on centroid generation is minimal, ensuring the subsequent training stability of DaaR.
-Cost-Efficiency and Flexibility As shown in Appendix G.7, our method demonstrates significant computational efficiency compared to baselines that rely on GPT-based evaluators or full-LLM inference.
-By operating on frozen hidden layers and generating regression scores instead of full vocabulary distributions, our approach achieves 70% lower GPU usage and 2.5x faster inference.
-6 Conclusions, Limitations & Future Works
+Figure 3: Semantic cosine similarity across different domains for generated samples of size 10.
+Ablation Studies on Centroid Construction
+As shown in Fig.3 and Appendix G.1, DaaR can generate domain-representative samples with clear distinctions.
+Notably, the diversity is consistently most pronounced between common sense, reasoning, and coding domains across model architectures and parameters, showing that despite differences in model architecture across LLMs, their discriminability between these domains exhibit remarkable similarity.
+Impacts of Hyper-Parameters
+In Appendix G.2-Appendix G.7, we conduct a series of ablation studies, including (1) selection of layer: the discrimination accuracy derived from layer 1-10 fluctuates in 91.2-93.6%; (2) the number of seed: the similarity between centroids remains highly consistent with variance ranging from 5.4E-7 to 1.8E-5; (3) the number of augmented data: the mean similarity of centroids generated from sizes of 10 and 30 is 0.983 and 0.981 on Qwen2.5 and Llama3.1 respectively; (4) the size of sliding window & similarity threshold: the impact of different hyper-parameters on centroid generation is minimal, ensuring the subsequent training stability of DaaR; (5) diversity threshold, where varying thresholds minimally affect centroid-guided clustering (with distinct generation attempt frequencies across thresholds) and empirically balanced thresholds are determined for different domains; (6) generalizability across model scales, where DaaR shows consistent improvements in representative scenario on both Llama3.2-3B and Qwen2.5-14B models.
+Cost-Efficiency and Flexibility
+As detailed in Appendix G.9, our method demonstrates significant computational efficiency compared to baselines that rely on GPT-based evaluators or full-LLM inference.
+By operating on frozen hidden layers and generating regression scores instead of full vocabulary, our approach achieves 70% lower GPU usage and 2.5x faster inference.
+6 Conclusion, Limitation & Future Works
 In this paper, we propose a new approach to fine-tuning LLMs with data that lacks clear domain labels, using diversity as a guiding principle.
@@
 Our experiments with various SOTA LLMs show notable superiority of the method over competitive baselines, highlighting the potential of data diversity to enhance model overall performance.
-There are several limitations to our study.
-Regarding the implementation of DaaR, the customization of selection ratios holds potential for broader application in real-world scenarios.
-Future work could include developing methods to efficiently adjust diversity measures and learning algorithms, thus supporting LLMs in self-evolving towards artificial general intelligence in dynamic environments.
+The implementation of DaaR has limitations that can be addressed further, such as the customization of the selection ratios to support broader application in real-world scenarios.
+Besides, DaaR has potential to efficiently adjust diversity measures, benefiting LLMs in self-evolving towards artificial general intelligence in dynamic environments trinity-rft .
+Acknowledgment
+This research is supported by Key-Area Research and Development Program of Guangdong Province, Granted No.
+2024B1111060004 and Guangdong Provincial Special Funds for Promoting High Quality Economic Development (Marine Economic Development) in Six Major Marine Industries, Granted No.
+GDNRC[2024]52.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4.
+Justification: The abstract and introduction appropriately outline the paper’s contributions and scope, presenting clear and accurate claims that align with the theoretical and experimental results.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: Section 6 discusses the limitations of our approach and outlines potential directions for future work to extend our findings.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes]
+  14. Justification: We follow the requirements and instructions of the theoretical analysis.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: We fully disclose all our implementation details in Appendix B & D for reproducibility.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: We fully release our implementation code via an anonymous link during the submission stage, and provide a public link in the camera-ready version.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: We fully disclose all our implementation details in Appendix B & D, alongside with several ablation study of hyper-parameters in Appendix F.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34. Justification: We independently replaced two seeds and used the average of the results as our final outcome in main experiments.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: We provide sufficient information on the computer resources in Appendix D.3.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: We fully align with the Code of Ethics and make sure our code is anonymous during submission period.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [N/A]
+  49. Justification: There is no societal impact of the work performed.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: The paper poses no such risks.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: We document the sources for all assets, including LLMs, datasets, training / evaluation platforms, and baselines.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets ·
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [Yes]
+  64. Justification: The released code is well documented in a anonymous url during submission period.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [N/A]
+  79. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
 Appendix
-Our appendix is organized as follows:
-  • Sec. A: The symbols used in this paper.
-  • Sec. B: The details about experimental setup, including data pool construction (Sec. B.1), implementation of embedding (Sec. B.2), benchmarks (Sec. B.3) and train / eval details (Sec. B.4).
-  • Sec. C: The implementation details of DaaR, including hyper-parameter setting (Sec. C.1) and the process of model-aware training data generation (Sec. C.2).
-  • Sec. D: The details of utilized baselines, including the implementation of baselines (Sec. D.1) and visualization of the data selected by baselines (Sec. D.2).
-  • Sec. E: The detailed derivation of Proposition 3.2.
-  • Sec.
-F: More comprehensive experiments about DaaR, including the evaluation on MMLU (Sec.
-F.1), the validation on Qwen3 (Sec.
-F.2), the customized selection (Sec.
-F.3) and the utilization of entropy (Sec.
-F.4).
-  • Sec.
-G: Complete ablation studies and discussions of DaaR, including the centroid similarity (Sec.
-G.1), LLM’s layer selection (Sec.
-G.2).
-As well as several ablation studies of the number of seeds (Sec.
-G.3), the number of diversity augmentation data (Sec.
-G.4), the size of sliding window (Sec.
-G.5) and diversity threshold (Sec.
-G.6).
-Additionally, we verify the cost-efficiency (Sec.
-G.7) and the stability of DaaR (Sec.
-G.8).
-  • Sec H: The comprehensive experimental results, including the visualization results of embeddings on Qwen2-7B, Llama3.1-8B & Qwen2.5-7B (Sec H.1), the visualization of inter- & intra-diversity distributions (Sec H.2), the comprehensive training dynamics (Sec.
-H.3), the validation of inter-diversity and intra-diversity (Sec.
-H.4) and complete results of DaaR with baselines (Sec.
-H.5).
 Appendix A Notation
@@
 $\psi_{\text{dom}}$ A MLP-based domain predictor $\psi_{\text{div}}$ A MLP-based entropy predictor
-$p_{k}(\tilde{x})$ Domain predicted probability by $\psi_{\text{dom}}$ $\hat{H}(\mathbf{x})$ Predictive entropy by $\psi_{\text{dom}}$
+[…]
 $\mathcal{L}_{\text{dom}}$ Cross-entropy loss of $\psi_{\text{dom}}$ $\mathcal{L}_{\text{div}}$ MSE loss of $\psi_{\text{dom}}$
-Appendix B Experimental Setup Details
-B.1 Construction of Data Pool
+Appendix B Implementation Details of DaaR
+B.1 Hyper-Parameters
+For the generation of synthetic data, we utilize the model.generate function from the PyTorch library.
+Key parameters included setting max-new-tokens to 2048 to control the length of the output, enabling do-sample to True to allow sampling, and configuring top-p to 0.95 and temperature to 0.9 to ensure diversity in the generated content.
+During the content extraction phase, we employ regular expressions to efficiently extract and structure the desired information.
+[…]
+B.2 Implementation of Embedding
+Given the Alpaca-based [43] data format employed in our study, we concatenate the ‘instruction’, ‘input’, and ‘output’ components of each data sample to capture domain-aware semantic representations.
+These concatenated sequences are then fed into the LLM, where we compute the average of all token embeddings extracted from the embedding layer to derive the semantic vector characterizing each sample.
+The embedding layer is deliberately selected due to its shallow architectural position: (1) it preserves precise semantic information capture through raw input representations, and (2) it incurs reduced computational overhead compared to deeper layers.
+As evidenced by the visualization in Fig.
+13 (Appendix G.2), this approach demonstrates robust domain discrimination capabilities across different LLM layers, thereby validating the robustness of our methodology.
+B.3 The Phase of Model-Aware Training Data Generation
+Seed Generation
+This phase aims to provide zero-shot LLM-generated seed data that reflects how LLMs inherently perceive the characteristics of domain-specific data.
+We employ the LLM itself to generate concise domain descriptions across diverse fields illustrated in Prompt B.3, and the prompt shown in Prompt B.3.
+Due to the difficulty of pre-trained LLMs in providing accurate and coherent responses, we utilize their corresponding Instruct versions for centroid data generation.
+Specifically, we apply Llama3.1-8B-Instruct for Llama3.1-8B, Qwen2-7B-Instruct for Qwen2-7B, and Qwen2.5-7B-Instruct for Qwen2.5-7B.
+Notably, these descriptions are not required to be optimal or prescriptive, rather, they serve as representative examples compatible with our experimental scenarios.
+Users retain the flexibility to customize target domain specifications according to specific requirements.
+The associated stability analysis regarding this design choice will be elaborated in Appendix G.10.
+Diversity Augmentation
+While seed data effectively captures the intrinsic domain knowledge of LLMs, its limited quantity and the constrained diversity of single-prompt generation methods necessitate enhancement.
+We address these limitations through two complementary strategies including a sliding window mechanism and a rule-based filter.
+The implementation details are formalized in Prompt B.3 and Eq.
+(7).
+Domain-Aware Clustering
+During reward model training data clustering, we rigorously curate a subset of 5,000 entries that are distribution-matched to the fine-tuning dataset while separate from it.
+Common Sense: Common sense generally includes a knowledge-based question and its corresponding answer, without reasoning.
+Reasoning: Reasoning involves the ability to think logically about a situation or problem, to draw conclusions from available information, and to apply knowledge in new situations.
+Mathematics: Mathematical skills include the ability to perform calculations, understand mathematical concepts, solve hard and professional math problems, and apply mathematical reasoning.
+Coding: Design and generate specific code programs, or apply algorithms and data structures, with code generation in the Output.
+You are an AI model with expertise in {selected_domain}.
+Here’s a brief description of this domain: {Prompt 1} Generate 5 different instruction pairs related to this field with various lengths.
+Maintain the index format: Instruction [1 … 5].
+The response should include three parts:
+1. Instruction: A clear command or question that can be understood by the assistant.
+2.
+Input: Any information provided to help it understand the instruction.
+If there is no need to generate, just keep it empty.
+3.
+Output: The expected answer or action.
+Keep the generated content focused on {selected_domain}.
+And do not involve {unselected_domains} related knowledge.
+You are an AI model with expertise in {selected_domain}. Here’s a brief description of this domain:{Prompt 1}
+Generate only an instruction pair related to this field. The response should include three parts:
+Instruction: A clear command or question that can be understood by the assistant.
+Input: Any information provided to help it understand the instruction. If there is no need to generate, just keep empty.
+Output: The expected answer or action.
+Keep the generated content focused on {selected_domain}.
+Do not involve {unselected_domain} related knowledge.
+Note that you should generate content strongly unrelated and different to these examples to ensure diversity in the generated output:
+Counterexample: {}
+The format of the generated content should be: Instruction: [], Input: [], Output: [].
+Appendix C Supplementary Theoretical Analysis
+This appendix provides additional details on the theoretical foundation of our work, as discussed in Section 3.4.
+We first present the full derivation of the importance weight decomposition (Proposition 3.2) and then offer an analysis justifying our entropy-based selection strategy.
+C.1 Derivation of Proposition 3.2
+We derive the density ratio by systematically expanding and substituting terms:
+$\frac{q(\mathbf{x})}{p(\mathbf{x})}=\frac{\sum_{k=1}^{K}q(c=k)q(\mathbf{x}|c=k)}{p(\mathbf{x})}$ (10)
+Expand $q(\mathbf{x})$ via the law of total probability over domains.
+$\frac{\sum_{k=1}^{K}q(c=k)p(\mathbf{x}|c=k)}{p(\mathbf{x})}$ (11)
+Apply Assumption 3.1 ($q(\mathbf{x}|c=k)=p(\mathbf{x}|c=k)$) to unify domain-conditional densities.
+$\frac{\sum_{k=1}^{K}q(c=k)\frac{p(c=k|\mathbf{x})p(\mathbf{x})}{p(c=k)}}{p(\mathbf{x})}$ (12)
+Express $p(\mathbf{x}|c=k)$ via Bayes’ rule inversion: $p(\mathbf{x}|c=k)=\frac{p(c=k|\mathbf{x})p(\mathbf{x})}{p(c=k)}$.
+$\sum_{k=1}^{K}\frac{q(c=k)}{p(c=k)}p(c=k|\mathbf{x})\cdot\frac{p(\mathbf{x})}{p(\mathbf{x})}$ (13)
+Factor out $p(\mathbf{x})$ from the numerator and the denominator.
+$\sum_{k=1}^{K}\underbrace{\frac{q(c=k)}{p(c=k)}}_{\lambda_{k}}p(c=k|\mathbf{x})$ (14)
+Define domain proportion ratio $\lambda_{k}:=\frac{q(c=k)}{p(c=k)}$.
+This demonstrates that the density ratio decomposes linearly into domain-specific components weighted by their proportion shifts ($\lambda_{k}$) and posterior probabilities ($p(c=k|\mathbf{x})$).
+C.2 Supplementary Analysis of the Theoretical Foundation
+Our theoretical framework motivates a data selection strategy that privileges high-entropy samples.
+Here, we theoretically analyze why this is not merely a heuristic to move away from a known suboptimal point, but rather a principled approach to minimize approximation error.
+Approximation Error of Deterministic Assignment
+As shown in Section 3.4, a simplified selection strategy might rely on a deterministic domain assignment, where only the most likely domain $k^{*}(x)=\arg\max_{k}p(c=k|x)$ is considered.
+This reduces the true importance weight $w(x)=\sum_{k=1}^{K}\lambda_{k}p(c=k|x)$ to an approximation $w_{\text{approx}}(x)=\lambda_{k^{*}(x)}$.
+We analyze the squared error of this approximation, $\text{Err}(x)=(w(x)-w_{\text{approx}}(x))^{2}$.
+Proposition C.1.
+The squared approximation error $\text{Err}(x)$ can be measured by:
+$\text{Err}(x)=\left(\sum_{j\neq k^{*}(x)}(\lambda_{j}-\lambda_{k^{*}(x)})\cdot p(c=j|x)\right)^{2}$ (15)
+Implication for Data Selection
+This proposition reveals that the approximation error is a direct function of the probability mass distributed across non-dominant domains.
+The error is zero only if $p(c=k^{*}(x)|x)=1$, which corresponds to zero conditional entropy, $H(C|X=x)=0$.
+Conversely, the error is maximized precisely when the probability is spread across multiple domains—the very definition of high predictive entropy.
+Therefore, our strategy of selecting high-entropy samples is a principled approach to preferentially select samples where the deterministic approximation is most erroneous.
+These are exactly the samples for which the full diversity information encoded in $p(c|x)$ is most critical for an accurate estimation of the true importance weight $w(x)$.
+Our method, DaaR, employs a dedicated domain discrimination probe to model $p(c|x)$ and uses its predictive entropy as a computationally feasible proxy to identify these high-error, high-value samples for training.
+Appendix D Experimental Setup Details
+D.1 Construction of Data Pool
 To investigate data selection from large data pools and its impact on the mixture of downstream tasks, we construct a data pool with distinct properties to mimic practical settings.
@@
 Specifically, for the Math-Instruct dataset, due to its inclusion of CoT and certain coding capabilities, we extract a highly mathematics-related subset and use regular expressions to filter out the coding-related content (including ’program’, ’python’, ’def’, ’import’, ’print’, ’return’), ensuring it remains within the domain of mathematics.
-B.2 Implementation of Embedding
-Given the Alpaca-based [38] data format employed in our study, we concatenate the ’instruction’, ’input’, and ’output’ components of each data sample to capture domain-aware semantic representations.
-These concatenated sequences are then fed into the LLM, where we compute the average of all token embeddings extracted from the embedding layer to derive the semantic vector characterizing each sample.
-The embedding layer is deliberately selected due to its shallow architectural position: (1) it preserves precise semantic information capture through raw input representations, and (2) it incurs reduced computational overhead compared to deeper layers.
-As evidenced by the visualization in Fig.
-12 (Appendix G.2), this approach demonstrates robust domain discrimination capabilities across different LLM layers, thereby validating the robustness of our methodology.
-B.3 Benchmarks
+D.2 Benchmarks
 To evaluate the models’ true capabilities and performance across different domains, we follow the approach of major open-source LLMs (e.g., the Llama3 series [18] and Qwen2 series [53]) and select the following widely used evaluation sets.
@@
 Utilized 3,610 5,000 10,042 500 1,000 500 164
-B.4 Training and Evaluation Details
+D.3 Training and Evaluation Details
 Platform
@@
 Experiments are conducted on a computing platform equipped with four NVIDIA A100 GPUs (40GB), with pre-trained LLMs loaded as 16-bit floating-point numbers.
-The specific data-model development processes are completed in Data-Juicer Sandbox [9], via integration with the ms-swift [54] training repository, and the OpenCompass [15] evaluation repository.
+The specific data-model development processes are completed in Data-Juicer Sandbox [9, 8], via integration with the ms-swift [61] training repository, and the OpenCompass [16] evaluation repository.
 Training Details
@@
 For most tasks, we employ the gen mode, while for the Hellaswag task, we opt for the ppl mode to better assess perplexity.
-Appendix C Implementation of DaaR
-C.1 Implementation of the Synthetic Data Generation
-For the generation of synthetic data, we utilize the model.generate function from the PyTorch library.
-Key parameters included setting max-new-tokens to 2048 to control the length of the output, enabling do-sample to True to allow sampling, and configuring top-p to 0.95 and temperature to 0.9 to ensure diversity in the generated content.
-During the content extraction phase, we employ regular expressions to efficiently extract and structure the desired information.
-For implementation details of DaaR’s hyper-parameters, we set the number of seed samples $\mathcal{S}_{k}^{(0)}$ as 5 (ablation in Appendix G.3), the number of anchors in sliding window as 3 (ablation in Appendix G.5), the similarity threshold $\tau$ as 0.9 for Mathematics and 0.85 for others (ablation in Appendix G.6), the terminated iteration as 30 (ablation in Appendix G.4), the hidden layer $\mathcal{M}(\tilde{x})$ as layer-3 (ablation in Appendix G.2).
-C.2 Implementation of Model-Aware Training Data Generation
-Seed Generation
-This phase aims to provide zero-shot LLM-generated seed data that reflects how LLMs inherently perceive the characteristics of domain-specific data.
-We employ the LLM itself to generate concise domain descriptions across diverse fields illustrated in Prompt C.2, and the prompt shown in Prompt C.2.
-Due to the difficulty of pre-trained LLMs in providing accurate and coherent responses, we utilize their corresponding Instruct versions for centroid data generation.
-Specifically, we apply Llama3.1-8B-Instruct for Llama3.1-8B, Qwen2-7B-Instruct for Qwen2-7B, and Qwen2.5-7B-Instruct for Qwen2.5-7B.
-Notably, these descriptions are not required to be optimal or prescriptive, rather, they serve as representative examples compatible with our experimental scenarios.
-Users retain the flexibility to customize target domain specifications according to specific requirements.
-The associated stability analysis regarding this design choice will be elaborated in Appendix G.8.
-Diversity Augmentation
-While seed data effectively captures the intrinsic domain knowledge of LLMs, its limited quantity and the constrained diversity of single-prompt generation methods necessitate enhancement.
-We address these limitations through two complementary strategies including a sliding window mechanism and a rule-based filter.
-The implementation details are formalized in Prompt C.2 and Eq.
-(7).
-Domain-Aware Clustering
-During reward model training data clustering, we rigorously curate a subset of 5,000 entries that are distribution-matched to the fine-tuning dataset while separate from it.
-[…]
-[…]
-Reasoning: Reasoning involves the ability to think logically about a situation or problem, to draw conclusions from available information, and to apply knowledge in new situations.
-Mathematics: Mathematical skills include the ability to perform calculations, understand mathematical concepts, solve hard and professional math problems, and apply mathematical reasoning.
-Coding: Design and generate specific code programs, or apply algorithms and data structures, with code generation in the Output.
-[…]
-[…]
-Generate 5 different instruction pairs related to this field with various lengths. Maintain the index format: Instruction [1 … 5].
-The response should include three parts:
-1. Instruction: A clear command or question that can be understood by the assistant.
-2. Input: Any information provided to help it understand the instruction. If there is no need to generate, just keep it empty.
-3. Output: The expected answer or action.
-Keep the generated content focused on {selected_domain}. And do not involve {unselected_domains} related knowledge.
-[…]
-[…]
-Generate only an instruction pair related to this field. The response should include three parts:
-Instruction: A clear command or question that can be understood by the assistant.
-Input: Any information provided to help it understand the instruction. If there is no need to generate, just keep empty.
-Output: The expected answer or action.
-Keep the generated content focused on {selected_domain}. Do not involve {unselected_domain} related knowledge.
-Note that you should generate content strongly unrelated and different to these examples to ensure diversity in the generated output:
-Counterexample: {}
-The format of the generated content should be: Instruction: [], Input: [], Output: [].
-Appendix D Details of Baselines
-D.1 Implementation of Baselines
+Appendix E Details of Baselines
+E.1 Implementation of Baselines
 We select the following representative methods and works related to data selection as our baselines to evaluate their performance in the context of a mixture of downstream tasks.
-Random Selection(Rand)
-A random selection of 8,000 data samples from the data pool was made, which to some extent reflects the distribution characteristics of the original data pool.
-Instruction Length (IL)
-The length of the instruction can be considered a measure of input complexity.
+Random Selection(Rand) A random selection of 8,000 data samples from the data pool was made, which to some extent reflects the distribution characteristics of the original data pool.
+Instruction Length (IL) The length of the instruction can be considered a measure of input complexity.
 It is widely believed [5, 60] that more complex data is beneficial for enhancing model capabilities.
 Therefore, we select a subset of 8,000 entries with the maximum number of words (based on spaces) in the concatenation of Instruction and Input as part of the filtering process.
-AlpaGasus [10]
-Using prompts to directly score and annotate the quality of the data leverages the powerful cognitive abilities of LLMs for evaluation and selection.
+AlpaGasus [10] Using prompts to directly score and annotate the quality of the data leverages the powerful cognitive abilities of LLMs for evaluation and selection.
 Based on the original work, we use the GPT-3.5-Turbo API to score the data with the following prompt:
-[…]
-[…]
+System Prompt:
 We would like to request your feedback on the performance of AI assistant in response to the instruction and the given input displayed following.
@@
 The distribution of direct ratings for the 40,000 data pool is shown in Fig.
-3.
+4.
 Although the original paper’s prompts were strictly followed and efforts were made to minimize potential bias, most scores still clustered around 5.
@@
 Figure 4: The distribution of the score by Alpagasus.
-Instag [30]
-first utilizes ChatGPT to tag the samples based on semantics and intentions, then trains a LLaMA-based tagger on the ChatGPT tags to tag data.
+Instag [34] first utilizes ChatGPT to tag the samples based on semantics and intentions, then trains a LLaMA-based tagger on the ChatGPT tags to tag data.
 They use the number of tags as a proxy for complexity.
@@
 Following the original paper, the prompt is as follows.
-[…]
-[…]
+System Prompt:
 You are a tagging system that provides useful tags for instruction intentions to distinguish instructions for a helpful AI assistant. Below is an instruction:
@@
 This approach captures a broader range of unique tags, thereby enhancing the diversity and representativeness of the dataset.
-SuperFilter [26]
-introduces a method that astonishingly utilizes a small GPT2 model to successfully filter out the high-quality subset from the existing GPT4-generated instruction tuning dataset.
+SuperFilter [29] introduces a method that astonishingly utilizes a small GPT2 model to successfully filter out the high-quality subset from the existing GPT4-generated instruction tuning dataset.
 The core concept is the instruction-following difficulty (IFD) score.
 By meticulously following this methodology, we utilize GPT-2 to select 8,000 data samples.
-Deita [29]
-is an open-sourced project designed to facilitate Automatic Data Selection for instruction tuning in LLMs.
+Deita [33] is an open-sourced project designed to facilitate Automatic Data Selection for instruction tuning in LLMs.
 It delves into the relationships between data complexity, quality, and diversity, and develops a series of methodologies.
@@
 Following its methodology, we gain 8,000 entries as a baseline.
-D.2 Visualization of Baseline’s Selected Data
+E.2 Visualization of Baseline’s Selected Data
 We project all baseline-selected data samples through Llama3.1-8B’s, Qwen2-7B’s and Qwen2.5-7B’s embedding layer and visualize their distributions via t-SNE dimensionality reduction in Fig.
-5, Fig.
-6, and Fig.
-7, respectively.
+6, Fig.
+7, and Fig.
+8, respectively.
 This visualization provides intuitive insights into how different data selection strategies handle domain-mixed data.
@@
 Figure 8: t-SNE visualization of data samples selected by different baselines using Qwen2.5-7B embeddings. While original labels were removed during training, we preserve them for interpretability.
-Appendix E Derivation of Proposition 3.2
-We derive the density ratio by systematically expanding and substituting terms:
-$\frac{q(\mathbf{x})}{p(\mathbf{x})}=\frac{\sum_{k=1}^{K}q(c=k)q(\mathbf{x}|c=k)}{p(\mathbf{x})}$ (10)
-Expand $q(\mathbf{x})$ via the law of total probability over domains.
-$=\frac{\sum_{k=1}^{K}q(c=k)p(\mathbf{x}|c=k)}{p(\mathbf{x})}$ (11)
-Apply Assumption 3.1 ($q(\mathbf{x}|c=k)=p(\mathbf{x}|c=k)$) to unify domain-conditional densities.
-$=\frac{\sum_{k=1}^{K}q(c=k)\frac{p(c=k|\mathbf{x})p(\mathbf{x})}{p(c=k)}}{p(\mathbf{x})}$ (12)
-Express $p(\mathbf{x}|c=k)$ via Bayes’ rule inversion: $p(\mathbf{x}|c=k)=\frac{p(c=k|\mathbf{x})p(\mathbf{x})}{p(c=k)}$.
-$=\sum_{k=1}^{K}\frac{q(c=k)}{p(c=k)}p(c=k|\mathbf{x})\cdot\frac{p(\mathbf{x})}{p(\mathbf{x})}$ (13)
-Factor out $p(\mathbf{x})$ from the numerator and the denominator.
-$=\sum_{k=1}^{K}\underbrace{\frac{q(c=k)}{p(c=k)}}_{\lambda_{k}}p(c=k|\mathbf{x})$ (14)
-Define domain proportion ratio $\lambda_{k}:=\frac{q(c=k)}{p(c=k)}$.
-This demonstrates that the density ratio decomposes linearly into domain-specific components weighted by their proportion shifts ($\lambda_{k}$) and posterior probabilities ($p(c=k|\mathbf{x})$).
 Appendix F Comprehensive Experiments
@@
 The t-SNE visualization of Qwen3’s embeddings (Fig.
-8) demonstrates comparable domain discrimination capabilities.
+9) demonstrates comparable domain discrimination capabilities.
 Figure 9: The t-SNE visualization of embeddings for data samples with different distributions on Qwen3-8B.
@@
 Consistent with observations in other models, the two-stage MLP training framework demonstrated robust stability and convergence, shown in Fig.
-9.
+10.
 Figure 10: Training loss and validation process of the two training stages of DaaR on Qwen3-8B, showing the model gradually converging.
 Evaluation Results of DaaR with Baselines
-Following the identical experimental setup outlined in Appendix B, we conducted experiments on Qwen3, with the results presented in Table 9.
+Following the identical experimental setup outlined in Appendix D, we conducted experiments on Qwen3, with the results presented in Table 10.
 As demonstrated, DaaR achieved the top comprehensive score on Qwen3, surpassing all baselines by up to 4.7%.
@@
 Coding-Major 10.76 57.02 72.78 78.20 12.80 65.20 73.22 52.85
-F.4 Rationale for Entropy over Diversity as the Reward Signal
+F.4 Robustness to Labeling Schemes and Embedding Representations
+A core principle of DaaR is its "model-aware" design, which relies on the model’s own representations (via its embedding layer) and internal data structure (via clustering-based pseudo-labels) rather than external annotations or models.
+This section presents two ablation studies that investigate the validity and effectiveness of these design choices.
+Analysis 1: Model-Aware Pseudo-Labels vs. Ground-Truth Labels
+To analyze the impact of our clustering-based pseudo-labels, we conducted an experiment where we replaced them with human-annotated, ground-truth (GT) domain labels.
+This allowed us to directly compare a "model-aware" perspective against a "human-aware" one within our framework.
+Setup.
+We reran the entire DaaR pipeline, using the GT labels to train the domain discrimination probe and subsequently to compute the predictive entropy for data selection.
+All other components remained identical.
+Observations.
+  • Data Selection Divergence: The choice of labels significantly altered the data selection process.
+As shown in Table 12, the overlap between the data selected using pseudo-labels and GT labels was only around 83-87%, indicating that they prioritize different data subsets.
+    Table 12: Overlap of selected data between using pseudo-labels and ground-truth (GT) labels.
+    Model Overlap (Pseudo vs. GT)
+    Llama3.1-8B 87.3%
+    Qwen2-7B 83.1%
+    Qwen2.5-7B 84.9%
+  • Final Performance: Critically, using GT labels led to a consistent, albeit small, decrease in average end-task performance across all models, as detailed in Table 13.
+This counter-intuitive result suggests that for fine-tuning, an LLM’s internal perception of data domains (captured by pseudo-labels) can be a more effective guide than human-defined categories.
+The GT labels may introduce biases that, while semantically correct to a human, are suboptimal for optimizing the model’s capabilities.
+Table 13: Performance comparison: DaaR using model-aware pseudo-labels vs. GT labels. The "Diff" row shows the performance change when switching from pseudo-labels to GT labels.
+Model Setting nq triviaqa hellaswag gsm8k math mbpp humaneval Avg
+Llama3.1-8B DaaR w/ Pseudo 20.08 64.55 74.88 54.80 15.30 4.70 37.50 38.83
+DaaR w/ GT 22.54 65.52 73.43 53.80 14.35 4.40 36.50 38.65
+Diff +2.46 +0.97 -1.45 -1.00 -0.95 -0.30 -1.00 -0.18
+Qwen2-7B DaaR w/ Pseudo 16.88 57.58 73.03 75.40 38.10 52.00 64.94 53.99
+DaaR w/ GT 18.75 58.35 72.02 73.40 35.75 51.50 64.01 53.40
+Diff +1.87 +0.77 -1.01 -2.00 -2.35 -0.50 -0.93 -0.59
+Qwen2.5-7B DaaR w/ Pseudo 15.83 58.65 72.48 80.20 16.70 64.20 68.29 53.76
+DaaR w/ GT 16.04 58.75 71.87 79.40 16.30 64.25 67.76 53.48
+Diff +0.21 +0.10 -0.61 -0.80 -0.40 +0.05 -0.53 -0.28
+Analysis 2: Internal vs. External Embedding
+We also investigated the choice of using the model’s own embedding layer versus relying on powerful, external SOTA embedding models.
+While external models could provide an alternative "semantic space," we evaluated the trade-offs.
+Setup.
+We processed our dataset using two SOTA external embedding models (GTE-Qwen2-7B and Qwen3-8B-Embedding) and compared the time cost against using the native embedding layer of the base LLM (Qwen2-7B).
+Observations.
+While the external models were also capable of producing separable domain clusters, they introduced prohibitive computational overhead.
+As shown in Table 14, using external models was 175x to 192x slower and led to out-of-memory (OOM) errors on a 40GB GPU.
+This highlights the significant efficiency advantage of DaaR’s internal approach.
+Table 14: Time cost for extracting embeddings from a 40K dataset.
+Model/Layer Time Cost (Relative Speed)
+Qwen2-7B (Internal Embedding Layer) 9.23s (1x)
+GTE-Qwen2-7B (External Model) 1662s (175x slower)
+Qwen3-8B-Embedding (External Model) 1774s (192x slower)
+F.5 Analysis of End-to-End Stability
+Given that DaaR is a multi-stage pipeline, it is important to analyze its end-to-end stability and robustness against potential compounding errors.
+To this end, we conducted three fully independent, end-to-end experimental runs of our entire framework and evaluated the consistency at critical stages of the pipeline.
+We present results for Qwen2-7B as a representative case.
+Stability of Centroid Generation
+The initial stage of our pipeline involves generating domain centroids.
+We measured the pairwise cosine similarities of the final centroid embeddings across the three independent runs.
+As shown in Table 15, the resulting centroids are remarkably consistent, with all similarity scores exceeding 0.98.
+This indicates that the foundational stage of our method is highly stable and is not a significant source of variance.
+Table 15: Pairwise cosine similarity of domain centroids across three independent runs.
+Domain Sim(Run 1-2) Sim(Run 1-3) Sim(Run 2-3)
+Common Sense 0.9912 0.9895 0.9921
+Reasoning 0.9845 0.9911 0.9887
+Mathematics 0.9880 0.9803 0.9854
+Coding 0.9905 0.9853 0.9918
+Stability of Final Data Selection
+We then examined whether minor variations in the initial stages propagate to the data selection outcome.
+We calculated the overlap ratio of the top 20% (8,000) selected samples across the three runs.
+Table 16 shows an average overlap of 96.4%, confirming that the final data selection is highly robust and largely deterministic, with minimal amplification of randomness.
+Table 16: Overlap ratio of selected data subsets across three independent runs.
+Metric Overlap (Run 1-2) Overlap (Run 1-3) Overlap (Run 2-3)
+Overlap Ratio 95.7% 97.3% 96.1%
+Consistency of Final Performance
+Finally, the end-to-end stability of the framework is reflected in the final benchmark performance.
+As reported in Table 17, the results across the three runs are highly consistent, with a minimal standard deviation of 0.08 in the average score.
+This provides strong empirical evidence that our framework is robust and its outcomes are reliable, without significant compounding error effects.
+Table 17: End-task performance across three independent runs on Qwen2-7B. Scores are reported with mean and standard deviation.
+Run nq triviaqa hellaswag gsm8k math mbpp humaneval Avg
+Run 1 17.92 56.78 72.91 75.00 39.60 51.40 64.80 54.06
+Run 2 15.84 58.38 73.14 75.80 36.60 52.60 65.07 53.92
+Run 3 15.18 57.96 72.99 76.70 38.80 51.40 65.24 54.04
+Avg 16.31 57.71 73.01 75.83 38.33 51.80 65.04 54.01
+Std. Dev. (1.43) (0.83) (0.12) (0.85) (1.55) (0.69) (0.22) (0.08)
+F.6 Rationale for Entropy over Diversity as the Reward Signal
 As discussed in Section 4, the explicit diversity metric defined in Eq.
@@
 Experimental result (Fig.
-10) demonstrates persistent convergence challenges, evidenced by consistently high training loss and subpar classification accuracy capped at 34%.
+11) demonstrates persistent convergence challenges, evidenced by consistently high training loss and subpar classification accuracy capped at 34%.
 These findings substantiate our hypothesis that LLMs fail to learn cross-sample characteristics, thereby necessitating entropy-based proxy metrics for effective reward formulation.
@@
 Using Qwen2-7B as an example, we conduct t-SNE visualizations of Layer-5, Layer-10, Layer-15, Layer-20, Layer-25, and the last hidden layer, based on the average token vectors, as shown in Fig.
-12.
+13.
 Figure 13: t-SNE visualization of mean token embeddings across Layer-5, Layer-10, Layer-15, Layer-20, Layer-25, and the last hidden layer of Qwen2-7B.
@@
 The training dynamics across the first ten layers are visualized in Fig.
-13, demonstrating consistent accuracy within the 91.2–93.6% range, verifying that various layers can effectively support DaaR first-stage training.
+14, demonstrating consistent accuracy within the 91.2–93.6% range, verifying that various layers can effectively support DaaR first-stage training.
 Figure 14: Validation loss and accuracy over steps on different layers.
@@
 However, the training results showed persistently high training loss and a validation accuracy of only about 0.6 (shown in Fig.
-14), indicating suboptimal learning performance.
+15), indicating suboptimal learning performance.
 Furthermore, related studies indicate that LLMs can learn semantic understanding and perception in the early layers [45], consequently, we select Layer-3 for DaaR.
@@
 Reasoning 135 61 44 15
-Mathematics $>1000$ 804 236 25
+Mathematics >1000>1000 804 236 25
 Coding 74 21 16 13
@@
 Coding 0.9646 0.9740 0.9691
-G.7 Cost-Efficiency and Flexibility
+G.7 Ablation of Model Scales
+To assess the generalizability of DaaR beyond the 7-8B parameter range, we conducted additional experiments on models of different scales: a smaller model, Llama3.2-3B, and a larger model, Qwen2.5-14B.
+Due to computational constraints, we performed a targeted comparison of DaaR against two key baselines: Raw (using the original, unselected data) and Random (a strong baseline involving random data selection).
+The results are presented in Table 23.
+On the larger Qwen2.5-14B model, DaaR achieves the highest average score (61.54), outperforming the strong Random baseline.
+Similarly, on the smaller Llama3.2-3B model, DaaR (29.64) maintains a consistent advantage over both Raw and Random baselines.
+These findings provide evidence that the effectiveness of DaaR is not confined to a specific model size and generalizes to both smaller and larger language models.
+Table 23: Performance comparison on Llama3.2-3B and Qwen2.5-14B, demonstrating the generalization capability of DaaR across different model scales.
+Model Method NQ TriviaQA Hellaswag GSM8K MATH MBPP HumanEval Avg
+Qwen2.5-14B Raw 10.89 66.56 76.86 86.00 19.70 5.40 78.66 49.15
+Random 20.06 65.80 76.76 86.40 36.90 64.20 75.46 60.80
+DaaR 19.73 66.26 77.19 85.00 39.10 66.80 76.69 61.54
+LLaMA3.2-3B Raw 7.62 53.10 68.77 26.60 4.50 3.80 23.17 26.79
+Random 16.03 53.56 68.46 28.70 6.15 4.65 29.12 29.52
+DaaR 15.13 53.79 68.23 29.10 5.63 5.30 30.34 29.64
+G.8 Ablation of Downstream Sample Injection in Seed Generation
+The seed generation process described in Section 4.1 includes a minor injection of downstream task samples.
+This component is intended to provide an initial diversity signal to accelerate the subsequent augmentation phase.
+To rigorously evaluate its impact, we present an ablation study comparing our standard method against a variant where this injection is completely removed.
+It is important to note that in all settings, our experimental protocol maintains strict data partitioning.
+The downstream samples used for seeding, when present, are drawn from a set entirely disjoint from the datasets used for final evaluation.
+Our analysis focuses on three key aspects: the impact on the final centroid representations, the effect on end-task model performance, and the change in data generation efficiency.
+Impact on Centroid Similarity
+We measured the cosine similarity between domain centroids generated with our standard method (w/ injection) and the ablation setting (w/o injection).
+The results in Table 24 show that the centroids are nearly identical, with an average similarity score of 0.987.
+This indicates that the minor injection has a negligible impact on the final learned domain representations.
+Table 24: Cosine similarity of domain centroids generated with and without downstream seed injection, evaluated on Qwen2-7B.
+Domain Similarity (w/ vs. w/o Injection)
+Common Sense 0.9922
+Reasoning 0.9865
+Mathematics 0.9843
+Coding 0.9878
+Average 0.9877
+Impact on Final Performance
+As shown in Table 25, the end-task performance of models trained using data selected by DaaR is statistically indistinguishable between the two settings.
+This demonstrates that the effectiveness of our method is not dependent on the seed injection.
+Table 25: End-task performance comparison between DaaR with and without downstream seed injection. The results show statistically insignificant differences.
+Model Setting nq triviaqa hellaswag gsm8k math mbpp humaneval Avg
+Llama3.1-8B DaaR w/ Inject 20.08 64.55 74.88 54.80 15.30 4.70 37.50 38.83
+DaaR w/o Inject 20.39 64.80 76.05 55.40 13.65 5.75 36.48 38.93
+Qwen2-7B DaaR w/ Inject 16.88 57.58 73.03 75.40 38.10 52.00 64.94 53.99
+DaaR w/o Inject 16.22 58.33 73.41 75.20 36.37 52.95 65.14 53.95
+Qwen2.5-7B DaaR w/ Inject 15.83 58.65 72.48 80.20 16.70 64.20 68.29 53.76
+DaaR w/o Inject 14.91 58.32 72.55 79.70 16.30 63.70 70.65 53.73
+Impact on Generation Efficiency
+While not critical for performance, the injection significantly improves the efficiency of the diversity augmentation process.
+As detailed in Table 26, removing the initial seed diversity required approximately 4 times more generation attempts to meet the same diversity threshold ($\tau$).
+This confirms its functional role as a practical accelerator.
+Table 26: Number of generation attempts required to meet the diversity threshold during seed augmentation.
+Domain Attempts (w/ Injection) Attempts (w/o Injection)
+Common Sense 31 134
+Reasoning 61 248
+Mathematics 236 819
+Coding 21 105
+G.9 Cost-Efficiency and Flexibility
 Compared to baseline methods requiring GPT-based evaluators (Alpagasus, Instag) or full-LLM inference (Deita), our approach achieves superior efficiency through data-model co-optimizations.
@@
 Experimental results on 5,000 samples demonstrate this efficiency advantage: our method requires approximately 40 minutes for sample-level inference, whereas conventional LLM inference approaches demand nearly two hours.
-G.8 Stability of DaaR
+G.10 Stability of DaaR
 Our dual-MLP design of DaaR strategically separates entropy regression (second MLP) from neural softmax-based measurement (first MLP), both showing stable convergence in Fig 2.
@@
 We present the t-SNE visualization of data samples from Llama3.1-8, Qwen2-7B and Qwen2.5-7B with different distributions in Fig.
-15, Fig.
+16, Fig.
 17 and Fig.
-17.
+18.
 It is evident that, despite differences in architecture or model, the method of filtering data through Inter-Diversity and Intra-Diversity is effective.
@@
 (a) The data pool of all 40K samples, (b) Randomly selected subset, (c) Distribution of data farthest from other domain centroids on Inter-Diversity, (d) Distribution of data closest to other domain centroids on Inter-Diversity, (e) Distribution of data closest to its own domain centroid on Inter-Diversity, (f) Distribution of data farthest from its own domain centroid on Inter-Diversity.
-Figure 16: The t-SNE visualization of embeddings for data samples with different distributions on Qwen2-7B.
-(a) The data pool of all 40K samples, (b) Randomly selected subset, (c) Distribution of data farthest from other domain centroids on Inter-Diversity, (d) Distribution of data closest to other domain centroids on Inter-Diversity, (e) Distribution of data closest to its own domain centroid on Inter-Diversity, (f) Distribution of data farthest from its own domain centroid on Inter-Diversity.
-Figure 17: The t-SNE visualization of embeddings for data samples with different distributions on Qwen2.5-7B.
-(a) The data pool of all 40K samples, (b) Randomly selected subset, (c) Distribution of data farthest from other domain centroids on Inter-Diversity, (d) Distribution of data closest to other domain centroids on Inter-Diversity, (e) Distribution of data closest to its own domain centroid on Inter-Diversity, (f) Distribution of data farthest from its own domain centroid on Inter-Diversity.
+Figure 17: The t-SNE visualization of embeddings for data samples on Qwen2-7B.
+Figure 18: The t-SNE visualization of embeddings for data samples on Qwen2.5-7B.
 H.2 Visualization of Inter- & Intra-Diversity Distribution Data
@@
 19 and Fig.
-19.
+20.
 As seen in the figures, the data gradually transitions from domain-aware diverse to domain-aware closed, indicating that our data construction method effectively controls the distribution of different data.
-However, it is important to note that t-SNE dimensionality reduction is used only for visualization purposes; during the reduction process, positional information is lost, and it cannot fully capture the true high-dimensional relationships between the data.
 Figure 19: Data visualization (t-SNE) based on different Inter-Diversity distributions on Qwen2-7B.
@@
 H.3 Comprehensive Training Dynamics
-We present the training loss and validation results for two distinct models, Llama3.1-8B and Qwen2.5-7B, in Fig 20.
+We present the training and validation dynamics for Llama3.1-8B and Qwen2.5-7B, in Fig 21.
 It can be observed that across different models, the training process of DaaR method consistently ensures gradual convergence, achieving high domain predictability and calibrated entropy fitting.